### v1

# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

Abstract
The energy correction associated with the self-energy diagram is the leading (in magnitude) contribution to the Lamb shift in hydrogen-like ions. All conventional approaches to this correction rely on partial-wave expansions, which are a stumbling block limiting accuracy. To elucidate an issue, we perform a comparative analysis of partial-wave-expansion convergence in two gauges: Feynman and Coulomb. Some tricks for improving convergence are also discussed.
I Introduction
Accurate theoretical predictions for simple atomic systems must include QED effects mohr_main; indelicato_qed_2019; SHABAEV202494; volotka_progress_2013. One of the leading QED corrections to the bound-electron energy levels corresponds to the one-electron self-energy (SE) diagram, shown in Fig. 1. A high-precision evaluation of the corresponding contribution holds significance, however, not only for studies of the Lamb shift and transition energies in hydrogen-like ions and other few-electron systems Stohlker; PhysRevLett.94.223001; PhysRevLett.66.1434; PhysRevLett.91.073202; PhysRevLett.95.233003; Trassinelli_2009; Trassinelli_2011; Loetzsch2024 (for related theory, see, e.g., Refs. PhysRevLett.97.253004; PhysRevA.77.032501; PhysRevA.83.012504; se_yerokh; PhysRevA.71.062104; PhysRevA.100.062506), but also for consideration of a wide range of radiative corrections. Accurate calculations of this contribution is a cornerstone in the ab initio treatment of the two-electron SE diagrams PhysRevLett.76.204; PhysRevA.60.3522; PhysRevA.63.052507, QED corrections to the $g$-factor yerokhin_self-energy_2002; PhysRevA.62.032510, quadratic Zeeman splitting agababaev2025qedeffectsquadraticzeeman, and hyperfine structure PhysRevA.64.012506; PhysRevA.63.032506; PhysRevLett.100.163001; PhysRevA.81.012502. In addition, the SE correction is essential when studying the E1 E1_Kozlov and M1 M1_Volotka transition amplitudes. In these applications it is often necessary to calculate off-diagonal matrix elements of the SE operator. However, one-electron (diagonal) SE is a simple example that one can and should use to experiment in order to understand how to calculate it with the highest precision.
There are two fundamentally different approaches to the SE calculations. The first one is based on the nonrelativistic formulation of QED and implies an expansion in $\alpha Z$ caswell ($\alpha$ is the fine-structure constant and $Z$ is the nuclear charge). It is suitable for light nuclei, in which this parameter is much less than unity. In contrast, the second approach is non-perturbative in $\alpha Z$, i.e., all calculations are carried out to all orders in this parameter, see, e.g., Refs. SHABAEV2002119; mohr_main. These approaches can be applied as complementary RevModPhys.97.025002. In the present paper, we employ the second approach, also known as the Furry picture of QED PhysRev.81.115, to calculate the SE correction.
In this regard, an issue of representing the bound-electron propagator in calculations naturally arises. Unlike the free-electron propagator, the propagator in a binding potential does not have a closed-form expression. However, when the spherical symmetry is present, as is the case in the atomic calculations, such the propagator can be constructed, e.g., using a partial-wave (PW) expansion. With this in mind, we use two state-of-the-art approaches: the Green’s function (GF) method MOHR197426; MOHR197452; PhysRevA.38.5066, and the finite-basis-set (FBS) method PhysRevA.37.307; DKB.
Within the first approach, the bound-electron propagator is represented in terms of the Dirac-equation solutions bounded at infinity and at origin. For the point-nucleus Coulomb potential, the corresponding solutions are expressed analytically via the Whittaker functions PhysRev.101.843. For more realistic models of nuclear-charge distribution, e.g., the Fermi model, or other spherically symmetric potentials, e.g., some local screening potentials, these solutions can be found numerically. This approach enables one to perform calculations with a high numerical accuracy when large numerical cancellations occur or when the PW expansion does not converge rapidly. Note that the Green’s function, considered as a function of any of its radial arguments, is discontinuous, when the arguments coincide. Therefore, obtaining accurate results for the matrix elements of the Green’s function requires special care.
Within the second approach, the electron Green’s function is built using the FBS method. Basis sets for the Dirac equation could be constructed, e.g., from the B splines J_Sapirstein_1996; johnson or Gaussians boys_gauss; se_gauss. The corresponding numerical procedures can easily incorporate any spherically symmetric potential. However, for systems that do not have spherical symmetry, the FBSs can also be readily prepared PhysRevA.106.012813; DIRAC25. The advantage of this method is that it allows one to easily separate out and exculde, if necessary, the contribution of a specific bound state to the Green’s function. Note also that this method provides an approximation to the Green’s function, which is a continuous function of the radial arguments. Nevertheless, the basis-set method has some important drawbacks when compared to the GF approach. Because of an additional parameter, the number of basis functions $N$, the corresponding results should be extrapolated to $N\rightarrow\infty$. This sets a limit on the computational accuracy. Furthermore, the number of partial waves, usually considered in practical calculations by means of the FBS method is limited and inferior to that in the Green’s function method sym12050800. In this paper we use the FBS method with the basis obtained from B splines within the double-kinetic-balance (DKB) approach DKB. The latter method excludes the so-called spurious states and establishes the correct asymptotics of wave functions in the non-relativistic limit. The effectiveness of this basis as applied to the calculations of the SE correction has been demonstrated, e.g., in Ref. DKB.
The SE diagram contains ultraviolet (UV) divergences. The conventional approach to renormalizing these divergences, which is used in the present work, involves the potential expansion of the bound-electron propagator snyderman. Within this expansion, three contributions naturally arise: zero-, one-, and many-potential terms. Only the zero- and one-potential terms, which correspond to the two first terms of the potential expansion, are UV divergent. They are treated separately in momentum space. The many-potential term, which is given by a remainder of the potential expansion, is considered in coordinate space using the PW expansion. The truncation of the latter and subsequent estimation of the residual is the main source of numerical uncertainty of the result.
As is well known, the total SE contribution is gauge invariant, unlike the individual terms of its potential expansion. Therefore, an intriguing idea is to compare the magnitude of these terms in different gauges. Traditionally, the SE correction is calculated in the Feynman gauge, see, e.g., Ref. base_1999. In the Coulomb gauge, similar calculations were carried out in hen; yer_2025. These works suggest that the many-potential contribution is significantly smaller in the Coulomb gauge than in the Feynman one. However, a smaller absolute value does not imply a better convergence of the PW expansions. Therefore, an important task is to conduct a thorough comparative analysis of many-potential contributions in both gauges. Note that other covariant gauges, aside from the Feynman one, can also be applied to the SE calculations. For example, the Fried-Yennie PhysRevD.47.3647 gauge was also considered in Ref. yer_2025. In the present work, however, we restrict our consideration only to the Feynman and Coulomb gauges.
Another important task in the SE calculations is to somehow accelerate the convergence of employed PW expansions. It is reasonable to assume that the poor convergence of the many-potential contribution is due to the leading term of its potential expansion. Therefore, the first natural idea is to additionally subtract the two-potential term from it, which should improve the PW-expansion convergence of the remainder. The subtracted term has to be evaluated separately with high precision and added to the final result. In the following, we will refer to this approach as the two-potential scheme. It should be noted, however, that the direct calculation of the two-potential term in momentum space in a closed form, i.e., not resorting to the PW expansion, turns out to be a challenging problem PhysRevLett.100.163001; Yer2010.
Another acceleration-convergence approach, which we will further refer to as the Sapirstein-Cheng (SC) scheme, was proposed recently in Ref. sap_orig. Within the framework of this promising scheme, an approximation to the two-potential term, the quasi-two-potential contribution, can be subtracted from the many-potential term instead. In contrast to the exact two-potential term, its quasi-two-potential counterpart can be easily calculated in momentum space. For the point-nucleus case, calculations of the SE correction using this acceleration scheme have been performed in Ref. yer_2025 for different gauges. However, it is also noteworthy to examine the case of an extended nuclues.
We also note another acceleration-convergence approach developed earlier for the SE correction to the Lamb shift in Ref. PhysRevA.72.042502, which is referred to as the Yerokhin, Pachucki, and Shabaev (YPS) scheme. It also improves the convergence by subtracting an approximation for the many-potential contribution, that can be calculated in a closed form. The YPS scheme was initially created for the calculations in the Feynman gauge, and was later generalized to the Coulomb-gauge case in Ref. yer_2025. Judging by the results of the latter paper, it is an extremely powerful scheme, however, its drawback is that it is difficult to generalize it to more complex diagrams, unlike other acceleration-convergence approaches. For instance, the two-potential scheme has recently been applied in the QED calculations of the quadratic Zeeman effect agababaev2025qedeffectsquadraticzeeman, and some modifications of the SC scheme have been applied to the calculations of the two-electron SE sap_alexey and two-loop SE PhysRevLett.133.251803; PhysRevA.111.042820 contributions. Due to this fact, we do not consider the YPS scheme in the current study.
In this paper, we thoroughly examine the application of several different schemes to the SE calculations in both the Coulomb and Feynman gauges. We conduct an independent of Ref. yer_2025 analysis, consider in details the issue of the PW-expansion convergence, perform the SE-correction calculation for nonpoint nuclei, and collect convenient formulas that allow one to make the corresponding calculations.
Throughout this article the relativistic units ($\hbar=c=1$) are used. We use roman style ($\rm{p}$) for four-vectors, bold style ($\bf{p}$) for space vectors, and italic style ($p$) for scalars.
II Self-energy contribution
II.1 Basic formalism
Within the framework of the Furry picture, the description of an atom or ion starts with the Dirac equation:
$\left[-i(\boldsymbol{\alpha}\cdot\boldsymbol{\nabla})+\beta m+V(\mathbf{r})\right]|a\rangle=\varepsilon_{a}|a\rangle,$ (1)
where $|a\rangle$ and $\varepsilon_{a}$ are the Dirac wave function and energy, $\boldsymbol{\alpha}$ and $\beta$ are the Dirac matrices, $V(\mathbf{r})$ is a binding potential, which is a function of the position vector $\mathbf{r}$. In the following, we consider only hydrogen-like ions for which the potential is spherically symmetric, $V(\mathbf{r})=V(r)$, $r=|\mathbf{r}|$. Then, the solution of Eq. (1) can be represented in the form:
$|a\rangle\leftrightarrow\psi_{a}(\mathbf{r})={\begin{pmatrix}g_{a}(r)\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{r}})\\ if_{a}(r)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{r}})\end{pmatrix}}$, (2)
[…]
Figure 1: Self-energy diagram with the related mass counterterm. The double line indicates the electron propagator in the external field of the nucleus, the wavy line denotes the photon propagator, and the cross stands for the counterterm.
The energy shift $\Delta\varepsilon_{a}$ of the bound state $|a\rangle$ due to the first-order self-energy correction, which is graphically represented in Fig. 1, is given base_1999 by the real part of the expression:
$\begin{gathered}\Delta\varepsilon_{a}=2i\alpha\int_{-\infty}^{\infty}d\omega\int d^{3}\mathbf{r}_{1}d^{3}\mathbf{r}_{2}\psi^{{\dagger}}_{a}(\mathbf{r}_{1})\alpha^{\mu}\\ \times G(\varepsilon_{a}-\omega,\mathbf{r}_{1},\mathbf{r}_{2})\alpha^{\nu}\psi_{a}(\mathbf{r}_{2})D_{\mu\nu}(\omega,\mathbf{r}_{12})\\ -\delta m\int d^{3}\mathbf{r}{\psi}^{\dagger}_{a}(\mathbf{r})\beta\psi_{a}(\mathbf{r}),\end{gathered}$ (3)
where $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$, $D_{\mu\nu}(\omega,\mathbf{r}_{12})$ is the photon propagator, $\alpha_{\mu}=(1,{\boldsymbol{\alpha}})$, $G(\varepsilon,\mathbf{r}_{1},\mathbf{r}_{2})$ is the bound-electron Green’s function, and $\delta m$ is the mass counterterm.
Let us define the gauge-dependent operator $I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ according to
$I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=e^{2}\alpha^{\nu}\alpha^{\mu}D_{\mu\nu}(\omega,\mathbf{r}_{12}).$ (4)
In the Feynman and Coulomb gauges, this operator has the forms $I_{F}$ and $I_{C}$, respectively:
$I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[1-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\right]\frac{e^{i\hat{\omega}r_{12}}}{r_{12}},$ (5)
$I_{C}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[\frac{1}{r_{12}}-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\frac{e^{i\hat{\omega}r_{12}}}{r_{12}}+\frac{(\boldsymbol{\alpha}_{1}\cdot\boldsymbol{\nabla}_{1})(\boldsymbol{\alpha}_{2}\cdot\boldsymbol{\nabla}_{2})}{\hat{\omega}^{2}}\frac{e^{i\hat{\omega}r_{12}}-1}{r_{12}}\right],$ (6)
where $r_{12}=|\mathbf{r}_{12}|$, $\hat{\omega}=\sqrt{\omega^{2}+i0}$, and the branch of the square root is fixed with the condition $\Im\sqrt{\omega^{2}+i0}>0$.
In its turn, the bound-electron propagator can be expressed using the spectral representation:
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{n}\frac{\psi_{n}(\mathbf{r}_{1})\psi_{n}^{\dagger}(\mathbf{r}_{2})}{\omega-\varepsilon_{n}^{-}},$ (7)
where the summation runs over the complete Dirac spectrum and $\varepsilon_{n}^{-}=\varepsilon_{n}(1-i0)$. In shortened form, the formula (7) can be written as follows:
$G(\omega)=\sum_{n}\frac{|n\rangle\langle n|}{\omega-\varepsilon_{n}^{-}}.$ (8)
Then, the energy shift (3) can be expressed as:
$\Delta\varepsilon_{a}=\langle a|\gamma^{0}\Big[\Sigma(\varepsilon_{a})-\delta m\Big]|a\rangle,$ (9)
where the diagonal matrix element of the one-loop self-energy operator $\Sigma(E)$ is given by:
$\langle a|\gamma^{0}\Sigma(E)|a\rangle=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle an|I(\omega)|na\rangle}{E-\omega-\varepsilon_{n}^{-}},$ (10)
and $\gamma^{0}\equiv\beta$. The expression (10) contains the UV divergences. To properly treat them, we expand the bound-electron propagator in powers of the binding potential blundell_snyderman. The potential expansion can conveniently be written as:
$\begin{gathered}G(\omega)=\sum_{f}\frac{|f\rangle\langle f|}{\omega-\varepsilon_{f}^{-}}+\sum_{f_{1},f_{2}}\frac{|f_{1}\rangle\langle f_{1}|V|f_{2}\rangle\langle f_{2}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{f_{2}}^{-})}+\\ +\sum_{f_{1},f_{2},n}\frac{|f_{1}\rangle\langle f_{1}|V|n\rangle\langle n|V|f_{2}\rangle\langle f_{2}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{n}^{-})(\omega-\varepsilon_{f_{2}}^{-})},\end{gathered}$ (11)
where the states $|n\rangle$ correspond to the bound-electron spectrum, the states $|f\rangle$, $|f_{1}\rangle$, and $|f_{2}\rangle$ correspond to the free-electron spectra, and the last term is expressed via the bound-electron Green’s function $G(\omega)$ itself and encompasses an infinite number of terms corresponding to the two or more interactions with the nucleus. As before, all summations in Eq. (11), are carried out over the complete spectra for the corresponding Dirac hamiltonians.
Figure 2: Potential expansion of the self-energy diagram. The dashed line ended by a rhombus denotes the interaction with the nuclear field. The single line denotes the free-electron propagator. The mass counterterm is not shown.
Substituting the potential expansion (11) into Eq. (10) splits the SE diagram into three terms, as shown graphically in Fig. 2. These contributions are commonly referred to as the zero-, one-, and many-potential terms, respectively. The zero- and one-potential terms diverge, whereas the many-potential term is UV finite. The UV-divergent terms have to be renormalized along with the mass counterterm. For this aim, they are considered in momentum space, where the dimensional regularization is applied to isolate and eliminate the UV divergences. This renormalization procedure yields UV-finite contributions $\Delta\varepsilon_{a}^{\text{0p}}$ and $\Delta\varepsilon_{a}^{\text{1p}}$ to the Lamb shift. The many-potential term $\Delta\varepsilon_{a}^{\text{Mp}}$ is calculated in the coordinate space. Therefore, the SE correction to the energy levels is expressed as the gauge-invariant sum of three contributions:
$\begin{gathered}\Delta\varepsilon_{a}=\Delta\varepsilon_{a}^{\text{0p}}+\Delta\varepsilon_{a}^{\text{1p}}+\Delta\varepsilon_{a}^{\text{Mp}}.\end{gathered}$ (12)
II.2 Zero- and one-potential terms
The Fourier transform of the wave function (1), which solves the Dirac equation (1) in a spherically-symmetric potential, reads as:
$\psi_{a}(\mathbf{p})=\int d^{3}{\mathbf{r}}e^{-i({\mathbf{p\cdot r}})}\psi_{a}({\mathbf{r}})=i^{-l_{a}}{\begin{pmatrix}\widetilde{g}_{a}({p})\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{p}})\\ \widetilde{f}_{a}(p)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{p}})\end{pmatrix}}$, (13)
where $\mathbf{\hat{p}}=\mathbf{p}/p$ and for a bound state $\widetilde{g}_{a}(p)=\widetilde{g}_{{n_{r}}_{a},\kappa_{a}}(p)$ and $\widetilde{f}_{a}(p)=\widetilde{f}_{{n_{r}}_{a},\kappa_{a}}(p)$. For the point-nucleus case, there are analytical expressions for $\widetilde{g}_{a}(p)$ and $\widetilde{f}_{a}(p)$, see, e.g., mohr_main. For an arbitrary potential, they can be obtained numerically according to:
$\widetilde{g}_{a}(p)=4\pi\int_{0}^{\infty}drr^{2}j_{l_{a}}(pr)g_{a}(r),$ (14)
$\widetilde{f}_{a}(p)=-4\pi\frac{\kappa_{a}}{|\kappa_{a}|}\int_{0}^{\infty}drr^{2}j_{2j_{a}-l_{a}}(pr)f_{a}(r),$ (15)
[…]
The zero-potential term is expressed as:
$\begin{gathered}\Delta\varepsilon_{a}^{\text{0p}}=\int\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\psi_{a}^{\dagger}(\mathbf{p})\gamma^{0}\Sigma^{(0)}_{R}(\mathrm{p})\psi_{a}(\mathbf{p}),\end{gathered}$ (16)
[…]
$\Delta\varepsilon_{a}^{\text{0p}}=\frac{\alpha}{4\pi}\int_{0}^{\infty}\frac{p^{2}dp}{(2\pi)^{3}}\{a(\varepsilon_{a},p)[\widetilde{g}_{a}^{2}-\widetilde{f}_{a}^{2}]\\ +b(\varepsilon_{a},p)[\varepsilon_{a}(\widetilde{g}_{a}^{2}+\widetilde{f}_{a}^{2})+2p\widetilde{g}_{a}\widetilde{f}_{a}]+c(\varepsilon_{a},p)[\widetilde{g}_{a}^{2}+\widetilde{f}_{a}^{2}]\},$ (17)
where the dependence of $\widetilde{g}_{a}$ and $\widetilde{f}_{a}$ on $p$ is omitted for brevity.
The one-potential term is expressed as:
$\begin{gathered}\Delta\varepsilon_{a}^{\text{1p}}=\int\frac{d^{3}\mathbf{p}^{\prime}d^{3}\mathbf{p}}{(2\pi)^{6}}\psi_{a}^{\dagger}(\mathbf{p}^{\prime})\gamma^{0}\Gamma^{(0)}_{R}(\mathrm{p}^{\prime},\mathrm{p})\widetilde{V}(|\mathbf{p}^{\prime}-\mathbf{p}|)\psi_{a}(\mathbf{p}),\end{gathered}$ (18)
where $\Gamma^{(0)}_{R}$ is the renormalized free-particle vertex function, the explicit form of which with the gauge-dependent coefficients $A$-$G_{2}$ is given in the Appendix B in Eq. (60), $\widetilde{V}$ is the Fourier transform of the binding potential, and $\mathrm{p}=(\varepsilon_{a},\mathbf{p})$ and $\mathrm{p}^{\prime}=(\varepsilon_{a},\mathbf{p}^{\prime})$. In this case, the angular integration can be performed using the identity:
$\frac{1}{2j+1}\sum_{m}\chi_{\kappa m}^{\dagger}(\mathbf{\hat{p}}^{\prime})\chi_{\kappa m}(\mathbf{\hat{p}})=\frac{1}{4\pi}P_{l}(z),$ (19)
where $z$ is the cosine of the angle between the vectors $\mathbf{p}$, $\mathbf{p}^{\prime}$, namely, $z=(\mathbf{\hat{p}}\cdot\mathbf{\hat{p}^{\prime}})$, $P_{l}$ is the Legendre polynomial. The final expression for the one-potential term reads as
$\begin{gathered}\Delta\varepsilon_{a}^{\text{1p}}=\frac{\alpha}{2(2\pi)^{6}}\int_{0}^{\infty}p^{\prime 2}dp^{\prime}\int_{0}^{\infty}p^{2}dp\int_{-1}^{1}dz\widetilde{V}(q)\{X_{1}P_{l_{a}}(z)+X_{2}P_{2j_{a}-l_{a}}(z)\},\end{gathered}$ (20)
where $q^{2}=p^{2}+p^{\prime 2}-2pp^{\prime}z$. The coefficients $X_{1}$ and $X_{2}$ are defined as:
$\begin{gathered}X_{1}=A\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{1}^{\prime}\widetilde{g}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{g}^{\prime}_{a}K_{1}\\ +DK_{1}K_{1}^{\prime}+\varepsilon_{a}(H_{1}+H_{2})\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+G_{1}K_{1}^{\prime}\widetilde{g}_{a}+G_{2}\widetilde{g}^{\prime}_{a}K_{1},\\ X_{2}=A\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{2}^{\prime}\widetilde{f}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{f}^{\prime}_{a}K_{2}\\ +DK_{2}K_{2}^{\prime}-\varepsilon_{a}(H_{1}+H_{2})\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}-G_{1}K_{2}^{\prime}\widetilde{f}_{a}-G_{2}\widetilde{f}^{\prime}_{a}K_{2},\\ K_{1}=\varepsilon_{a}\widetilde{g}_{a}+p\widetilde{f}_{a},\;K_{1}^{\prime}=\varepsilon_{a}\widetilde{g}^{\prime}_{a}+p^{\prime}\widetilde{f}^{\prime}_{a},\;K_{2}=\varepsilon_{a}\widetilde{f}_{a}+p\widetilde{g}_{a},\;K_{2}^{\prime}=\varepsilon_{a}\widetilde{f}^{\prime}_{a}+p^{\prime}\widetilde{g}^{\prime}_{a}.\end{gathered}$ (21)
For brevity, the dependence of wave functions on $p$ and $p^{\prime}$ is omitted. For the functions of $p^{\prime}$, an additional prime is added: $\widetilde{g}^{\prime}_{a}=\widetilde{g}_{a}(p^{\prime})$ and $\widetilde{f}^{\prime}_{a}=\widetilde{f}_{a}(p^{\prime})$.
II.3 Many-potential term
As was said above, the many-potential contribution is calculated in coordinate space using the PW expansion. Therefore, let us start with the PW expansion of the bound-electron propagator (7), inspired by the explicit form of the Dirac-equation solution (2):
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{\kappa_{n}}{\begin{pmatrix}G_{{\kappa}_{n}}^{11}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{++}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})&-iG_{\kappa_{n}}^{12}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{+-}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})\\ iG_{\kappa_{n}}^{21}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{-+}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})&G_{\kappa_{n}}^{22}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{--}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})\end{pmatrix}}$, (22)
where $\pi^{\pm\pm}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})=\sum_{m_{n}}\chi_{\pm{\kappa_{n}},{m_{n}}}(\mathbf{\hat{r}}_{1})\chi_{\pm{\kappa_{n}},{m_{n}}}^{\dagger}(\mathbf{\hat{r}}_{2})$ and the components of the radial Green’s function $G_{\kappa_{n}}^{ij}$ are obtained in different ways depending on the method used. For fixed $\kappa_{n}$, the radial Dirac-equation spectrum is non-degenerate, so it is convenient to introduce the index $i_{n}$ that enumerates the solutions with the given angular symmetry. For example, for $G^{11}$ Eq. (7) results in
$G_{\kappa_{n}}^{11}(\omega,{r_{1},r_{2}})=\sum_{i_{n}}\frac{g_{{i_{n}},{\kappa_{n}}}(r_{1})g_{{i_{n}},{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}.$ (23)
The index $i_{n}$ in Eq. (23) and similar expressions runs over both the positive- and negative-energy Dirac continua as well as all bound states. When using the DKB approach, the complete Dirac spectrum is replaced with the finite set of solutions of the equation (1). In this case, the index $i_{n}$ runs over this finite set. When using the GF approach, the radial Green’s function $G_{\kappa_{n}}^{ij}$ can be constructed from the homogeneous Dirac-equation solutions, bounded at zero, $\phi_{\kappa_{n}}^{0}=(g_{\kappa_{n}}^{0}\;\;f_{\kappa_{n}}^{0})^{\rm{T}}$, and at infinity, $\phi_{\kappa_{n}}^{\infty}=(g_{\kappa_{n}}^{\infty}\;\;f_{\kappa_{n}}^{\infty})^{\rm{T}}$, (see, e.g., Ref. sym12050800):
$G_{{\kappa_{n}}}(\omega,r_{1},r_{2})=\phi_{{\kappa_{n}}}^{\infty}(r_{1}){\phi_{{\kappa_{n}}}^{0}}^{\rm{T}}(r_{2})\theta(r_{1}-r_{2})+\phi_{{\kappa_{n}}}^{0}(r_{1}){\phi_{{\kappa_{n}}}^{\infty}}^{\rm{T}}(r_{2})\theta(r_{2}-r_{1}).$ (24)
Note that the solutions $\phi^{0}$ and $\phi^{\infty}$ are supposed to be normalized so that their Wronskian equals one. Expressions similar to Eqs. (22)-(24) can also be obtained for the free-electron Green’s function, which we will denote as $G^{(0)}$.
With this in mind, we construct the many-potential part of the bound-electron Green’s function which involves two or more interactions with the binding potential and corresponds to the last term in the expression (11):
$G^{(2+)}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\int d^{3}\mathbf{x}\int d^{3}\mathbf{y}G^{(0)}(\omega,\mathbf{r}_{1},\mathbf{x})V(x)G(\omega,\mathbf{x},\mathbf{y})V(y)G^{(0)}(\omega,\mathbf{y},\mathbf{r}_{2}).$ (25)
Since the potential $V$ is assumed to be spherically symmetric, angular integrations in Eq. (25) can be easily carried out. As a result, one obtains the same PW expansion for $G^{(2+)}$ as that for $G$ in Eq. (22). The radial components of $G^{(2+)}$ read as
$G^{(2+)ij}_{\kappa_{n}}(\omega,r_{1},r_{2})=\sum_{k,m}\int dxx^{2}\int dyy^{2}G^{(0)ik}_{\kappa_{n}}(\omega,r_{1},x)V(x)G^{km}_{\kappa_{n}}(\omega,x,y)V(y)G^{(0)mj}_{\kappa_{n}}(\omega,y,r_{2}).$ (26)
In the GF case, because of the $\theta$-functions, radial integrations need to be handled carefully in Eq. (26), see, e.g., Ref. Art_trick. In the framework of the DKB approach, it is convenient to rewrite $G^{(2+)}$ in the form:
$G^{(2+)}(\omega)=\sum_{n}\frac{|\check{n}\rangle\langle\check{n}|}{\omega-\varepsilon_{n}^{-}},$ (27)
where new states $|\check{n}\rangle$ are defined by:
$|\check{n}\rangle=\sum_{f}\frac{|f\rangle\langle f|V|n\rangle}{\omega-\varepsilon_{f}^{-}}\equiv{\begin{pmatrix}\check{g}_{n}(r)\chi_{\kappa_{n}m_{n}}(\mathbf{\hat{r}})\\ i\check{f}_{n}(r)\chi_{-\kappa_{n}m_{n}}(\mathbf{\hat{r}})\end{pmatrix}}$, (28)
and the multi-index $n$, as discussed above, corresponds to the complete set of quantum numbers: $n=(i_{n},\kappa_{n},m_{n})$. Then, using Eq. (28), the many-potential contribution is expressed as:
$\Delta\varepsilon_{a}^{\text{Mp}}=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle a\check{n}|I(\omega)|\check{n}a\rangle}{\varepsilon_{a}-\omega-\varepsilon_{n}^{-}}.$ (29)
The expression (29) can be readily adjusted to be used within the GF approach. For this aim, one need to reverse an analogue of Eq. (23) for $G^{(2+)}$:
$\sum_{i_{n}}\frac{\check{g}_{{i_{n}},{\kappa_{n}}}(r_{1})\check{g}_{{i_{n}},{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}\rightarrow G^{(2+)11}_{\kappa_{n}}(\omega,r_{1},r_{2}).$ (30)
We use the following formula for the matrix element of the operator $I(\omega)$:
[…]
where the reduced matrix elements on the right-hand side do not depend on the angular-momentum projections. Their explicit form in the Feynman and Coulomb gauges is given in Appendix C in Eqs. (71) and (72), respectively. The summation over the projections $m_{n}$ in Eq. (29) can be done analytically, and we obtain:
[…]
The sums over $J$ and $\kappa_{n}$ in Eq. (32) are not independent due to the triangular inequality. In our calculations, the sum over $\kappa_{n}$ is considered as the primary one. As a result, the many-potential contribution is represented by the sum of the often poorly converging PW-expansion series in $\kappa_{n}$. To calculate it, one has to resort to the procedure of extrapolation in $k=|\kappa_{n}|$. In what follows, we for brevity will use $\kappa$ instead of $\kappa_{n}$, if this does not lead to misunderstandings.
II.4 Acceleration schemes
The convergence of the PW-expansion series could be improved using different acceleration schemes. Their general idea is to subtract the term $\Delta\varepsilon_{a,x}^{\text{substr.}}$, which contains the slowly converging part of the PW expansion of the many-potential contribution. The index ‘‘$x$’’ indicates that the subtracted term should be calculated exactly in the same manner as Delta $\varepsilon_{a}^{\text{Mp}}$, that is in coordinate space using the PW expansion. The same term, evaluated separately but not necessarily within the same approach, so we denote it as $\Delta\varepsilon_{a}^{\text{substr.}}$ without the additional index ‘‘$x$’’, should then be added back. This idea can be illustrated by the formula:
$\Delta\varepsilon_{a}^{\text{Mp}}\rightarrow(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{\text{substr.}})+\Delta\varepsilon_{a}^{\text{substr.}}.$ (33)
It is implied that the PW-expansion series for the two contributions in the brackets in Eq. (33) are subtracted term by term and only then extrapolated. The resulting difference is expected to demonstrate a better convergence than the initial series for the many-potential contribution. Therefore, If one can calculate $\Delta\varepsilon_{a}^{\text{substr.}}$ with high accuracy, it will determine the overall improvement in the calculations of the SE correction.
Within the two-potential acceleration scheme (see the Introduction section), the subtracted term is the two-potential contribution, $\Delta\varepsilon_{a,x}^{\text{substr.}}=\Delta\varepsilon_{a,x}^{2\text{p}}$. When adding this term, one must evaluate it with high precision in order to benefit from applying the scheme. A natural idea is to perform such calculations in momentum space without using the PW expansions. However, this problem is complicated and leads to complex multidimensional integrals, see, e.g., Ref. Yer2010, where a similar approach was employed in the point-nucleus case for the SE correction to the hyperfine splitting. For this reason, an alternative was proposed in Refs. PhysRevLett.98.173004; Art_trick, where the difference between the many- and two-potential terms was addressed using the FBS method, while the two-potential term itself was also considered in coordinate space but employing the Green’s functions method and truncating the corresponding PW summations at higher values of angular momenta. We note that in this case, it is not necessary to solve the system of differential equations numerically for an arbitrary potential, since the PW expansion of the free-electron propagator can be expressed in terms of the spherical Bessel functions and the desired two-potential term can be obtained by combining three such propagators. In the present work, we study this two-potential scheme but for simplicity the difference between the many- and two-potential terms is also calculated based on the Green’s function approach. We refer to this term as the three-plus-potential contribution.
The Sapirstein-Cheng (SC) scheme sap_orig is based on the idea that the dominant contribution from the free-electron propagators comes from the region where both spatial arguments are close to each other. Therefore, one can obtain an approximation to the two-potential term by ‘‘transposing’’ the potentials from the inner electron line to the vertices, where the photon propagator is attached, see Fig. 3. We refer to this approximation as the quasi-two-potential term and denote it as $\Delta\varepsilon_{a,x}^{\text{substr.}}=\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$.
Figure 3: Diagrammatic representation of the quasi-two-potential term separation. The last term on the right side is calculated in momentum space without using any PW expansion.
A solid advantage of this scheme is the fact that it is fairly easy to calculate the quasi-two-potential contribution with high accuracy in momentum space. We denote the result as $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ and use it as $\Delta\varepsilon_{a}^{\text{substr.}}$. Let us discuss the SC scheme in more detail.
In the coordinate-space part of the calculations, we should construct the PW expansion
$\widetilde{G}^{(3+)}=G^{(2+)}-\sum_{f}\frac{|Vf\rangle\langle fV|}{(\varepsilon_{a}-\omega-\varepsilon_{f}^{-})^{3}},$ (34)
which provides an approximation for the contribution involving three or more interactions with the binding potential. Here $|Vf\rangle=V|f\rangle$. Having this expansion, one can directly evaluate the expression $\Delta\tilde{\varepsilon}_{a,x}^{3+}=\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$, which we will refer as the quasi-three-plus-potential contribution. Then, it remains to calculate the $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ in momentum space. The potentials that have been transposed into vertices can be formally taken into account by multiplying the wave function of the external state $|a\rangle$ by them. An issue arises of calculating the Fourier transform of such products. Since $V$ is assumed to be spherically symmetric, this can be done by a straightforward generalization of Eqs. (13)-(15). We represent the result in the following form:
$\phi_{a}(\mathbf{p})=\int d^{3}{\mathbf{r}}e^{-i({\mathbf{p\cdot r}})}V(r)\psi_{a}({\mathbf{r}})=i^{-l_{a}}{\begin{pmatrix}\widetilde{t}_{a}(p)\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{p}})\\ \widetilde{s}_{a}(p)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{p}})\end{pmatrix}}$, (35)
$\widetilde{t}_{a}(p)=4\pi\int_{0}^{\infty}drr^{2}j_{l_{a}}(pr)V(r)g_{a}(r),$ (36)
$\widetilde{s}_{a}(p)=-4\pi\frac{\kappa_{a}}{|\kappa_{a}|}\int_{0}^{\infty}drr^{2}j_{2j_{a}-l_{a}}(pr)V(r)f_{a}(r).$ (37)
Using the identity
$\int d\mathbf{y}d\mathbf{z}G^{(0)}(\omega,\mathbf{x_{1}},\mathbf{y})G^{(0)}(\omega,\mathbf{y},\mathbf{z})G^{(0)}(\omega,\mathbf{z},\mathbf{x_{2}})=\frac{1}{2}\frac{\partial^{2}}{\partial\omega^{2}}G^{(0)}(\omega,\mathbf{x_{1}},\mathbf{x_{2}}),$ (38)
which readily follows from the spectral representation for $G^{(0)}$ and the orthogonality of the Dirac basis, one obtains
$\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}=\frac{1}{2}\int\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\bar{\phi}_{a}(\mathbf{p})\frac{\partial^{2}\Sigma_{R}^{(0)}(\mathrm{p})}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\phi_{a}(\mathbf{p}),$ (39)
where $\mathrm{p}=(p_{0},\mathbf{p})$. The explicit form of the derivatives of the free-electron self-energy operator $\Sigma_{R}^{(0)}$ in both gauges is given in Appendix D. After performing the angular integration, which is carried out similarly to that in the zero-potential contribution, one can obtain:
$\begin{gathered}\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}=\frac{\alpha}{8\pi}\int\frac{dpp^{2}}{(2\pi)^{3}}\left[N_{1}\left(\widetilde{t}_{a}^{2}-\widetilde{s}_{a}^{2}\right)+N_{2}\left(\widetilde{t}_{a}^{2}+\widetilde{s}_{a}^{2}\right)+2pN_{3}\widetilde{t}_{a}\widetilde{s}_{a}\right]\end{gathered}$ (40)
where $N_{1}$, $N_{2}$, and $N_{3}$ are defined in Appendix D, $\widetilde{t}_{a}=\widetilde{t}_{a}(p)$ and $\widetilde{s}_{a}=\widetilde{s}_{a}(p)$.
III Computational details
The calculations are carried out by means of two approaches: the FBS method within the DKB basis and the GF method. The methods differ only in the treatment of the terms evaluated in coordinate space. The momentum-space calculations are the same in both approaches.
III.1 Momentum-space calculations
The contributions evaluated in momentum space include the zero- and one-potential terms resulting from the renormalization procedure, as well as the quasi-two-potential terms corresponding to the SC convergence-acceleration scheme. All the integrals in the zero-, one-, and quasi-two-potential terms are calculated using the Gauss-Legendre quadratures. Compared to the Feynman gauge, the zero-potential contribution in the Coulomb gauge contains an additional integration over the Feynman parameter $x$, see Eqs. (17), (56), (57). The coefficient $F_{2}$ in Eq. (57) must be calculated as a principal-value integral. From a numerical point of view, it is necessary to isolate the singularity. We do that by adding and subtracting the same integrand but with the numerator evaluated at $x=a$:
$\int dx\frac{f(x)}{x-a+i0}=\int dx\frac{f(x)-f(a)}{x-a+i0}+\int dx\frac{f(a)}{x-a+i0}.$ (41)
The first term in the right-hand side of Eq. (41) becomes regular. To avoid numerical problems at $x=a$, the exact expression can be replaced with its Taylor series in a small vicinity of this point. The second term in Eq. (41) is calculated analytically. The domain of integration over $p$ is divided into two parts, $p\leq p_{0}$ and $p\geq p_{0}$, where $p_{0}=\alpha Z$. The integration over the second domain can be carried out using a change of variables of the form $t=\alpha Z/p$, as a result of which the ray $\alpha Z,\infty)$, is mapped onto the interval $[1,0]$. Similar methods can be used to calculate $\Delta\tilde{\varepsilon}_{a}^{2\text{p}}$ in momentum space. To avoid possible numerical problems due to large cancellations, one should use the Taylor-series expansion for the corresponding integrands.
The evaluation of the one-potential contributions involves the integrations over the variables $p$, $p^{\prime}$, and $z$ and a number of Feynman parameters. When integrating over the Feynman parameters in both gauges, to avoid possible numerical errors, the Taylor series for the integrands are employed. In the Coulomb-gauge case we use the results of Ref. OP_CG, but expand all the coefficients $F_{1}$ - $F_{21}$ at ‘‘suspicious’’ points. The expression (20) has an integrable singularity in $q=0$ at $p=p^{\prime}$ and $z=1$. Therefore it is necessary to properly choose integration nodes for the variable $z$ to ensure the integral convergence. For this goal, we use the transformation $z=1-2t^{2}$. To calculate the integrals over the variables $p$ and $p^{\prime}$, we change the variables according to $P=(p+p^{\prime})/\sqrt{2}$ and $P^{\prime}=(p-p^{\prime})/\sqrt{2}$. The integral over $P$ is calculated similar to the integral over $p$ for the zero-potential contribution, but the separation point $p_{0}=2\alpha Z$ is chosen instead. The integral over $P^{\prime}$ is taken within the finite interval [P,P][-P,P].
To estimate an uncertainty of the momentum-space calculations, the number of quadrature nodes and position of the separation points $p_{0}$ is varied.
III.2 Coordinate-space calculations
The evaluation of all coordinate-space contributions, many-potential, two-potential, and (quasi-) three-plus-potential, has many common points. Let us illustrate them with an example of the many-potential contribution. Its evaluation essentially comes down to the calculation of the three integrals: over the radial variables $r_{1}$ and $r_{2}$, see Eqs. (71) and (72), and over the energy parameter $\omega$. Taking into account the exponential decrease of bound-electron wave functions at large distances, all radial integrations are performed over the finite interval $0,R_{max}]$. The value of the parameter $R_{max}$ is chosen so that the results are independent of it within the desired accuracy. All the radial integrations are carried out using the Gauss–Legendre quadratures. However, the specific details differ slightly in the DKB and GF approaches.
In the DKB approach, the domain $0,R_{max}]$ is divided into a finite number of intervals by the nodes used to construct the B splines. The integration method for the double integral over $r_{1}$ and $r_{2}$ depends on whether these variables belong to the same interval or not. If not, the integrations are carried out independently using the same quadrature for both variables. However, when $r_{1}$ and $r_{2}$ fall into the same interval, a more delicate integration is employed to take into account the fact that the integrand depends on $r_{<}=\min\{r_{1},r_{2}\}$ and $r_{>}=\max\{r_{1},r_{2}\}$. Namely, for some fixed value of $r_{1}$ the corresponding interval is divided into two subitervals lying to the left and to the right from $r_{1}$. The integration over $r_{2}$ is performed by using the Gauss-Legendre quadratures for both subintervals. Note that within the DKB approach, integration over radial variables turns out to be ‘‘tied’’ to the size of the basis $N$. As the basis-set size grows, the number of the nodes also increases, that leads to the improvement of the integration accuracy. Thus, the single parameter $N$ affects both the accuracy of the Green’s function representation and the integration accuracy.
In the GF approach, the radial integrations become more complicated due to the aforementioned discontinuities of the electron Green’s functions. The integrations over $x$ and $y$ in Eq. ((25)) cannot be performed independently of the integrations over $r_{1}$ and $r_{2}$. To evaluate the radial integrals, we employ a variation of the method described, e.g., in Ref. sym12050800. Namely, a hierarchy of the radial integration grids is prepared, with each subsequent grid being finer and "lying" inside the previous one. Then, all the radial integrations are arranged in the following order: $r_{1}$, $r_{2}$, $x$, and $y$. Therefore, the finest grid corresponds to the variable $y$. The accuracy of the GF approach is determined mainly by the accuracy of the radial integrations.
$\text{Re}(\omega)$$\text{Im}(\omega)$$\varepsilon_{a}-\varepsilon_{1s}$
Figure 4: Deformed contour (blue line) in the complex plane $\omega$. Singularities and branching cuts of the electron Green’s function (black lines and black dots) and photon propagator (black dashed lines) are shown.
The photon and electron propagators have a complicated analytical structure if considered as functions of the energy parameter $\omega$ in the complex plane. They are two-valued functions and are defined on the corresponding Riemann surfaces with two sheets (for each propagator). Fig. 4 shows the complex plane of the integration variable $\omega$, the branch points with the cuts starting from them, and the poles of the electron propagator corresponding to the bound states. In the expression for the SE correction (3), the integration contour goes along the real axis, but it can be bended to ‘‘simplify’’ the calculations. Because of the zero photon mass, the contour is squeezed between two infinitely close branching points. The introduction of a finite mass would allow to push them apart, which can be easily demonstrated in the Feynman gauge. Therefore, the contour is always obliged to pass through $\omega=0$. In principle, there is no other restrictions on the contour. Thus, it can be changed as needed, respecting the analytical structure of the integrand.
In this work, the approach proposed in Ref. base_1999 is used. Namely, we employ the transformation that results in the contour shown in Fig. 4. The original contour along the real axis is rotated and deformed. The rotation is done to avoid fast oscillations of the integrand encountered on the real axis: on the vertical parts of the deformed contour the integrand decays exponentially. The deformation of the contour near zero and its shift to the right from the imaginary axis is done to avoid going around the poles of the electron Green’s function, which correspond to the bound states with binding energies no less than that of the state under consideration. The integral over $\omega$ is evaluated using the Simpson quadratures. The number of integration points is increased until the required accuracy is achieved.
III.3 Extrapolation of the PW-expansion sums
We obtain a coordinate-space contribution as a series in $\kappa$, which should inevitably be truncated at some $|\kappa_{\text{max}}|$. To estimate the remainder of the series, we must resort to an extrapolation procedure. The extrapolation methods discussed below are applied to all treated coordinate-space contributions.
Here and in what follows, when discussing partial sums, we will denote by $k>0$ the sum of terms of the PW-expansion with $|\kappa|=k$. When using the DKB method, the individual contributions strongly depend on the basis-set size $N$, therefore, an extrapolation over this parameter is also required. In this case, the most stable results are obtained by first extrapolating with respect to the basis size. Let us denote as $S_{k}(N)$ the partial sum obtained for the given basis-set size $N$. For each $k$, the extrapolation over $N$ is performed using a non-linear least-squares method using an ansatz of the form:
$S_{k}(N)=A_{k}\left(\frac{1}{N}\right)^{B_{k}}+S_{k},$ (42)
where $S_{k}$ is the extrapolated to the infinite basis set partial sum, and $A_{k}$ and $B_{k}$ are numerical coefficients. In the GF approach, the sums $S_{k}$ are obtained directly from the calculations, and this extrapolation is not required.
To extrapolate over $k$ in both DKB and GF approaches, we implement the following anzatz for the partial sums $S_{k}$:
$\tilde{S}(k)=S_{\infty}+\frac{C_{2}}{k^{2}}+\frac{C_{3}}{k^{3}}+...+\frac{C_{m}}{k^{m}}\penalty 10000\ ,$ (43)
where the parameter $S_{\infty}$ and coefficients $C_{i}$ for $i=2\ldots m$ (with $m$ being typically about $5$) are obtained by a statistical method. Namely, we select a random ‘‘kappa’’ subset $\{k_{i}\}_{i=1}^{m}$ a certain number $N_{\text{iter}}=100-1000$ of times. An additional condition is imposed that there are no consecutive $k$ in the chosen subset. We then construct and solve the system of $m$ linear equations of the form
$\tilde{S}(k_{i})=S_{k_{i}},\;i=1\dots m$ (44)
to determine $m$ variables $S_{\infty}$, $C_{2}$, … $C_{m}$. This procedure is repeated $N_{\text{iter}}$ times, resulting in a set of approximations for $S_{\infty}$. Evaluating the mean value $\overline{S}_{\infty}$ and variance $\sigma_{k}$ one obtains the desired extrapolated sum and an estimation for its uncertainty.
Within DKB approach, the uncertainties of extrapolation in k, $\sigma_{k}$, and extrapolation to an infinite basis set, $\sigma_{n}$, are summed quadratically
$\sigma_{\text{tot}}=\sqrt{\sigma_{\kappa}^{2}+\sigma_{N}^{2}}.$ (45)
Example of the use of the two-step extrapolation procedure can be found in the ‘‘results’’ section below.
IV Results
We conduct two sets of SE correction calculations. First we consider the $1S_{1/2}$ state in three hydrogen-like ions: argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). We adopt the results from the work hen as a benchmark for the comparison. For this reason, the values of nuclear radii as well as the values of fundamental constants, e.g. the fine-structure constant $\alpha$ and the Hartree energy in eV, are taken to be the same as there. The obtained data are used to analyze the convergence of the many-potential contribution.
We then reproduce the results from the work se_yerokh, the rms radii and models of the nuclei are chosen to coincide with the corresponding work. In that part, the SE correction is presented in terms of the function $F(\alpha Z)$, defined as:
$\Delta\varepsilon_{a}=\frac{\alpha}{\pi}\frac{(\alpha Z)^{4}}{n^{3}}F(\alpha Z)mc^{2}.$ (46)
We use the numerical value for $\alpha$ from the latest CODATA (2018) codata_2018. The data obtained in the second part of the work are used to analyze acceleration-convergence schemes.
IV.1 Many-potential term: convergence analysis
Let us start with the example of the use of the two-step extrapolation procedure for the ground state of hydrogen-like xenon ($Z=54$), see Fig. 5. The many-potential term is considered, the results for both gauges are presented. At the first stage, the extrapolation with respect to the size of the basis $N$ is performed. For convenience, the extrapolation curves $k=10$, $15$ and $20$ are shown. The corresponding extrapolated sums $S_{k}$ are depicted as stars. The other extrapolated partial sums are depicted with dashes as well. At the second stage, the approximation over $k$ is performed. The final value ($\Delta\varepsilon_{a}^{mp}=S_{\infty}$, black) is obtained. Note that the GF method yields results that are indistinguishable at this scale from those obtained by the DKB method.
Based on our experience, there is no significant difference in extrapolation by basis size across two gauges. This can be seen, e.g., in the Fig. 5, where the many-potential term differ greatly (by one and a half orders of magnitude) and the rate of convergence in the basis-set size is approximately the same.
Figure 5: The extrapolation of the many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like xenon ($Z=54$) obtained within the framework of the DKB approach in Feynman (a) and Coulomb (b) gauges.
Figure 6: The many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}(k)$: (a) normalized to the full self-energy correction $\Delta\varepsilon_{a}$ and (b) additionally multiplied by $k^{3}$. Note the log-log scale.
Let us examine the $k$-dependence of the many-potential contributions in more detail and perform the comparative analysis of the $k$ – convergence of the many-potential contribution in two gauges. The zero-, one-, and many-potential contributions to the SE correction in two gauges for $Z=18$ and $Z=92$ as a series in $k$ are presented in Tab. 1. There, at the bottom of the table, the results obtained by applying the extrapolation procedure to the corresponding PW-expansion series are shown. Note that the notation $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ is introduced. The latter is the contribution to $\Delta\varepsilon_{a}^{\text{Mp}}$ corresponding to the given $k$ (thus, the partial sum $S_{k}$ for many-potential term is $S_{k}=\sum_{k^{\prime}=1}^{k}\Delta\varepsilon_{a}^{\text{Mp}}(k^{\prime})$).
In Fig. 6 (a) the individual $k$ terms of the PW expansion of the many-potential contribution from Tab. 1 normalized to the full self-energy are plotted. From both the table and the graph it can be noticed that for all $Z$ both the many-potential contribution and the first PW contributions are significantly smaller in the Coulomb gauge (in the case of the $Z=18$ argon – by two orders of magnitude). This difference, however, becomes much smaller in the large-$k$ area, although the contributions are still larger in the Feynman gauge.
Multiplying by the $k^{3}$ factor allows a much closer comparison of the decay rate of contributions, see Fig. 6 (b). From our study it follows that the individual contributions are decreasing as $\Delta\varepsilon_{a}^{\text{Mp}}(k)\sim\frac{1}{k^{3}}$. That justifies that the smallest power used in $k$ extrapolation in (43) is $\frac{1}{k^{2}}$, since $\sum_{k^{\prime}=1}^{k}\frac{1}{k^{\prime 3}}\sim\frac{1}{k^{2}}$. At the same time, convergence to $\frac{1}{k^{3}}$ is clearly slower in the Coulomb gauge. An indirect interpretation is that the coefficients at the higher powers in the expansion (43) are larger in the Coulomb gauge case.
The latter circumstance apparently balances out the smallness of the many-potential contribution itself in the Coulomb gauge. In practice, given the same resource costs (number of basis functions, number of integration nodes for radial variables, etc.), Coulomb and Feynman gauges are equally suitable for estimating the many-potential term of the SE correction within a wide range of nuclear charges $Z$.
IV.2 Notes on the acceleration-convergence schemes
Now let us turn to the comparative analysis of the SC and two-potential schemes used to accelerate the convergence of the many-potential term in two gauges. In Fig. 8 (a) the ratio of the (quasi-) three-plus-potential to the many-potential contributions is plotted. In the small-$Z$ region the two-potential scheme in Coulomb gauge yields the smallest three-plus-potential contribution to be extrapolated. However, the accuracy there is limited by the remainder – the two-potential contribution itself. The non-monotonicity of the Coulomb gauge SC graph is explained by the fact that the quasi-three-plus potential contribution passes through zero and changes sign.
Figure 7: Comparison of two schemes in two gauges, extrapolated sum of the series (a) and leading power ($\sim\frac{1}{k^{2}}$) coefficient $C_{2}^{3+}$ (b) for (quasi-) three-plus-potential terms. Note the normalization to the many-potential contribution.
Let us discuss the extrapolation coefficients $C_{i}$ from Eq. (43). From our study it follows that their relative error grows quite rapidly with increasing $i$. Therefore we consider the first expansion coefficient $C_{2}$, plotted in Fig. 7 (b). This graph partially explains the convergence acceleration mechanism. In the area of small $Z$, the coefficient $C_{2}$ decreases by one or two orders of magnitude; in the region of large $Z$, it also decreases, but less effectively. Within the same gauge regardless of the tricks $C_{2}$ decreases almost identically illustrating the viability of the SC approximation.
Figure 8: Comparison of two schemes in two gauges, normalized PW expansion contributions. Data for $Z=18$ argon (a) and for $Z=92$ uranium (b) is shown.
We cannot conclude that the extraction of two-potential contributions reduces the minimum power in the expansion (43) to 3; we argue that the opposite situation occurs. The last statement has been verified empirically – the extrapolation procedure gave incorrect results. It can also be reasoned by the contributions graphs scaled by $k^{3}$, where the (quasi-)three-plus-potential contribution come out to a constant in all cases, see Fig. 8 (b). This graph also demonstrates that any acceleration scheme in the Coulomb gauge performs better than in the Feynman gauge, as the convergence to the asymptotics is much faster.
From the above analysis, it is difficult to conclude which specific acceleration scheme is better. However the SC scheme has a significant advantage: the ability to obtain a two-potential contribution without a PW expansion. In practice almost always the combination of the Coulomb gauge and the SC scheme allows one to obtain the most accurate self-energy correction for all $Z$.
IV.3 Data: summary
The calculated SE corrections with the parameters from the Ref. hen are summarized in Tab. 2. In this table the zero-, one-, and most accurate many-potential contributions for three nuclei are listed in two gauges. One can see that the total uncertainty of the SE correction is mainly determined by the uncertainty of the many-potential contribution. Even taking into account the acceleration schemes. We present separately the many-potential, two-potential, and quasi-two-potential contributions in Tab. 3. Using this table one can compare the accuracy of the many-potential contribution obtained. Overall, it is clear that the acceleration tricks do reduce the extrapolation error and allow to obtain the many-potential contribution more accurately. In addition, there is a small but noticeable difference between our data and the data from Ref. hen.
Results with the parameters from the Ref. se_yerokh are presented in the similar manner in Tab. 5. An attractive feature of the Coulomb gauge can be seen from that table – the absence of the large reductions between the momentum (zero-, one-potential) contributions in the low-$Z$ region. It is clearly visible from the Fig. 9 (a) (compare the green and yellow curves). Various coordinate contributions are given in a separate table 6. In some cases we are able to reduce the numerical error compared with the work se_yerokh and the results obtained are in reasonable agreement with the corresponding work. Calculations for excited states using the Green’s function method are also performed, see Tab. LABEL:tab:we_yerokh_excited_states.
V Conclusion
The self-energy correction for various states of hydrogen-like ions is calculated. Two gauges are considered, acceleration tricks are discussed and convenient formulas for the corresponding calculations are collected.
Overall, the total numerical uncertainty of the SE correction is mainly determined by the numerical uncertainty of the many-potential contribution. The contribution itself is significantly smaller in the Coulomb gauge than in the Feynman one. It could be seen, firstly, by the blue curves in the Fig. 9 (a), and secondly, in Fig. 9 (b), where the sum of the zero- and one-potential contributions is shown separately. Despite the smallness of the many-potential contribution, it does not allow obtaining a full self-energy correction with greater accuracy if no acceleration schemes are involved. An important feature of the Coulomb gauge SE calculations is the absence of the large reductions between the momentum (zero-, one-potential) contributions in the low-$Z$ region.
The acceleration schemes do improve the accuracy of the calculations performed. Using either the two-potential scheme or the SC scheme reduces the extrapolation error and allows one to obtain up to two new significant figures in the many-potential contribution. The combination of Coulomb gauge and SC scheme is most suitable (among those considered) for calculating the self-energy correction in hydrogen-like ions, especially when considering light nuclei. The next goal is to consider the Fried-Yennie gauge; we believe that the analysis performed will simplify the calculation of a wide range of more complex radiative corrections with the self-energy loop.
VI Acknowledgments
This work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics ‘‘BASIS’’.
Table 1: Contributions to the self-energy correction for the ground state of hydrogen-like argon and uranium. obtained in the Feynman (F) and Coulomb (C) gauges (in eV). The individual terms of the partial-wave expansion for the many-potential contribution as well as the results of applying the extrapolation procedure are shown. The nuclear-charge distribution is described by the homogeneously-charged-sphere model with the root-mean-square radii taken from hen.
$Z=18$ $Z=92$
F C F C
$\Delta\varepsilon_{a}^{0\text{p}}$ 67.924836-67.924836 $1.341667$ 516.318629-516.318629 $210.068167$
$\Delta\varepsilon_{a}^{1\text{p}}$ $49.511442$ $0.054770$ $472.000467$ $213.738955$
$\Delta\varepsilon_{a}^{\text{Mp}}(k),$ $k=1$ $18.954139$ 0.098735-0.098735 $392.056752$ 54.030900-54.030900
22 $0.447224$ 0.044326-0.044326 $2.922927$ 11.168699-11.168699
33 $0.115257$ 0.011632-0.011632 $2.001276$ 1.902586-1.902586
44 $0.046658$ 0.006453-0.006453 $0.915478$ 0.685643-0.685643
55 $0.023375$ 0.004153-0.004153 $0.478148$ 0.330469-0.330469
66 $0.013267$ 0.002872-0.002872 $0.278392$ 0.186004-0.186004
77 $0.008186$ 0.002080-0.002080 $0.175634$ 0.115450-0.115450
88 $0.005368$ 0.001558-0.001558 $0.117699$ 0.076706-0.076706
99 $0.003688$ 0.001199-0.001199 $0.082646$ 0.053600-0.053600
$10$ $0.002630$ 0.000943-0.000943 $0.060225$ 0.038946-0.038946
$11$ $0.001933$ 0.000754-0.000754 $0.045228$ 0.029195-0.029195
$12$ $0.001458$ 0.000613-0.000613 $0.034822$ 0.022452-0.022452
$13$ $0.001124$ 0.000505-0.000505 $0.027378$ 0.017638-0.017638
$14$ $0.000882$ 0.000420-0.000420 $0.021913$ 0.014110-0.014110
$15$ $0.000704$ 0.000353-0.000353 $0.017811$ 0.011464-0.011464
$16$ $0.000570$ 0.000300-0.000300 $0.014672$ 0.009441-0.009441
$17$ $0.000468$ 0.000257-0.000257 $0.012229$ 0.007868-0.007868
$18$ $0.000388$ 0.000221-0.000221 $0.010300$ 0.006626-0.006626
1919 $0.000325$ 0.000192-0.000192 $0.008756$ 0.005632-0.005632
2020 $0.000275$ 0.000168-0.000168 $0.007506$ 0.004827-0.004827
2121 $0.000234$ 0.000147-0.000147 $0.006483$ 0.004169-0.004169
$22$ $0.000201$ 0.000130-0.000130 $0.005638$ 0.003625-0.003625
$23$ $0.000174$ 0.000115-0.000115 $0.004933$ 0.003172-0.003172
$24$ $0.000152$ 0.000103-0.000103 $0.004341$ 0.002792-0.002792
$\sum_{k=1}^{24}$ $19.628680$ 0.178229-0.178229 $399.311187$ 68.732014-68.732014
$\sum_{k=25}^{\infty}$ [extr.] $0.00161(1)$ 0.00130(1)-0.00130(1) $0.049954(11)$ 0.0321(2)-0.0321(2)
$\Delta\varepsilon_{a}$ $1.21689(1)$ $1.21690(1)$ $355.042979(11)$ $355.0430(2)$
hen $1.21690(1)$ $1.216901(3)$ $355.0432(2)$ $355.0430(1)$
Table 2: Individual contributions to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like ions in the Feynman (F) and Coulomb (C) gauges (in eV). The nuclear-charge distribution is described by the homogeneously-charged-sphere model with the root-mean-square radii taken from hen.
$Z$ gauge $\Delta\varepsilon_{a}^{0\text{p}}$ $\Delta\varepsilon_{a}^{1\text{p}}$ $\Delta\varepsilon_{a}^{\text{Mp}}$ $\Delta\varepsilon_{a}$
18 C DKB 1.341 667 21.341\,667\,2 0.054 770 20.054\,770\,2 0.179 52(2)-0.179\;52(2) 1.216 91(2)1.216\;91(2)
GF 0.179 530 8(2)-0.179\;530\;8(2) 1.216 906 6(2)1.216\;906\;6(2)
hen 1.341 668 068(1)1.341\;668\;068(1) 0.054 770 997(7)0.054\;770\;997(7) 0.179 538(3)-0.179\;538(3) 1.216 901(3)1.216\;901(3)
F DKB 67.924 836 7-67.924\,836\,7 49.511 442 349.511\,442\,3 19.630 2(1)19.630\;2(1) 1.216 8(1)1.216\;8(1)
GF 19.630 301 2(4)19.630\;301\;2(4) 1.216 906 8(4)1.216\;906\;8(4)
hen 67.924 837 74(5)-67.924\;837\;74(5) 49.511 443 05(6)49.511\;443\;05(6) 19.630 296(10)19.630\;296(10) 1.216 90(1)1.216\;90(1)
54 C DKB 43.590 61043.590\,610 17.387 95817.387\,958 9.981 3(2)-9.981\;3(2) 50.997 2(2)50.997\;2(2)
GF 9.981 311(3)-9.981\;311(3) 50.997 257(2)50.997\;257(2)
hen 43.590 621 48(6)43.590\;621\;48(6) 17.387 986 7(3)17.387\;986\;7(3) 9.981 343(16)-9.981\;343(16) 50.997 27(2)50.997\;27(2)
F DKB 285.092 638-285.092\,638 190.356 023190.356\,023 145.733 8(7)145.733\;8(7) 50.997 1(7)50.997\;1(7)
GF 145.733 872(3)145.733\;872(3) 50.997 257(3)50.997\;257(3)
hen 285.092 638 6(1)-285.092\;638\;6(1) 190.356 052 0(3)190.356\;052\;0(3) 145.733 90(8)145.733\;90(8) 50.997 31(8)50.997\;31(8)
92 C DKB 210.068 167210.068\,167 213.738 955213.738\,955 68.764 8(8)-68.764\;8(8) 355.042 3(8)355.042\;3(8)
GF 68.764 144(5)-68.764\;144(5) 355.042 978(5)355.042\;978(5)
hen 210.068 220 5(7)210.068\;220\;5(7) 213.739 094(4)213.739\;094(4) 68.764 3(1)-68.764\;3(1) 355.043 0(1)355.043\;0(1)
F DKB 516.318 629-516.318\,629 472.000 467472.000\,467 $399.364(7)$ $355.045(7)$
GF 399.361 141(11)399.361\;141(11) 355.042 979(11)355.042\;979(11)
hen 516.318 598(4)-516.318\;598(4) 472.000 597(6)472.000\;597(6) 399.361 2(2)399.361\;2(2) 355.043 2(2)355.043\;2(2)
Table 3: Many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like ions (in eV), calculated in three ways: directly, (using the separation of the two-potential contribution), and (using the Sapirstein-Cheng (SC) approach). The nuclear-charge distribution is described by the homogeneously-charged-sphere model with the root-mean-square radii taken from hen.
direct two-pot. scheme SC scheme
$Z$ gauge $\Delta\varepsilon^{\text{Mp}}_{a}$ $\Delta\varepsilon^{2\text{p}}_{a}$ $\Delta\varepsilon^{3+}_{a}$ $\Delta\varepsilon^{\text{Mp}}_{a}$ $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ $\Delta\tilde{\varepsilon}_{a}^{3+}$ $\Delta\varepsilon_{a}^{\text{Mp}}$
18 C DKB 0.179 4(1)-0.179\;4(1) 0.218 341 8-0.218\,341\,8 0.038 82(2)0.038\;82(2) 0.179 52(2)-0.179\;52(2)
GF 0.179 53(1)-0.179\;53(1) 0.155 550 67(9)-0.155\;550\;67(9) 0.023 979 98(12)-0.023\;979\;98(12) 0.179 530 8(2)-0.179\;530\;8(2) 0.038 811 4(2)0.038\;811\;4(2) 0.179 530 6(2)-0.179\;530\;6(2)
F DKB $19.630(1)$ 8.722 252 38.722\,252\,3 10.908 0(1)10.908\;0(1) 19.630 2(1)19.630\;2(1)
GF 19.630 29(1)19.630\;29(1) 8.960 544 04(15)8.960\;544\;04(15) 10.669 756 5(6)10.669\;756\;5(6) 19.630 300 6(7)19.630\;300\;6(7) 10.908 048 8(4)10.908\;048\;8(4) 19.630 301 2(4)19.630\;301\;2(4)
54 C DKB 9.982(1)-9.982(1) 7.403 692-7.403\,692 2.577 7(2)-2.577\;7(2) 9.981 3(2)-9.981\;3(2)
GF 9.981 32(2)-9.981\;32(2) 6.727 056 62(3)-6.727\;056\;62(3) 3.254 253(2)-3.254\;253(2) 9.981 31(2)-9.981\;31(2) 2.577 618(3)-2.577\;618(3) 9.981 311(3)-9.981\;311(3)
F DKB $145.731(2)$ 54.420 12054.420\,120 91.313 7(7)91.313\;7(7) 145.733 8(7)145.733\;8(7)
GF 145.733 84(4)145.733\;84(4) 55.334 247 46(3)55.334\;247\;46(3) 90.399 623(7)90.399\;623(7) 145.733 871(7)145.733\;871(7) 91.313 751(3)91.313\;751(3) 145.733 872(3)145.733\;872(3)
92 C DKB 68.764(3)-68.764(3) 23.349 279-23.349\,279 45.415 6(8)-45.415\;6(8) 68.764 8(8)-68.764\;8(8)
GF 68.764 1(2)-68.764\;1(2) 31.829 260 86(19)-31.829\;260\;86(19) 36.934 883(6)-36.934\;883(6) 68.764 144(6)-68.764\;144(6) 45.414 865(5)-45.414\;865(5) 68.764 144(5)-68.764\;144(5)
F DKB $399.36(1)$ 149.189 230149.189\,230 $250.175(7)$ $399.364(7)$
GF 399.361 141(11)399.361\;141(11) 144.424 324 42(12)144.424\;324\;42(12) 254.936 814(17)254.936\;814(17) 399.361 138(17)399.361\;138(17) 250.171 912(15)250.171\;912(15) 399.361 142(15)399.361\;142(15)
Table 4: Many-potential term and the self-energy correction for the ground state of hydrogen-like argon. Comparison of two acceleration tricks and two gauges, Feynman (F) and Coulomb (C) (in eV). The individual terms of the partial-wave expansion for the many-potential contribution as well as the results of applying the extrapolation procedure are shown. The nuclear-charge distribution is described by the homogeneously-charged-sphere model with the root-mean-square radii taken from hen.
F C
$\Delta{\varepsilon}_{a}^{\text{Mp}}$ $\Delta{\varepsilon}_{a}^{3+}$ $\Delta\tilde{\varepsilon}_{a}^{3+}$ $\Delta{\varepsilon}_{a}^{\text{Mp}}$ $\Delta{\varepsilon}_{a}^{3+}$ $\Delta\tilde{\varepsilon}_{a}^{3+}$
substr. $8.96054404(15)$ $8.7222523$ 0.15555067(9)-0.15555067(9) 0.2183418-0.2183418
$|\kappa|=1$ $18.954139$ $10.6306576$ $10.8509659$ 0.098735-0.098735 0.01593568-0.01593568 $0.0408756$
22 $0.447224$ $0.0333412$ $0.0491332$ 0.044326-0.044326 0.00662752-0.00662752 0.0016441-0.0016441
33 $0.115257$ $0.0042399$ $0.0059313$ 0.011632-0.011632 0.00061625-0.00061625 $0.0000173$
44 $0.046658$ $0.0010336$ $0.0013902$ 0.006453-0.006453 0.00026991-0.00026991 0.0000776-0.0000776
55 $0.023375$ $0.0003404$ $0.0004394$ 0.004153-0.004153 0.00015126-0.00015126 0.0000734-0.0000734
66 $0.013267$ $0.0001294$ $0.0001612$ 0.002872-0.002872 0.00009440-0.00009440 0.0000571-0.0000571
77 $0.008186$ $0.0000519$ $0.0000628$ 0.002080-0.002080 0.00006308-0.00006308 0.0000431-0.0000431
88 $0.005368$ $0.0000201$ $0.0000238$ 0.001558-0.001558 0.00004428-0.00004428 0.0000326-0.0000326
99 $0.003688$ $0.0000062$ $0.0000017$ 0.001199-0.001199 0.00003229-0.00003229 0.0000250-0.0000250
$10$ $0.002630$ 0.0000000-0.0000000 $0.0000000$ 0.000943-0.000943 0.00002426-0.00002426 0.0000195-0.0000195
$11$ $0.001933$ 0.0000028-0.0000028 0.0000030-0.0000030 0.000754-0.000754 0.00001868-0.00001868 0.0000155-0.0000155
$12$ $0.001458$ 0.0000039-0.0000039 0.0000042-0.0000042 0.000613-0.000613 0.00001469-0.00001469 0.0000124-0.0000124
$13$ $0.001124$ 0.0000042-0.0000042 0.0000046-0.0000046 0.000505-0.000505 0.00001175-0.00001175 0.0000101-0.0000101
$14$ $0.000882$ 0.0000042-0.0000042 0.0000044-0.0000044 0.000420-0.000420 0.00000955-0.00000955 0.0000084-0.0000084
$15$ $0.000704$ 0.0000039-0.0000039 0.0000042-0.0000042 0.000353-0.000353 0.00000786-0.00000786 0.0000070-0.0000070
$16$ $0.000570$ 0.0000036-0.0000036 0.0000038-0.0000038 0.000300-0.000300 0.00000655-0.00000655 0.0000058-0.0000058
$17$ $0.000468$ 0.0000033-0.0000033 0.0000034-0.0000034 0.000257-0.000257 0.00000551-0.00000551 0.0000050-0.0000050
$18$ $0.000388$ 0.0000030-0.0000030 0.0000031-0.0000031 0.000221-0.000221 0.00000468-0.00000468 0.0000042-0.0000042
1919 $0.000325$ 0.0000027-0.0000027 0.0000028-0.0000028 0.000192-0.000192 0.00000400-0.00000400 0.0000036-0.0000036
2020 $0.000275$ 0.0000024-0.0000024 0.0000025-0.0000025 0.000168-0.000168 0.00000345-0.00000345 0.0000032-0.0000032
2121 $0.000234$ 0.0000021-0.0000021 0.0000022-0.0000022 0.000147-0.000147 0.00000300-0.00000300 0.0000028-0.0000028
$22$ $0.000201$ 0.0000019-0.0000019 0.0000020-0.0000020 0.000130-0.000130 0.00000262-0.00000262 0.0000024-0.0000024
$23$ $0.000174$ 0.0000017-0.0000017 0.0000018-0.0000018 0.000115-0.000115 0.00000230-0.00000230 0.0000021-0.0000021
$24$ $0.000152$ 0.0000016-0.0000016 0.0000016-0.0000016 0.000103-0.000103 0.00000203-0.00000203 0.0000019-0.0000019
$\sum_{k=1}^{24}$ $19.628680$ $10.6697790$ $10.9080659$ 0.178229-0.178229 0.02395560-0.02395560 $0.0388361$
$\sum_{k=25}^{\infty}$ [extr.] $0.00161(1)$ 0.0000225(6)-0.0000225(6) 0.0000171(4)-0.0000171(4) 0.00130(1)-0.00130(1) 0.00002438(12)-0.00002438(12) 0.0000247(2)-0.0000247(2)
$\Delta\varepsilon_{a}^{\text{Mp}}$ $19.63029(1)$ $19.6303006(7)$ $19.6303012(4)$ 0.17953(1)-0.17953(1) 0.1795308(2)-0.1795308(2) 0.1795306(2)-0.1795306(2)
$\Delta\varepsilon_{a}$ $1.21689(1)$ $1.2169062(7)$ $1.2169068(4)$ $1.21691(1)$ $1.2169066(2)$ $1.2169068(2)$
hen $1.21690(1)$ $1.21690(1)$ $1.21690(1)$ $1.216901(3)$ $1.216901(3)$ $1.216901(3)$
Table 5: Individual contributions to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like ions in the Feynman (F) and Coulomb (C) gauges in $F(\alpha Z)$ units. For $Z=10$ spherical model was used, and for other $Z$’s Fermi model was used, nuclear radii taken from se_yerokh.
$Z$ gauge $\Delta\varepsilon_{a}^{0\text{p}}$ $\Delta\varepsilon_{a}^{1\text{p}}$ $\Delta\varepsilon_{a}^{\text{Mp}}$ $\Delta\varepsilon_{a}$
10 C DKB 5.502 181 55.502\,181\,5 0.278 283 7-0.278\,283\,7 0.569 81(5)-0.569\;81(5) 4.654 08(5)4.654\;08(5)
GF 0.569 768 8(1)-0.569\;768\;8(1) 4.654 129 0(1)4.654\;129\;0(1)
F DKB 828.249 50-828.249\,50 644.228 14644.228\,14 188.675 50(2)188.675\;50(2) 4.654 14(2)4.654\;14(2)
GF 188.675 47(1)188.675\;47(1) 4.654 11(1)4.654\;11(1)
se_yerokh 4.654 1294.654\;129
18 C DKB 3.797 152 553.797\,152\,55 0.155 009 410.155\,009\,41 0.508 08(2)-0.508\;08(2) 3.444 08(2)3.444\;08(2)
GF 0.508 102 32(1)-0.508\;102\;32(1) 3.444 059 64(1)3.444\;059\;64(1)
F DKB 192.239 149-192.239\,149 140.126 028140.126\,028 55.557 0(1)55.557\;0(1) 3.443 9(1)3.443\;9(1)
GF 55.557 179(2)55.557\;179(2) 3.444 058(2)3.444\;058(2)
se_yerokh 3.444 059(1)3.444\;059(1)
26 C DKB 2.891 040 22.891\,040\,2 0.351 512 50.351\,512\,5 0.458 82(5)-0.458\;82(5) 2.783 73(5)2.783\;73(5)
GF 0.458 787 8(4)-0.458\;787\;8(4) 2.783 764 9(4)2.783\;764\;9(4)
F DKB 74.244 868 3-74.244\,868\,3 51.657 591 351.657\,591\,3 25.370 9(1)25.370\;9(1) 2.783 6(1)2.783\;6(1)
GF 25.371 041 5(6)25.371\;041\;5(6) 2.783 764 5(6)2.783\;764\;5(6)
se_yerokh 2.783 7652.783\;765
36 C DKB 2.208 571 62.208\,571\,6 0.481 092 90.481\,092\,9 0.410 38(5)-0.410\;38(5) 2.279 28(5)2.279\;28(5)
GF 0.410 342 8(3)-0.410\;342\;8(3) 2.279 321 7(3)2.279\;321\;7(3)
F DKB 31.021 248 6-31.021\,248\,6 20.805 926 920.805\,926\,9 12.494 5(1)12.494\;5(1) 2.279 2(1)2.279\;2(1)
GF 12.494 643 3(4)12.494\;643\;3(4) 2.279 321 6(4)2.279\;321\;6(4)
se_yerokh 2.279 3222.279\;322
54 C DKB 1.523 080 461.523\,080\,46 0.607 564 150.607\,564\,15 0.348 78(2)-0.348\;78(2) 1.781 86(2)1.781\;86(2)
GF 0.348 757 76(4)-0.348\;757\;76(4) 1.781 886 85(4)1.781\;886\;85(4)
F DKB 9.961 275 6-9.961\,275\,6 6.651 147 56.651\,147\,5 5.091 98(4)5.091\;98(4) 1.781 85(4)1.781\;85(4)
GF 5.092 014 8(2)5.092\;014\;8(2) 1.781 886 7(2)1.781\;886\;7(2)
se_yerokh 1.781 887(2)(2)
82 C DKB 0.994 837 250.994\,837\,25 0.787 437 130.787\,437\,13 0.294 98(2)-0.294\;98(2) 1.487 29(2)1.487\;29(2)
GF 0.294 988 79(5)-0.294\;988\;79(5) 1.487 285 59(5)1.487\;285\;59(5)
F DKB 2.952 357 937-2.952\,357\,937 2.371 873 4372.371\,873\,437 2.067 76(2)2.067\;76(2) 1.487 27(2)1.487\;27(2)
GF 2.067 770 12(5)2.067\;770\;12(5) 1.487 285 62(5)1.487\;285\;62(5)
se_yerokh 1.487 286(16)(3)
92 C DKB 0.871 179 700.871\,179\,70 0.886 500 900.886\,500\,90 0.285 19(1)-0.285\;19(1) 1.472 49(1)1.472\;49(1)
GF 0.285 192 17(2)-0.285\;192\;17(2) 1.472 488 43(2)1.472\;488\;43(2)
F DKB 2.141 311 3-2.141\,311\,3 1.957 559 51.957\,559\,5 1.656 23(1)1.656\;23(1) 1.472 48(1)1.472\;48(1)
GF 1.656 240 1(1)1.656\;240\;1(1) 1.472 488 3(1)1.472\;488\;3(1)
se_yerokh 1.472 50(1)(2)
100 C GF 0.780 926 900.780\;926\;90 0.996 469 020.996\;469\;02 0.281 256 93(1)-0.281\;256\;93(1) 1.496 138 99(1)1.496\;138\;99(1)
F GF 1.737 050 7-1.737\;050\;7 1.795 888 81.795\;888\;8 1.437 300 7(2)1.437\;300\;7(2) 1.496 138 8(1)1.496\;138\;8(1)
se_yerokh 1.496 14(7)(47)
Table 6: Many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like ions, calculated in three ways: directly, (using the separation of the two-potential contribution), and (using the Sapirstein-Cheng (SC) approach). Results are given in units $F(\alpha Z)$. For $Z=10$ spherical model was used, and for other $Z$’s Fermi model was used, nuclear radii taken from se_yerokh.
direct two-pot. scheme SC scheme
$Z$ gauge $\Delta\varepsilon^{\text{Mp}}_{a}$ $\Delta\varepsilon^{2\text{p}}_{a}$ $\Delta\varepsilon^{3+}_{a}$ $\Delta\varepsilon^{\text{Mp}}_{a}$ $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ $\Delta\tilde{\varepsilon}_{a}^{3+}$ $\Delta\varepsilon_{a}^{\text{Mp}}$
10 C DKB 0.570 0(8)-0.570\;0(8) 0.791 761 7-0.791\,761\,7 0.221 95(5)0.221\;95(5) 0.569 81(5)-0.569\;81(5)
GF 0.569 8(1)-0.569\;8(1) 0.519 43(2)-0.519\;43(2) 0.050 350 0(1)-0.050\;350\;0(1) 0.569 78(2)-0.569\;78(2) 0.221 992 9(1)0.221\;992\;9(1) 0.569 768 8(1)-0.569\;768\;8(1)
F DKB $188.7(6)$ 87.239 7587.239\,75 101.435 75(2)101.435\;75(2) 188.675 50(2)188.675\;50(2)
GF 188.675 3(2)188.675\;3(2) 89.835 999(2)89.835\;999(2) 98.839 47(1)98.839\;47(1) 188.675 47(1)188.675\;47(1) 101.435 74(2)101.435\;74(2) 188.675 49(2)188.675\;49(2)
18 C DKB 0.508 2(19)-0.508\;2(19) 0.617 945 60-0.617\,945\,60 0.109 86(2)0.109\;86(2) 0.508 08(2)-0.508\;08(2)
GF 0.508 11(2)-0.508\;11(2) 0.440 235 3(3)-0.440\;235\;3(3) 0.067 867 7(5)-0.067\;867\;7(5) 0.508 103 0(7)-0.508\;103\;0(7) 0.109 843 28(1)0.109\;843\;28(1) 0.508 102 32(1)-0.508\;102\;32(1)
F DKB 55.556 9(8)55.556\;9(8) 24.685 49724.685\,497 30.871 6(1)30.871\;6(1) 55.557 0(1)55.557\;0(1)
GF 55.557 2(1)55.557\;2(1) 25.359 90(2)25.359\;90(2) 30.197 273(2)30.197\;273(2) 55.557 17(2)55.557\;17(2) 30.871 682(2)30.871\;682(2) 55.557 179(2)55.557\;179(2)
26 C DKB 0.458 72(9)-0.458\;72(9) 0.497 070 7-0.497\,070\,7 0.038 25(5)0.038\;25(5) 0.458 82(5)-0.458\;82(5)
GF 0.458 79(1)-0.458\;79(1) 0.377 228 12(3)-0.377\;228\;12(3) 0.081 559 7(5)-0.081\;559\;7(5) 0.458 787 8(5)-0.458\;787\;8(5) 0.038 282 9(4)0.038\;282\;9(4) 0.458 787 8(4)-0.458\;787\;8(4)
F DKB 25.371 3(3)25.371\;3(3) 10.809 479 410.809\,479\,4 14.561 5(1)14.561\;5(1) 25.370 9(1)25.370\;9(1)
GF 25.371 03(1)25.371\;03(1) 11.080 133 05(4)11.080\;133\;05(4) 14.290 907(2)14.290\;907(2) 25.371 040(2)25.371\;040(2) 14.561 562 1(6)14.561\;562\;1(6) 25.371 041 5(6)25.371\;041\;5(6)
36 C DKB 0.410 2(1)0.410\;2(1) 0.388 558 5-0.388\,558\,5 0.021 83(5)-0.021\;83(5) 0.410 38(5)-0.410\;38(5)
GF 0.410 34(1)-0.410\;34(1) 0.315 349 962 3(19)-0.315\;349\;962\;3(19) 0.094 992 9(4)-0.094\;992\;9(4) 0.410 342 8(4)-0.410\;342\;8(4) 0.021 784 3(3)-0.021\;784\;3(3) 0.410 342 8(3)-0.410\;342\;8(3)
F DKB 12.494 8(4)12.494\;8(4) 5.054 839 95.054\,839\,9 7.439 7(1)7.439\;7(1) 12.494 5(1)12.494\;5(1)
GF 12.494 63(2)12.494\;63(2) 5.168 459 166(5)5.168\;459\;166(5) 7.326 183 4(14)7.326\;183\;4(14) 12.494 642 6(14)12.494\;642\;6(14) 7.439 803 3(4)7.439\;803\;3(4) 12.494 643 3(4)12.494\;643\;3(4)
54 C DKB 0.348 74(3)-0.348\;74(3) 0.258 689 31-0.258\,689\,31 0.090 09(2)-0.090\;09(2) 0.348 78(2)-0.348\;78(2)
GF 0.348 758(1)-0.348\;758(1) 0.235 050 084 3(11)-0.235\;050\;084\;3(11) 0.113 707 7(2)-0.113\;707\;7(2) 0.348 757 8(2)-0.348\;757\;8(2) 0.090 068 45(5)0.090\;068\;45(5) 0.348 757 76(4)-0.348\;757\;76(4)
F DKB 5.092 1(3)5.092\;1(3) 1.901 467 11.901\,467\,1 3.190 52(4)3.190\;52(4) 5.091 98(4)5.091\;98(4)
GF 5.092 013(2)5.092\;013(2) 1.933 405 160(1)1.933\;405\;160(1) 3.158 609 4(6)3.158\;609\;4(6) 5.092 014 6(6)5.092\;014\;6(6) 3.190 547 7(2)3.190\;547\;7(2) 5.092 014 8(2)5.092\;014\;8(2)
82 C DKB 0.294 98(2)-0.294\;98(2) 0.132 956 05-0.132\,956\,05 0.162 04(2)-0.162\;04(2) 0.294 99(2)-0.294\;99(2)
GF 0.294 988 8(2)-0.294\;988\;8(2) 0.153 987 845 8(9)-0.153\;987\;845\;8(9) 0.141 000 96(11)-0.141\;000\;96(11) 0.294 988 81(11)-0.294\;988\;81(11) 0.162 032 73(5)-0.162\;032\;73(5) 0.294 988 79(5)-0.294\;988\;79(5)
F DKB 2.067 76(2)2.067\;76(2) 0.745 897 720.745\,897\,72 1.321 87(2)1.321\;87(2) 2.067 76(2)2.067\;76(2)
GF 2.067 770 03(15)2.067\;770\;03(15) 0.738 029 089 8(5)0.738\;029\;089\;8(5) 1.329 740 9(2)1.329\;740\;9(2) 2.067 770 0(2)2.067\;770\;0(2) 1.321 872 39(5)1.321\;872\;39(5) 2.067 770 12(5)2.067\;770\;12(5)
92 C DKB 0.285 2(1)-0.285\;2(1) 0.096 833 87-0.096\,833\,87 0.188 36(1)-0.188\;36(1) 0.285 19(1)-0.285\;19(1)
GF 0.285 192 0(2)-0.285\;192\;0(2) 0.132 008 645(1)-0.132\;008\;645(1) 0.153 183 54(7)-0.153\;183\;54(7) 0.285 192 19(7)-0.285\;192\;19(7) 0.188 358 30(2)-0.188\;358\;30(2) 0.285 192 17(2)-0.285\;192\;17(2)
F DKB 1.656 26(5)1.656\;26(5) 0.618 726 40.618\,726\,4 1.037 51(1)1.037\;51(1) 1.656 23(1)1.656\;23(1)
GF 1.656 240(2)1.656\;240(2) 0.598 961 3(1)0.598\;961\;3(1) 1.057 278(1)1.057\;278(1) 1.656 239(1)1.656\;239(1) 1.037 513 7(1)1.037\;513\;7(1) 1.656 240 1(1)1.656\;240\;1(1)
100 C GF 0.281 256 9(1)-0.281\;256\;9(1) 0.116 019 00(1)-0.116\;019\;00(1) 0.165 237 9(1)-0.165\;237\;9(1) 0.281 256 9(1)-0.281\;256\;9(1) 0.068 173 06-0.068\;173\;06 0.213 083 87(1)-0.213\;083\;87(1) 0.281 256 93(1)-0.281\;256\;93(1)
F GF 1.437 300 8(2)1.437\;300\;8(2) 0.534 332 3(2)0.534\;332\;3(2) 0.902 968 5(1)0.902\;968\;5(1) 1.437 300 8(3)1.437\;300\;8(3) 0.565 473 40.565\;473\;4 0.871 827 3(2)0.871\;827\;3(2) 1.437 300 7(2)1.437\;300\;7(2)
Table 7: Contributions to the self-energy correction for $2S_{1/2}$, $2P_{1/2}$, and $2P_{3/2}$ states of the hydrogen-like ions. Results are given in units $F(\alpha Z)$. For $Z=10$ spherical model was used, and for other $Z$’s Fermi model was used, nuclear radii taken from se_yerokh.
$Z$ gauge $2S_{1/2}$ $2P_{1/2}$ $2P_{3/2}$
10 C $\Delta\varepsilon_{a}^{0\text{p}}$ 7.190 427.190\;42 2.624 892.624\;89 2.775 882.775\;88
$\Delta\varepsilon_{a}^{1\text{p}}$ 1.534 43-1.534\;43 2.119 89-2.119\;89 2.038 94-2.038\;94
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.761 61(1)-0.761\;61(1) 0.619 81(1)-0.619\;81(1) 0.606 58(1)-0.606\;58(1)
$\Delta\varepsilon_{a}$ 4.894 38(1)4.894\;38(1) 0.114 82(1)-0.114\;82(1) 0.130 35(1)0.130\;35(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 2 075.579 23-2\;075.579\;23 2 196.693 6-2\;196.693\;6 2 192.070 77-2\;192.070\;77
$\Delta\varepsilon_{a}^{1\text{p}}$ 1 719.050 471\;719.050\;47 1 818.840 11\;818.840\;1 1 815.473 221\;815.473\;22
$\Delta\varepsilon_{a}^{\text{Mp}}$ 361.423 19(3)361.423\;19(3) 377.738 6(1)377.738\;6(1) 376.727 91(2)376.727\;91(2)
$\Delta\varepsilon_{a}$ 4.894 42(3)4.894\;42(3) 0.114 7(1)-0.114\;7(1) 0.130 36(2)0.130\;36(2)
$\Delta\varepsilon_{a}$, se_yerokh 4.894 384(7)4.894\;384(7) 0.114 84(2)-0.114\;84(2) 0.130 36(2)0.130\;36(2)
18 C $\Delta\varepsilon_{a}^{0\text{p}}$ 5.135 2445.135\;244 2.027 252.027\;25 2.157 9522.157\;952
$\Delta\varepsilon_{a}^{1\text{p}}$ 0.736 198-0.736\;198 1.519 14-1.519\;14 1.443 991-1.443\;991
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.699 149-0.699\;149 0.605 61(1)-0.605\;61(1) 0.573 139(2)-0.573\;139(2)
$\Delta\varepsilon_{a}$ 3.699 896(5)3.699\;896(5) 0.097 51(1)-0.097\;51(1) 0.140 821(2)0.140\;821(2)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 510.933 71-510.933\;71 547.071 56-547.071\;56 543.528 24-543.528\;24
$\Delta\varepsilon_{a}^{1\text{p}}$ 405.527 77405.527\;77 431.854 90431.854\;90 429.503 30429.503\;30
$\Delta\varepsilon_{a}^{\text{Mp}}$ 109.105 84(2)109.105\;84(2) 115.119 15(2)115.119\;15(2) 114.165 74(1)114.165\;74(1)
$\Delta\varepsilon_{a}$ 3.699 90(2)3.699\;90(2) 0.097 50(2)-0.097\;50(2) 0.140 81(1)0.140\;81(1)
$\Delta\varepsilon_{a}$, se_yerokh 3.699 8923.699\;892 0.097 511(4)-0.097\;511(4) 0.140 8190.140\;819
26 C $\Delta\varepsilon_{a}^{0\text{p}}$ 4.030 4224.030\;422 1.681 601.681\;60 1.790 2741.790\;274
$\Delta\varepsilon_{a}^{1\text{p}}$ 0.321 172-0.321\;172 1.166 08-1.166\;08 1.100 585-1.100\;585
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.649 957(1)-0.649\;957(1) 0.591 74(1)-0.591\;74(1) 0.536 070(1)-0.536\;070(1)
$\Delta\varepsilon_{a}$ 3.059 292(1)3.059\;292(1) 0.076 21(1)-0.076\;21(1) 0.153 619(1)0.153\;619(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 207.664 30-207.664\;30 224.482 21-224.482\;21 221.561 94-221.561\;94
$\Delta\varepsilon_{a}^{1\text{p}}$ 159.521 30159.521\;30 169.968 23169.968\;23 168.173 15168.173\;15
$\Delta\varepsilon_{a}^{\text{Mp}}$ 51.202 29(1)51.202\;29(1) 54.437 76(1)54.437\;76(1) 53.542 39(1)53.542\;39(1)
$\Delta\varepsilon_{a}$ 3.059 29(1)3.059\;29(1) 0.076 21(1)-0.076\;21(1) 0.153 60(1)0.153\;60(1)
$\Delta\varepsilon_{a}$, se_yerokh 3.059 292(1)3.059\;292(1) 0.076 218(1)-0.076\;218(1) 0.153 6200.153\;620
36 C $\Delta\varepsilon_{a}^{0\text{p}}$ 3.191 6173.191\;617 1.402 981.402\;98 1.483 7731.483\;773
$\Delta\varepsilon_{a}^{1\text{p}}$ 0.003 718-0.003\;718 0.871 02-0.871\;02 0.823 673-0.823\;673
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.602 925(1)-0.602\;925(1) 0.576 95(1)-0.576\;95(1) 0.488 292(1)-0.488\;292(1)
$\Delta\varepsilon_{a}$ 2.584 973(1)2.584\;973(1) 0.044 99(1)-0.044\;99(1) 0.171 807(1)0.171\;807(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 92.012 084-92.012\;084 100.545 37-100.545\;37 98.122 65-98.122\;65
$\Delta\varepsilon_{a}^{1\text{p}}$ 68.478 91368.478\;913 72.554 1872.554\;18 71.176 2471.176\;24
$\Delta\varepsilon_{a}^{\text{Mp}}$ 26.118 14(1)26.118\;14(1) 27.946 19(1)27.946\;19(1) 27.118 19(1)27.118\;19(1)
$\Delta\varepsilon_{a}$ 2.584 96(1)2.584\;96(1) 0.044 99(1)-0.044\;99(1) 0.171 79(1)0.171\;79(1)
$\Delta\varepsilon_{a}$, se_yerokh 2.584 9722.584\;972 0.044 991-0.044\;991 0.171 8080.171\;808
54 C $\Delta\varepsilon_{a}^{0\text{p}}$ 2.346 175 32.346\;175\;3 1.103 6171.103\;617 1.136 0731.136\;073
$\Delta\varepsilon_{a}^{1\text{p}}$ 0.361 787 90.361\;787\;9 0.516 470-0.516\;470 0.523 758-0.523\;758
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.547 351 8(1)-0.547\;351\;8(1) 0.561 532(3)-0.561\;532(3) 0.403 752(1)-0.403\;752(1)
$\Delta\varepsilon_{a}$ 2.160 611 4(1)2.160\;611\;4(1) 0.025 614(3)0.025\;614(3) 0.208 562(1)0.208\;562(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 32.616 71-32.616\;71 36.324 81-36.324\;81 34.414 650-34.414\;650
$\Delta\varepsilon_{a}^{1\text{p}}$ 23.445 3523.445\;35 24.194 7424.194\;74 23.197 27323.197\;273
$\Delta\varepsilon_{a}^{\text{Mp}}$ 11.331 96(1)11.331\;96(1) 12.155 67(1)12.155\;67(1) 11.425 936(5)11.425\;936(5)
$\Delta\varepsilon_{a}$ 2.160 60(1)2.160\;60(1) 0.025 60(1)0.025\;60(1) 0.208 559(5)0.208\;559(5)
$\Delta\varepsilon_{a}$, se_yerokh 2.160 612(2)(3) 0.025 6170.025\;617 0.208 5630.208\;563
82 C $\Delta\varepsilon_{a}^{0\text{p}}$ 1.703 650 81.703\;650\;8 0.860 7500.860\;750 0.829 247 300.829\;247\;30
$\Delta\varepsilon_{a}^{1\text{p}}$ 0.875 870 90.875\;870\;9 0.073 015-0.073\;015 0.274 656 46-0.274\;656\;46
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.514 155 6(2)-0.514\;155\;6(2) 0.582 124(1)-0.582\;124(1) 0.282 870 24(5)-0.282\;870\;24(5)
$\Delta\varepsilon_{a}$ 2.065 366 1(2)2.065\;366\;1(2) 0.205 611(1)0.205\;611(1) 0.271 720 60(5)0.271\;720\;60(5)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 11.119 815 49-11.119\;815\;49 12.768 56-12.768\;56 11.152 418-11.152\;418
$\Delta\varepsilon_{a}^{1\text{p}}$ 8.151 898 868.151\;898\;86 7.689 867.689\;86 6.797 1626.797\;162
$\Delta\varepsilon_{a}^{\text{Mp}}$ 5.033 279(5)5.033\;279(5) 5.284 29(1)5.284\;29(1) 4.626 973(3)4.626\;973(3)
$\Delta\varepsilon_{a}$ 2.065 363(5)2.065\;363(5) 0.205 60(1)0.205\;60(1) 0.271 718(3)0.271\;718(3)
$\Delta\varepsilon_{a}$, se_yerokh 2.065 367(23)(5) 0.205 613(1)(0) 0.271 7210.271\;721
92 C $\Delta\varepsilon_{a}^{0\text{p}}$ 1.553 6291.553\;629 0.800 3520.800\;352 0.755 3920.755\;392
$\Delta\varepsilon_{a}^{1\text{p}}$ 1.134 0511.134\;051 0.128 0270.128\;027 0.216 677-0.216\;677
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.517 176(2)-0.517\;176(2) 0.611 516(2)-0.611\;516(2) 0.243 672(1)-0.243\;672(1)
$\Delta\varepsilon_{a}$ 2.170 505(2)2.170\;505(2) 0.316 863(2)0.316\;863(2) 0.295 042(1)0.295\;042(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 8.389 628-8.389\;628 9.729 68-9.729\;68 8.095 70-8.095\;70
$\Delta\varepsilon_{a}^{1\text{p}}$ 6.410 1466.410\;146 5.762 415.762\;41 4.775 604.775\;60
$\Delta\varepsilon_{a}^{\text{Mp}}$ 4.149 985(2)4.149\;985(2) 4.284 12(1)4.284\;12(1) 3.615 13(1)3.615\;13(1)
$\Delta\varepsilon_{a}$ 2.170 502(2)2.170\;502(2) 0.316 86(1)0.316\;86(1) 0.295 03(1)0.295\;03(1)
$\Delta\varepsilon_{a}$, se_yerokh 2.170 52(2)(2) 0.316 869(2)(2) 0.295 043(0)(1)
100 C $\Delta\varepsilon_{a}^{0\text{p}}$ 1.438 749 7381.438\;749\;738 0.749 9050.749\;905 0.704 9550.704\;955
$\Delta\varepsilon_{a}^{1\text{p}}$ 1.413 259 461.413\;259\;46 0.344 7710.344\;771 0.177 247-0.177\;247
$\Delta\varepsilon_{a}^{\text{Mp}}$ 0.527 333 1(2)-0.527\;333\;1(2) 0.649 663(2)-0.649\;663(2) 0.214 176(1)-0.214\;176(1)
$\Delta\varepsilon_{a}$ 2.324 676 0(2)2.324\;676\;0(2) 0.445 013(2)0.445\;013(2) 0.313 531(1)0.313\;531(1)
F $\Delta\varepsilon_{a}^{0\text{p}}$ 6.953 875-6.953\;875 8.114 51-8.114\;51 6.400 104-6.400\;104
$\Delta\varepsilon_{a}^{1\text{p}}$ 5.597 1895.597\;189 4.824 504.824\;50 3.682 8143.682\;814
$\Delta\varepsilon_{a}^{\text{Mp}}$ 3.681 359(3)3.681\;359(3) 3.735 01(1)3.735\;01(1) 3.030 819(2)3.030\;819(2)
$\Delta\varepsilon_{a}$ 2.324 673(3)2.324\;673(3) 0.445 00(1)0.445\;00(1) 0.313 529(2)0.313\;529(2)
$\Delta\varepsilon_{a}$, se_yerokh 2.324 7(1)(8) 0.445 01(1)(9) 0.313 53(0)(2)
Table 7: Self-energy, excited states (continued).
Figure 9: Contributions to the self-energy, plotted on a logarithmic scale ($F(\alpha Z)$ units) in two gauges as a function of the nuclear charge $Z$. The zero-, one-, and many-potential contributions are shown in Fig. (a). Sum of zero- and one-potential contributions is shown in Fig. (b). Note that for negative contributions the absolute value is taken.
Appendix A Free-electron self-energy operator
We define the free-electron self-energy operator as:
[…]
$D_{\mu\nu}^{F}(\mathrm{k})=-\frac{g_{\mu\nu}}{\mathrm{k}^{2}}$ (48)
in the Feynman gauge and
$D_{\mu\nu}^{C}(\mathrm{k})=\frac{1}{\mathrm{k}^{2}}(-g_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{\mathbf{k}^{2}}+\frac{k^{0}(k_{\mu}\delta_{\nu 0}+k_{\nu}\delta_{\mu 0})}{\mathbf{k}^{2}})$ (49)
in the Coulomb gauge. In Eq. 49, $\delta_{\nu\mu}$ is the four-dimensional Kronecker delta. The expression (47) is UV divergent in both gauges. To eliminate the UV divergences, we apply the renormalization procedure. To make sense of divergent expressions, we use the dimensional regularization ($D=4-2\varepsilon$). In both gauges, we separate out only the UV divergent part by writing the self-energy in the form:
[…]
Note that we use this definition instead of writing a more familiar expression in terms of the renormalization constant $Z_{2}$. In the Feynman gauge, the second approach leads to the IR singularities of individual contributions. In the Coulomb gauge, both approaches are equally convenient, but, for consistency, we write the expression for the free-electron self-energy operator in this form. The notation $\Delta_{\varepsilon}=\frac{1}{\varepsilon}-\gamma_{E}+\ln 4\pi+\ln m$ is introduced, and the mass counterterm is
$\delta m=\frac{3\alpha}{4\pi}m(\Delta_{\varepsilon}+\frac{4}{3}).$ (51)
The renormalized part of the free-electron self-energy operator on the right-hand side of Eq. (50) can be written in the form:
$).\Sigma^{(0)}_{R}(\mathrm{p})=\frac{\alpha}{4\pi}(a(p_{0},p)+\not{\mathrm{p}}b(p_{0},p)+\gamma^{0}c(p_{0}$,p)). (52)
The coefficients $a$, $b$ and $c$ are gauge-dependent. In the Feynman gauge, they can be found, e.g., in Ref. base_1999:
$\begin{gathered}a(p_{0},p)=2m\left(1+\frac{2\rho}{1-\rho}\ln\rho\right),\\ b(p_{0},p)=-\frac{2-\rho}{1-\rho}\left(1+\frac{\rho}{1-\rho}\ln\rho\right),\\ c(p_{0},p)=0,\end{gathered}$ (53)
where $\rho=1-{\mathrm{p^{2}}}/{m^{2}}$. For the Coulomb gauge, they could be derived from the expression base_adkins of the form:
[…]
where
$\begin{gathered}X=1+\frac{p^{2}}{m^{2}}(1-x),\\ Y=1-\frac{\mathrm{p}^{2}}{m^{2}}(1-x)-i0,\\ Z=1-\frac{p_{0}^{2}}{m^{2}}(1-u)+\frac{p^{2}}{m^{2}}(1-xu)-i0.\end{gathered}$ (55)
Some integrals in this expression can be evaluated analytically, which was done, e.g., in Ref. zp_final. The latter results allows one to write for the coefficients $a$, $b$, and $c$:
$\begin{gathered}a(p_{0},p)=2m\left(1-F_{0}+\frac{\rho\log\rho}{1-\rho}\right),\\ b(p_{0},p)=\frac{(\rho-2)(1-\rho+\rho\ln\rho)}{(1-\rho)^{2}}-2F_{2}\rho+\frac{2m^{2}(F_{1}\rho\ln\rho-F_{0})}{p^{2}},\\ c(p_{0},p)=2\frac{p_{0}}{p^{2}}(F_{0}m^{2}-F_{1}m^{2}\rho\ln\rho+F_{2}\rho p^{2}),\end{gathered}$ (56)
where $F_{0}$, $F_{1}$ and $F_{2}$ are
$\begin{gathered}F_{0}=\frac{\sqrt{p^{2}+m^{2}}}{p}\ln\left|\frac{\sqrt{p^{2}+m^{2}}+p}{\sqrt{p^{2}+m^{2}}-p}\right|-2,\\ F_{1}=\frac{p_{0}}{p}\ln\left|\frac{p_{0}+p}{p_{0}-p}\right|-2,\\ F_{2}=\int_{0}^{1}\frac{\sqrt{x}\ln Xdx}{X-\rho+i0}.\end{gathered}$ (57)
Appendix B Free-electron vertex operator
In order to calculate the one-potential term, one needs the free-electron vertex function, given by:
[…]
This expression is UV divergent in both gauges. After regularization, it can be written as:
$\Gamma^{\mu}(\mathrm{p}^{\prime},\mathrm{p})=\frac{\alpha}{4\pi}\Delta_{\varepsilon}\gamma^{\mu}+\Gamma_{R}^{\mu}(\mathrm{p}^{\prime},\mathrm{p}),$ (59)
where, as in the case of the free-electron self-energy operator, the UV finite part of $\Gamma_{R}$ also turns out to be IR finite. For brevity, we give below only the time component of the operator, $\Gamma^{0}_{R}$, for the case when $p_{0}=p^{\prime}_{0}=\varepsilon_{a}$:
[…]
In the Feynman gauge:
$\begin{gathered}A=C_{5}-2+\mathrm{p}^{\prime 2}C_{11}+\mathrm{p}^{2}C_{12}+4(\mathrm{p}^{\prime}\cdot\mathrm{p})(C_{00}+C_{11}+C_{12})+m^{2}(-2C_{00}+C_{11}+C_{12}),\\ B_{1}=-4(C_{11}+C_{23}),\;B_{2}=-4(C_{00}+C_{11}+C_{12}+C_{25}),\\ C_{1}=B_{2},\;C_{2}=-4(C_{12}+C_{24}),\\ D=2(C_{00}+C_{11}+C_{12}),\\ H_{1}=4m(C_{00}+2C_{11}),\;H_{2}=4m(C_{00}+2C_{12}),\;G_{1}=G_{2}=0,\end{gathered}$ (61)
where
$\begin{gathered}C_{ij}=\int_{0}^{1}\frac{dy}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}}S_{i}K_{j},\\ S_{0,1,2}=\{-\ln X^{\prime},1-Y^{\prime}\ln X^{\prime},-\frac{1}{2}+Y^{\prime}-{Y^{\prime}}^{2}\ln X^{\prime}\},\\ K_{0,1,2,3,4,5}=\{1,y,1-y,y^{2},(1-y)^{2},y(1-y)\},\\ C_{5}=-\int_{0}^{1}dy\ln(y^{2}\mathrm{k}^{2}/\penalty 50m^{2}-y\mathrm{k}^{2}/m^{2}+1),\\ X^{\prime}=1+\frac{1}{Y^{\prime}},\;Y^{\prime}=\frac{m^{2}-y\mathrm{p}^{\prime 2}-(1-y)\mathrm{p}^{2}}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}},\end{gathered}$ (62)
and $\mathrm{k}=\mathrm{p}-\mathrm{p}^{\prime}$. In the Coulomb gauge, these coefficients can be determined from Ref. OP_CG:
[…]
and finally $A$:
$\begin{gathered}A=\varepsilon_{a}^{2}(2F_{1}-F_{2})-F_{22}+m^{2}(2F_{1}-3F_{2})+p^{\prime 2}(F_{11}-F_{8}+4F_{5}-5F_{3})-\\ -4(\mathbf{p^{\prime}}\cdot\mathbf{p})^{2}F_{13}+p^{2}(F_{10}-F_{11}-5F_{2}+5F_{3}+4F_{4}-4F_{5}-F_{7}+\\ +F_{8}+(\mathbf{p^{\prime}}\cdot\mathbf{p})(-2F_{12}+4F_{13}))+(-4F_{5}+4F_{6}+2F_{8}-2F_{9})k^{2}+\\ +(\mathbf{p^{\prime}}\cdot\mathbf{p})(4F_{1}-2F_{19}-2F_{2}-2F_{14}k^{2})-\varepsilon_{a}^{2}(B_{1}+C_{1}-D).\end{gathered}$ (64)
The coefficients $F_{1-22}$ introduced above can be determined as follows. Let us define $B_{ij}^{k}$ as
$B_{i\,j}^{k}=\begin{cases}1&\text{if }i\leq k\leq j\\ 0&\text{otherwise}\end{cases}\,\,$, (65)
then for $F_{1}-F_{6}$ one obtains
$\begin{gathered}F_{i}=\int_{0}^{1}\frac{du}{\mathrm{t}^{2}}\Big[C_{1}^{i}\delta_{1}+C_{2}^{i}-C_{3}^{i}\frac{A}{\mathrm{t}^{2}}\Big]C_{4}^{i},\\ \delta_{1}=\ln\left(\frac{\mathrm{t}^{2}+A}{A}\right),\,\,C_{1}^{i}=[1-B_{2\,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)][1-B_{4\,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)],\\ C_{2}^{i}=B_{2\,3}^{i}+\frac{1}{2}B_{4\,6}^{i},\,\,C_{3}^{i}=B_{4\,6}^{i},\,\,C_{4}^{i}=\{1,1,u,1,u,u^{2}\}.\end{gathered}$ (66)
Here $A=m^{2}-u\mathrm{p}^{2}-(1-u)\mathrm{p}^{2}$, $\mathrm{t}=u\mathrm{p}^{\prime}+(1-u)\mathrm{p}$. For $F_{7-11}$ (below the coefficient $F_{7}$ is obtained by substituting $i=1$, and so on, i.e., $F_{8}\leftrightarrow i=2$, …):
$\begin{gathered}F_{i}=\int_{0}^{1}du\Big[C_{5}^{i}\delta_{2}-C_{6}^{i}\frac{1}{t^{2}}\Big]C_{7}^{i},\\ \delta_{2}=\frac{2}{\sqrt{t^{2}C}}\tanh^{-1}\left[\left(\frac{t^{2}}{C}\right)^{1/2}\right],\,\,C_{5}^{i}=\frac{C}{t^{2}}B_{1\,3}^{i}+B_{45}^{i},\\ C_{6}^{i}=2B_{1\,3}^{i},\,\,C_{7}^{i}=\{1,u,u^{2},1,u\},\end{gathered}$ (67)
where $C=m^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+t_{0}^{2}$. For $F_{12}-F_{21}$ (below $F_{12}\leftrightarrow i=1$):
$\begin{gathered}F_{i}=\int_{0}^{1}dsdu\Big[C_{8}^{i}\delta_{3}+C_{9}^{i}\delta_{4}\Big]C_{10}^{i},\\ \delta_{3}=\frac{1}{s(t^{2})^{2}}\left\{\left(\frac{B}{st^{2}}\right)^{1/2}\tanh^{-1}\Big[\left(\frac{st^{2}}{B}\right)^{1/2}\Big]-1\right\},\,\,\delta_{4}=\frac{1}{t^{2}(B-st^{2})},\\ C_{8}^{i}=-3B_{1\,7}^{i}+2t^{2}B_{8\,10}^{i},\,\,C_{9}^{i}=B_{1\,7}^{i},\\ C_{10}^{i}=\{1,u,u^{2},s,su,su^{2},su^{3},1,s,su\},\end{gathered}$ (68)
where $B=m^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+st_{0}^{2}$. And finally $F_{22}$ is given by
$F_{22}=\int_{0}^{1}du\ln(u^{2}\mathrm{k}^{2}/\penalty 50m^{2}-u\mathrm{k}^{2}/\penalty 50m^{2}+1).$ (69)
Appendix C Many-potential term
Reduced matrix elements can be calculated using the PW expansion of the photon propagator. Let us illustrate this with the example of the Feynman gauge, where, considering the operator $I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ from Eq. (5), one can use the standard expression
$I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})\sim\frac{e^{i\omega r_{12}}}{r_{12}}=4\pi i\omega\sum_{L=0}^{\infty}\sum_{M=-L}^{L}j_{L}(\omega r_{<})h_{L}^{(1)}(\omega r_{>})Y_{LM}^{*}(\mathbf{\hat{r}_{1}})Y_{LM}(\mathbf{\hat{r}_{2}}),$ (70)
where $j_{L}$ and $h_{L}^{(1)}$ are the spherical Bessel and Hankel functions of the first kind respectively and $r_{>}=\max(r_{1},r_{2})$, $r_{<}=\min(r_{1},r_{2})$. Then, one can obtain
$\begin{gathered}\langle ab||I_{F}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}dr_{2}\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(\omega,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}[J]g_{L}(\omega,r_{1},r_{2})D_{JL,ac}(r_{1})D_{JL,bd}(r_{2})\}\end{gathered}$ (71)
for the Feynman gauge, and
$\begin{gathered}\langle ab||I_{C}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}dr_{2}\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(0,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}a_{JL}g_{L}(\omega,r_{1},r_{2})D_{JL,ac}(r_{1})D_{JL,bd}(r_{2})\\ +(-1)^{J+1}b_{J}[g_{J}^{\text{ret}}(\omega,r_{1},r_{2})D_{JJ+1,ac}(r_{1})D_{JJ-1,bd}(r_{2})+g_{J}^{\text{ret}}(\omega,r_{2},r_{1})D_{JJ-1,ac}(r_{1})D_{JJ+1,bd}(r_{2})]\}\end{gathered}$ (72)
for the Coulomb gauge. In the above expressions, coefficients that include the radial parts of wave functions are:
$\begin{gathered}A_{ab}(r_{1})=G_{a}(r_{1})G_{b}(r_{1})+F_{a}(r_{1})F_{b}(r_{1}),\\ D_{JL,ab}(r_{1})=G_{a}(r_{1})F_{b}(r_{1})H_{L}^{J}(\kappa_{a},-\kappa_{b})-F_{a}(r_{1})G_{b}(r_{1})H_{L}^{J}(-\kappa_{a},\kappa_{b}),\\ G_{a}(r)=rg_{a}(r),\quad F_{a}(r)=rf_{a}(r).\end{gathered}$ (73)
The numerical coefficients are:
$\begin{gathered}a_{JL}=\begin{cases}J+1,&\text{for}\ L=J-1\\ 2J+1,&\text{for}\ L=J\\ J,&\text{for}\ L=J+1\end{cases}\;\;,\\ b_{J}=\sqrt{J(J+1)}\frac{\sqrt{[J+1][J-1]}}{[J]},\\ [a]=2a+1.\end{gathered}$ (74)
When considering the photon propagator, the function $g_{J}$ arises, which includes
$\begin{gathered}g_{J}(\omega,r_{1},r_{2})=i[J]\omega j_{J}(\omega r_{<})h_{J}^{(1)}(\omega r_{>}),\\ g_{J}(0,r_{1},r_{2})=\frac{r_{<}^{J}}{r_{<}^{J+1}}.\end{gathered}$ (75)
In the Coulomb gauge, we define the function $g_{J}^{\text{ret}}$:
$g_{J}^{\text{ret}}(\omega,r_{1},r_{2})=\begin{cases}i[J]\omega j_{J+1}(\omega r_{1})h_{J-1}^{(1)}(\omega r_{2}),&\text{for}\ r_{1}<r_{2}\\ i[J]\omega j_{J-1}(\omega r_{2})h_{J+1}^{(1)}(\omega r_{1})-\frac{[J]^{2}r_{2}^{J-1}}{\omega^{2}r_{1}^{J+2}},&\text{otherwise}\end{cases}\;\;.$ (76)
Angular coefficients are:
[…]
Note that (72) is consistent with Ref. mel_yerokh, except for the typo in Eq. (B6) there.
Appendix D Derivative of the free-electron self-energy operator
Let us represent the second derivative of the function $\Sigma^{(0)}_{R}$ with respect to the time component of the four-vector $\mathrm{p}$ in the form:
$\frac{\partial^{2}\Sigma_{R}^{(0)}(\mathrm{p})}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}=\frac{\alpha}{4\pi}\left\{(N_{1}+N_{2}\gamma^{0}-N_{3}(\boldsymbol{\gamma}\cdot\mathbf{p})\right\},$ (78)
where the coefficients $N_{1}$, $N_{2}$ and $N_{3}$ can be obtained by differentiating the formula (52):
$\begin{gathered}N_{1}=\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=\varepsilon_{a}\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial b(p_{0},p)}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial^{2}c(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{3}=\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\end{gathered}$ (79)
In the Feynman gauge, the derivatives of the coefficients $a$ and $b$ are:
$\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}=-\frac{8}{m(1-\rho)}\left\{1+\frac{1}{1-\rho}\left[\ln\rho-\frac{2p_{0}^{2}}{m^{2}}\left(\frac{1+\rho}{\rho}+\frac{2}{1-\rho}\ln\rho\right)\right]\right\},$ (80)
$\frac{\partial b(p_{0},p)}{\partial p_{0}}=\frac{2p_{0}}{m^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\ln\rho\right\},$ (81)
$\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}=\frac{2}{m^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\left[\ln\rho-\frac{p_{0}^{2}}{m^{2}}\left(\frac{2+5\rho-\rho^{2}}{\rho}+\frac{6}{1-\rho}\ln\rho\right)\right]\right\}.$ (82)
In the Coulomb gauge, it is more convenient to differentiate the free-electron self-energy operator in the form (54):
[…]
Then, for the coefficients $N_{1}$ - $N_{3}$ we have:
$\begin{gathered}N_{1}=-2m\int_{0}^{1}dx\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=2\int_{0}^{1}dx(1-x)\left(\varepsilon_{a}\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+2\frac{\partial\ln Y}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\\ N_{3}=2\int_{0}^{1}dx\left((1-x)\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}-\sqrt{x}\int_{0}^{1}du\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\end{gathered}$ (84)
where
$\begin{gathered}\frac{\partial\ln Y}{\partial p_{0}}=\frac{2p_{0}(x-1)}{m^{2}-(p_{0}^{2}-p^{2})(1-x)-i0},\\ \frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}=\frac{2(x-1)(m^{2}+(p_{0}^{2}+p^{2})(1-x))}{(m^{2}-(p_{0}^{2}-p^{2})(1-x)-i0)^{2}},\\ \frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}=\frac{2(u-1)(m^{2}+p_{0}^{2}(1-u)+p^{2}(1-xu))}{(m^{2}-p_{0}^{2}(1-u)+p^{2}(1-xu)-i0)^{2}}.\end{gathered}$ (85)
For the coefficient $Z$, the second derivative can be integrated analytically with respect to $u$:
$\begin{gathered}\int_{0}^{1}du\left(\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\right)=2\Big[\frac{p_{0}^{4}-3p_{0}^{2}(m^{2}+p^{2})-xp^{2}(m^{2}+p^{2}-3p_{0}^{2})}{m^{2}\rho(p_{0}^{2}-xp^{2})^{2}}\\ -\frac{(m^{2}+p^{2}(1-x))(3p_{0}^{2}+xp^{2})}{(p_{0}^{2}-xp^{2})^{3}}\ln\left(\frac{m^{2}+p^{2}-p_{0}^{2}}{m^{2}+p^{2}(1-x)}\right)\Big].\end{gathered}$ (86)
Some integrals over $x$ from the expressions (84) can be evaluated analytically using the following master integrals:
$\begin{gathered}\int_{0}^{1}dy\frac{y}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{p_{0}^{2}-p^{2}}\left(\frac{\ln\rho}{p_{0}^{2}-p^{2}}+\frac{1}{m^{2}\rho}\right),\\ \int_{0}^{1}dy\frac{y^{2}}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{m^{2}}{(p_{0}^{2}-p^{2})^{3}}\left(\frac{1}{\rho}+2\ln\rho-\rho\right),\\ \int_{0}^{1}dy\frac{y^{3}}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{(p_{0}^{2}-p^{2})^{4}}(p_{0}^{4}-2(p_{0}^{2}+2m^{2})p^{2}+\frac{2m^{4}}{\rho}+6m^{4}\ln\rho\\ +4m^{2}p_{0}^{2}+(p^{2})^{2}-2m^{4}),\\ \int_{0}^{1}dy\frac{y^{2}}{m^{2}-(p_{0}^{2}-p^{2})y}=-\frac{1}{2}\left(\frac{1}{p_{0}^{2}-p^{2}}+\frac{2m^{2}}{(p_{0}^{2}-p^{2})^{2}}+\frac{2m^{4}\ln\rho}{(p_{0}^{2}-p^{2})^{3}}\right).\end{gathered}$ (87)