### v2

# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

Abstract
The energy correction associated with the self-energy diagram is the leading (in magnitude) and fundamental (in significance) contribution to the Lamb shift in highly charged ions. Conventional approaches to this correction rely on partial-wave expansions, which is a stumbling block limiting accuracy. To elucidate the issue, we perform a comprehensive comparative analysis of partial-wave-expansion convergence between two gauges: Feynman and Coulomb. Some tricks for improving the convergence are discussed as well.
I Introduction
Quantum electrodynamics (QED) effects are indispensable for the accurate theoretical description of various atomic systems. To account for relativistic effects, the corresponding calculations are to be performed non-perturbatively in the nuclear-strength parameter $\alpha Z$ ($\alpha$ is the fine-structure constant and $Z$ is the nuclear charge). In the present work, we employ the related formalism known as the Furry picture of QED Furry (1951). We note that the corresponding approach is relevant not only for highly charged ions Mohr et al. (1998); Yerokhin et al. (2003); Sapirstein and Cheng (2008); Glazov et al. (2011); Volotka et al. (2013); Shabaev et al. (2018); Indelicato (2019); Shabaev (2024), but also finds application in the case of light systems including the hydrogen atom, see, e.g., Refs. Jentschura et al. (1999); Yerokhin and Shabaev (2015b); Yerokhin et al. (2024); Mohr et al. (2025). An alternative, and in some sense complementary, nonrelativistic QED approach Caswell and Lepage (1986), primarily applicable to low-$Z$ systems, is beyond the scope of this paper. One of the leading QED corrections to energy levels originates from the one-electron self-energy (SE) diagram, shown in Fig. 1.
However, high-precision evaluation of the corresponding contribution holds significance not only for studying the Lamb shift and transition energies in hydrogen-like ions and other few-electron systems Stöhlker et al. (2000); Gumberidze et al. (2005); Schweppe et al. (1991); Brandau et al. (2003); Beiersdorfer et al. (2005); Trassinelli et al. (2009, 2011); Loetzsch et al. (2024) (for related theory, see, e.g., Refs. Yerokhin et al. (2006); Kozhedub et al. (2008); Sapirstein and Cheng (2011); Yerokhin and Shabaev (2015a); Artemyev et al. (2005); Kozhedub et al. (2019)), but also for investigating a wide range of atomic properties. Accurate calculations of the SE contribution is a cornerstone in the ab initio treatment of the one-electron QED corrections to the $g$ factor Beier et al. (2000); Yerokhin et al. (2002), hyperfine structure Yerokhin and Shabaev (2001); Sapirstein and Cheng (2001); Yerokhin and Jentschura (2008, 2010a), quadratic Zeeman splitting Agababaev et al. (2025), nuclear-magnetic shielding Yerokhin et al. (2011, 2012), E1 Kozlov et al. (2024) and M1 Volotka et al. (2006) transition amplitudes. The SE contribution also plays an important role in the treatment of various many-electron QED effects, see, e.g., Yerokhin et al. (1997); Persson et al. (1996); Yerokhin et al. (1999); Volotka et al. (2009, 2014) and references therein. In all these applications it is often necessary to calculate off-diagonal matrix elements of the SE operator. However, the one-electron (diagonal) SE contribution is the simplest example that one can and should use to probe different methods and numerical approaches.
Within the Furry-picture formalism, an issue of representing the bound-electron propagator in calculations naturally arises. Unlike the free-electron propagator, the propagator in the presence of external field generally does not have a closed-form expression. However, when the spherical symmetry is present, as is the case of atomic electrons, the partial-wave (PW) expansion helps to construct the propagator. With this in mind, we use two state-of-the-art approaches: the Green’s function (GF) method Mohr (1974b, a); Soff and Mohr (1988); Yerokhin and Maiorova (2020), and the finite-basis-set (FBS) method Johnson et al. (1988b); Tupitsyn and Loginov (2003); Kotochigova et al. (2007); Shabaev et al. (2004).
Within the first approach, the bound-electron propagator is represented in terms of the Dirac-equation solutions bounded at infinity and at origin. For the point-nucleus Coulomb potential, the corresponding solutions are expressed analytically via the Whittaker functions Wichmann and Kroll (1956). For more realistic models of nuclear-charge distribution, e.g., the Fermi model, or other spherically symmetric potentials, e.g., some local screening potentials, these solutions can be found numerically. This approach enables one to perform calculations with a high numerical accuracy when large numerical cancellations occur or when the PW expansion does not converge rapidly. Note that the Green’s function, considered as a function of any of its radial arguments, is discontinuous, when the arguments coincide. Therefore, obtaining accurate results for the matrix elements of the Green’s function requires special care.
Within the second approach, the electron Green’s function is represented using the FBS’s. The basis sets for the Dirac equation can be constructed, e.g., from the B splines Sapirstein and Johnson (1996); Johnson et al. (1988a); Shabaev et al. (2004), Sturm orbitals Tupitsyn and Loginov (2003); Kotochigova et al. (2007); Ivanov et al. (2024) or Gaussians Boys and Egerton (1950); Ferenc et al. (2025). The corresponding numerical procedures can easily incorporate any spherically symmetric potential. Moreover, for systems that do not possess spherical symmetry, the FBSs can also be readily prepared E. B. Rozenbaum, D. A. Glazov, V. M. Shabaev, K. E. Sosnova, and D. A. Telnov (2014); A. N. Artemyev and A. Surzhykov (2015); 1. The advantage of this method is that it allows one to easily separate out and exclude, if necessary, the contribution of a specific bound state to the Green’s function. Note also that this method provides an approximation to the Green’s function, which is a continuous function of the radial arguments. Nevertheless, the basis-set method has some important drawbacks when compared to the GF approach. Because of an additional parameter, the number of basis functions $N$, the obtained results should be extrapolated to $N\rightarrow\infty$. Typically, this sets a limit on the computational accuracy. Furthermore, the number of partial waves, usually accessible in practical calculations by means of the FBS method, is limited and smaller than that in the Green’s function method Yerokhin and Maiorova (2020). In this paper, we use the FBS method with the basis obtained from B splines within the dual-kinetic-balance (DKB) approach Shabaev et al. (2004). The latter method eliminates the so-called spurious states and establishes the correct asymptotics of wave functions in the non-relativistic limit. The effectiveness of this basis for calculating the SE correction has been demonstrated, e.g., in Ref. Shabaev et al. (2004) for the Coulomb potential and in Ref. Glazov et al. (2006) for local screening potentials.
The SE diagram contains ultraviolet (UV) divergences. In the present work, these divergences are renormalized using the conventional approach based on the potential expansion of the bound-electron propagator Snyderman (1991). Within this expansion, three contributions naturally arise: zero-, one-, and many-potential terms. Only the zero- and one-potential terms, which correspond to the first and second terms of the potential expansion, are UV divergent. They are treated separately in momentum space together with the mass counterterm. The remaining many-potential term is considered in coordinate space using the PW expansion. The truncation of this expansion and subsequent estimation of the residual is generally the main source of numerical uncertainty of the results.
The total SE contribution is gauge invariant, unlike the individual terms of its potential expansion. It is therefore intriguing to compare the magnitude of these terms in different gauges of the photon propagator. Starting from the pioneering works Mohr (1974b, a), the SE correction is calculated predominantly in the Feynman gauge, see, e.g., Refs. Mohr (1974a, 1982); Yerokhin and Shabaev (1999); Jentschura et al. (2001); Jentschura and Mohr (2004); Yerokhin and Harman (2017) and references therein. In the Coulomb gauge, similar calculations were carried out in Refs. Hedendahl and Holmberg (2012); Yerokhin et al. (2025a). These works suggest that the many-potential contribution is significantly smaller in the Coulomb gauge. However, a smaller absolute value does not necessarily imply better convergence of the PW expansion. Therefore, a thorough comparative analysis of the many-potential contributions in both gauges is in demand. Note that other covariant gauges, aside from the Feynman one, can also be applied to the SE calculations. For example, the Fried-Yennie gauge Adkins (1993) was also considered in Ref. Yerokhin et al. (2025a). In the present work, however, we restrict our consideration only to the Feynman and Coulomb gauges.
Another important issue in the SE calculations is to somehow accelerate the convergence of PW expansions mentioned above. It is reasonable to assume that the poor convergence of the many-potential contribution is due to the leading term of its potential expansion. Therefore, the natural first step is to additionally subtract the two-potential term, which should improve the convergence of the remainder. The subtracted term has to be evaluated separately with high precision within some alternative method. In the following, we will refer to this approach as the two-potential scheme. It should be noted, however, that the direct calculation of the two-potential term in momentum space in a closed form, i.e., not resorting to the PW expansion, turns out to be a challenging problem Yerokhin and Jentschura (2008, 2010b).
Another convergence-acceleration approach, hereafter refered to as the Sapirstein-Cheng (SC) scheme, was proposed recently in Ref. Sapirstein and Cheng (2023). Within the framework of this promising scheme, the quasi-two-potential contribution, which is an approximation to the two-potential term, is subtracted from the many-potential term instead. In contrast to the exact two-potential term, its quasi-two-potential counterpart can be easily calculated in momentum space. For the point-nucleus case, calculations of the SE correction using this acceleration scheme have been performed in Ref. Yerokhin et al. (2025a) for different gauges. However, the case of an extended nucleus is also worth examining.
Yet another convergence-acceleration approach was developed earlier for the SE contribution to the Lamb shift in Ref. Yerokhin et al. (2005). We refer to it as the Yerokhin, Pachucki, and Shabaev (YPS) scheme. This scheme also improves the convergence by subtracting an approximation to the many-potential contribution, that can be calculated in a closed form. The YPS scheme was initially developed for the calculations in the Feynman gauge, and was later generalized to the Coulomb gauge in Ref. Yerokhin et al. (2025a). Judging by the results of the latter paper, it is an extremely powerful scheme, however, it is difficult to generalize it to more complex diagrams, unlike other convergence-acceleration approaches. For instance, the two-potential scheme has recently been applied in the QED calculations of the quadratic Zeeman effect Agababaev et al. (2025), and some modifications of the SC scheme have been applied to the calculations of the two-electron SE Malyshev et al. (2024) and two-loop SE Yerokhin et al. (2024, 2025b) contributions. For this reason, we do not consider the YPS scheme in the current study.
Thus we thoroughly examine the application of several different schemes to the SE calculations in both the Coulomb and Feynman gauges. We conduct an independent of Ref. Yerokhin et al. (2025a) analysis, consider in details the issue of the PW-expansion convergence, evaluate the SE correction for finite-size nuclei, and collect the necessary formulas in a convenient form that allow one to make the corresponding calculations.
Throughout this article the relativistic units ($\hbar=c=1$) and the Heaviside charge unit ($\alpha=e^{2}/(4\pi),e<0$) are used. We use roman style ($\rm{p}$) for four-vectors, bold style ($\bf{p}$) for space vectors, and italic style ($p$) for scalars.
II Self-energy contribution
II.1 Basic formalism
Within the framework of the Furry picture, the description of a bound electron starts with the Dirac equation:
$\left[-i(\boldsymbol{\alpha}\cdot\boldsymbol{\nabla})+\beta m_{e}+V(\mathbf{r})\right]|a\rangle=\varepsilon_{a}|a\rangle,$ (1)
where $|a\rangle$ and $\varepsilon_{a}$ are the Dirac wave function and relativistic energy, $\boldsymbol{\alpha}$ and $\beta$ are the Dirac matrices, $V(\mathbf{r})$ is the binding potential, which is a function of the position vector $\mathbf{r}$. In the following, we consider only hydrogen-like ions where the binding potential is spherically symmetric, $V(\mathbf{r})=V(r)$, $r=|\mathbf{r}|$. Then, the solution of Eq. (1) can be represented in the form:
$|a\rangle\leftrightarrow\psi_{a}(\mathbf{r})={\begin{pmatrix}g_{a}(r)\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{r}})\\ if_{a}(r)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{r}})\end{pmatrix}}$, (2)
[…]
Figure 1: Self-energy diagram with the related mass counterterm. The double line indicates the electron propagator in the external field of the nucleus, the wavy line denotes the photon propagator, and the cross stands for the mass counterterm.
The energy shift $\Delta\varepsilon_{a}$ of the bound state $|a\rangle$ due to the first-order self-energy correction, which is graphically represented in Fig. 1, is given by the real part of the following expression Mohr et al. (1998):
$\begin{gathered}\Delta\varepsilon_{a}=2i\alpha\int_{-\infty}^{\infty}d\omega\int d^{3}\mathbf{r}_{1}d^{3}\mathbf{r}_{2}\psi^{{\dagger}}_{a}(\mathbf{r}_{1})\alpha^{\mu}\\ \times G(\varepsilon_{a}-\omega,\mathbf{r}_{1},\mathbf{r}_{2})\alpha^{\nu}\psi_{a}(\mathbf{r}_{2})D_{\mu\nu}(\omega,\mathbf{r}_{12})\\ -\delta m\int d^{3}\mathbf{r}{\psi}^{\dagger}_{a}(\mathbf{r})\beta\psi_{a}(\mathbf{r}),\end{gathered}$ (3)
where $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$, $D_{\mu\nu}(\omega,\mathbf{r}_{12})$ is the photon propagator, $\alpha^{\mu}=(1,{\boldsymbol{\alpha}})$, $G(\varepsilon,\mathbf{r}_{1},\mathbf{r}_{2})$ is the bound-electron Green’s function, and $\delta m$ is the mass counterterm. Let us define the gauge-dependent operator $I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ according to
$I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=e^{2}\alpha^{\nu}\alpha^{\mu}D_{\mu\nu}(\omega,\mathbf{r}_{12}).$ (4)
In the Feynman and Coulomb gauges, this operator takes the forms $I_{\text{F}}$ and $I_{\text{C}}$, respectively,
$I_{\text{F}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[1-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\right]\frac{e^{i\hat{\omega}r_{12}}}{r_{12}},$ (5)
$I_{\text{C}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[\frac{1}{r_{12}}-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\frac{e^{i\hat{\omega}r_{12}}}{r_{12}}+(\boldsymbol{\alpha}_{1}\cdot\boldsymbol{\nabla}_{1})(\boldsymbol{\alpha}_{2}\cdot\boldsymbol{\nabla}_{2})\frac{e^{i\hat{\omega}r_{12}}-1}{\omega^{2}r_{12}}\right],$ (6)
where $r_{12}=|\mathbf{r}_{12}|$, $\hat{\omega}=\sqrt{\omega^{2}+i0}$, and the branch of the square root is fixed with the condition $\text{Im \,}\sqrt{\omega^{2}+i0}>0$.
The bound-electron propagator $G$ can be expressed using the spectral representation,
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{n}\frac{\psi_{n}(\mathbf{r}_{1})\psi_{n}^{\dagger}(\mathbf{r}_{2})}{\omega-\varepsilon_{n}(1-i0)},$ (7)
where the summation runs over the complete Dirac spectrum. In shortened form, the formula (7) can be written as follows:
$G(\omega)=\sum_{n}\frac{|n\rangle\langle n|}{\omega-\varepsilon_{n}^{-}},$ (8)
where the notation $\varepsilon_{n}^{-}=\varepsilon_{n}(1-i0)$ was introduced. Then, the energy shift (3) can be expressed as:
$\Delta\varepsilon_{a}=\langle a|\gamma^{0}\Big[\Sigma(\varepsilon_{a})-\delta m\Big]|a\rangle,$ (9)
where $\gamma^{0}\equiv\beta$ and the diagonal matrix element of the one-loop self-energy operator $\Sigma(E)$ is given by:
$\langle a|\gamma^{0}\Sigma(E)|a\rangle=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle an|I(\omega)|na\rangle}{E-\omega-\varepsilon_{n}^{-}}.$ (10)
The expression (10) contains the UV divergences. To properly treat them, we expand the bound-electron propagator $G$ in powers of the binding potential $V$ Blundell and Snyderman (1991). The potential expansion can be conveniently written as:
$\begin{gathered}G(\omega)=\sum_{f}\frac{|f\rangle\langle f|}{\omega-\varepsilon_{f}^{-}}+\sum_{f_{1},f_{2}}\frac{|f_{1}\rangle\langle f_{1}|V|f_{2}\rangle\langle f_{2}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{f_{2}}^{-})}+\\ +\sum_{f_{1},f_{2},n}\frac{|f_{1}\rangle\langle f_{1}|V|n\rangle\langle n|V|f_{2}\rangle\langle f_{2}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{n}^{-})(\omega-\varepsilon_{f_{2}}^{-})},\end{gathered}$ (11)
where the states $|n\rangle$ correspond to the bound-electron spectrum, the states $|f\rangle$, $|f_{1}\rangle$, and $|f_{2}\rangle$ correspond to the free-electron spectrum, and the last term is expressed via the bound-electron Green’s function $G(\omega)$ itself and encompasses an infinite number of terms corresponding to the two or more interactions with the nucleus. As before, all summations in Eq. (11), are carried out over the complete spectra for the corresponding Dirac Hamiltonians.
Figure 2: Potential expansion of the self-energy diagram. The dashed line ended by a rhombus denotes the interaction with the binding potential. The single line denotes the free-electron propagator. The mass counterterm is not shown.
Substituting the potential expansion (11) into Eq. (10) results in the SE diagram being split into three contributions, as shown graphically in Fig. 2. These contributions are commonly referred to as the zero-, one-, and many-potential terms. The zero- and one-potential terms diverge, whereas the many-potential term is UV finite. The UV-divergent terms have to be renormalized along with the mass counterterm. For this aim, they are considered in momentum space, and the dimensional regularization is applied to isolate and eliminate the UV divergences. This renormalization procedure yields the UV-finite contributions $\Delta\varepsilon_{a}^{\text{0p}}$ and $\Delta\varepsilon_{a}^{\text{1p}}$ to the Lamb shift. The many-potential term $\Delta\varepsilon_{a}^{\text{Mp}}$ is calculated in the coordinate space. Finally, the SE correction to energy levels is expressed as the gauge-invariant sum of three contributions:
$\begin{gathered}\Delta\varepsilon_{a}=\Delta\varepsilon_{a}^{\text{0p}}+\Delta\varepsilon_{a}^{\text{1p}}+\Delta\varepsilon_{a}^{\text{Mp}}.\end{gathered}$ (12)
In the following subsections we consider in detail these terms and some specific methods to improve the convergence of the partial-wave expansion in the many-potential part.
II.2 Zero- and one-potential terms
The Fourier transform of the wave function (2), which solves the Dirac equation (1) in a spherically-symmetric potential, reads as:
$\psi_{a}(\mathbf{p})=\int d^{3}{\mathbf{r}}e^{-i({\mathbf{p\cdot r}})}\psi_{a}({\mathbf{r}})=i^{-l_{a}}{\begin{pmatrix}\widetilde{g}_{a}({p})\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{p}})\\ \widetilde{f}_{a}(p)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{p}})\end{pmatrix}}$, (13)
where $\mathbf{\hat{p}}=\mathbf{p}/p$, $l_{a}=|\kappa_{a}+1/2|-1/2$, and for the bound state $\widetilde{g}_{a}(p)=\widetilde{g}_{{n_{r}}_{a}\kappa_{a}}(p)$ and $\widetilde{f}_{a}(p)=\widetilde{f}_{{n_{r}}_{a}\kappa_{a}}(p)$. In the point-nucleus case, there are analytical expressions for $\widetilde{g}_{a}(p)$ and $\widetilde{f}_{a}(p)$, which can be found, e.g., Ref. Mohr et al. (1998). For an arbitrary potential, they can be obtained numerically according to:
$\widetilde{g}_{a}(p)=4\pi\int_{0}^{\infty}drr^{2}j_{l_{a}}(pr)g_{a}(r),$ (14)
$\widetilde{f}_{a}(p)=-4\pi\frac{\kappa_{a}}{|\kappa_{a}|}\int_{0}^{\infty}drr^{2}j_{2j_{a}-l_{a}}(pr)f_{a}(r),$ (15)
where $j_{l}$ is the spherical Bessel function of the first kind.
The zero-potential term is expressed as:
$\begin{gathered}\Delta\varepsilon_{a}^{\text{0p}}=\int\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\psi_{a}^{\dagger}(\mathbf{p})\gamma^{0}\Sigma^{(0)}_{R}(\mathrm{p})\psi_{a}(\mathbf{p}),\end{gathered}$ (16)
[…]
$\Delta\varepsilon_{a}^{\text{0p}}=\frac{\alpha}{4\pi}\int_{0}^{\infty}\frac{p^{2}dp}{(2\pi)^{3}}\{a(\varepsilon_{a},p)[\widetilde{g}_{a}^{2}-\widetilde{f}_{a}^{2}]\\ +b(\varepsilon_{a},p)[\varepsilon_{a}(\widetilde{g}_{a}^{2}+\widetilde{f}_{a}^{2})+2p\widetilde{g}_{a}\widetilde{f}_{a}]+c(\varepsilon_{a},p)[\widetilde{g}_{a}^{2}+\widetilde{f}_{a}^{2}]\},$ (17)
where the dependence of $\widetilde{g}_{a}$ and $\widetilde{f}_{a}$ on $p$ is omitted for brevity.
The one-potential term is expressed as:
$\begin{gathered}\Delta\varepsilon_{a}^{\text{1p}}=\int\frac{d^{3}\mathbf{p}^{\prime}d^{3}\mathbf{p}}{(2\pi)^{6}}\psi_{a}^{\dagger}(\mathbf{p}^{\prime})\gamma^{0}\Gamma^{(0)}_{R}(\mathrm{p}^{\prime},\mathrm{p})\widetilde{V}(|\mathbf{p}^{\prime}-\mathbf{p}|)\psi_{a}(\mathbf{p}),\end{gathered}$ (18)
where $\Gamma^{(0)}_{R}$ is the renormalized free-particle vertex function, the explicit form of which with the gauge-dependent coefficients $A$-$G_{2}$ is given in the Appendix B in Eq. (64), $\widetilde{V}$ is the Fourier transform of the binding potential, and $\mathrm{p}=(\varepsilon_{a},\mathbf{p})$ and $\mathrm{p}^{\prime}=(\varepsilon_{a},\mathbf{p}^{\prime})$. In this case, the angular integration can be performed using the identity:
$\frac{1}{2j+1}\sum_{m}\chi_{\kappa m}^{\dagger}(\mathbf{\hat{p}}^{\prime})\chi_{\kappa m}(\mathbf{\hat{p}})=\frac{1}{4\pi}P_{l}(z),$ (19)
where $z$ is the cosine of the angle between the vectors $\mathbf{p}$ and $\mathbf{p}^{\prime}$, namely, $z=(\mathbf{\hat{p}}\cdot\mathbf{\hat{p}^{\prime}})$, and $P_{l}$ is the Legendre polynomial. The final expression for the one-potential term reads as
$\begin{gathered}\Delta\varepsilon_{a}^{\text{1p}}=\frac{\alpha}{2(2\pi)^{6}}\int_{0}^{\infty}p^{\prime 2}dp^{\prime}\int_{0}^{\infty}p^{2}dp\int_{-1}^{1}dz\widetilde{V}(q)\{X_{1}P_{l_{a}}(z)+X_{2}P_{2j_{a}-l_{a}}(z)\},\end{gathered}$ (20)
where $q^{2}=p^{2}+p^{\prime 2}-2pp^{\prime}z$. The coefficients $X_{1}$ and $X_{2}$ are defined as:
$X_{1}$ $=A\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{1}^{\prime}\widetilde{g}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{g}^{\prime}_{a}K_{1}$
$+DK_{1}K_{1}^{\prime}+\varepsilon_{a}(H_{1}+H_{2})\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+G_{1}K_{1}^{\prime}\widetilde{g}_{a}+G_{2}\widetilde{g}^{\prime}_{a}K_{1},$ (21)
$X_{2}$ $=A\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{2}^{\prime}\widetilde{f}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{f}^{\prime}_{a}K_{2}$
$+DK_{2}K_{2}^{\prime}-\varepsilon_{a}(H_{1}+H_{2})\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}-G_{1}K_{2}^{\prime}\widetilde{f}_{a}-G_{2}\widetilde{f}^{\prime}_{a}K_{2},$ (22)
where
$K_{1}=\varepsilon_{a}\widetilde{g}_{a}+p\widetilde{f}_{a},\;K_{1}^{\prime}=\varepsilon_{a}\widetilde{g}^{\prime}_{a}+p^{\prime}\widetilde{f}^{\prime}_{a},\;K_{2}=\varepsilon_{a}\widetilde{f}_{a}+p\widetilde{g}_{a},\;K_{2}^{\prime}=\varepsilon_{a}\widetilde{f}^{\prime}_{a}+p^{\prime}\widetilde{g}^{\prime}_{a}.$ (23)
For brevity, the dependence of wave functions on $p$ and $p^{\prime}$ is omitted. For the functions of $p^{\prime}$, an additional prime is added. Thus, the shorthand notations are $\widetilde{g}_{a}=\widetilde{g}_{a}(p)$, $\widetilde{f}_{a}=\widetilde{f}_{a}(p)$, $\widetilde{g}^{\prime}_{a}=\widetilde{g}_{a}(p^{\prime})$, and $\widetilde{f}^{\prime}_{a}=\widetilde{f}_{a}(p^{\prime})$.
II.3 Many-potential term
As noted above, the remaining many-potential contribution is calculated in coordinate space using the PW expansion. Therefore, let us start with the PW expansion of the bound-electron propagator (7), inspired by the explicit form of the Dirac-equation solution (2):
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{\kappa_{n}}{\begin{pmatrix}G_{{\kappa}_{n}}^{11}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{++}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})&-iG_{\kappa_{n}}^{12}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{+-}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})\\ iG_{\kappa_{n}}^{21}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{-+}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})&G_{\kappa_{n}}^{22}(\omega,{r_{1},r_{2}})\pi_{\kappa_{n}}^{--}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})\end{pmatrix}}$, (24)
where $\pi^{\pm\pm}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})=\sum_{m_{n}}\chi_{\pm{\kappa_{n}}{m_{n}}}(\mathbf{\hat{r}}_{1})\chi_{\pm{\kappa_{n}}{m_{n}}}^{\dagger}(\mathbf{\hat{r}}_{2})$ and the components of the radial Green’s function $G_{\kappa_{n}}^{ij}$ are obtained in different ways depending on the method used: FBS or GF.
For a fixed $\kappa_{n}$, the spectrum of the radial Dirac equation is non-degenerate, so it is convenient to introduce the index $i_{n}$ that enumerates solutions for the given angular symmetry:
$\begin{cases}\left(-\dfrac{d}{dr}+\dfrac{\kappa_{n}-1}{r}\right)f_{{i_{n}}{\kappa_{n}}}+\left(V+m_{e}\right)g_{{i_{n}}{\kappa_{n}}}=\varepsilon_{{i_{n}}{\kappa_{n}}}g_{{i_{n}}{\kappa_{n}}}\,,\\ \left(\dfrac{d}{dr}+\dfrac{\kappa_{n}+1}{r}\right)g_{{i_{n}}{\kappa_{n}}}+\left(V-m_{e}\right)f_{{i_{n}},{\kappa_{n}}}=\varepsilon_{{i_{n}}{\kappa_{n}}}f_{{i_{n}}{\kappa_{n}}}\,.\end{cases}$ (25)
Then, Eq. (7) results for, e.g., $G^{11}$ in
$G_{\kappa_{n}}^{11}(\omega,{r_{1},r_{2}})=\sum_{i_{n}}\frac{g_{{i_{n}}{\kappa_{n}}}(r_{1})g_{{i_{n}}{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{{i_{n}}{\kappa_{n}}}^{-}}.$ (26)
We stress that the index $i_{n}$ in Eq. (26) and similar expressions runs over both the positive- and negative-energy Dirac continua as well as all bound states for the chosen $\kappa_{n}$. When using the DKB approach, for each $\kappa_{n}$ the spectrum of the radial Dirac equation (25) is replaced with the finite set of solutions. In this case, the index $i_{n}$ runs over these finite sets. When using the GF approach, the radial Green’s function $G_{\kappa_{n}}^{ij}$ can be constructed from the solutions of the homogeneous Dirac equation, which are bounded at zero, $\phi_{\kappa_{n}}^{0}=(g_{\kappa_{n}}^{0}\;\;f_{\kappa_{n}}^{0})^{\text{{T}}}$, and at infinity, $\phi_{\kappa_{n}}^{\infty}=(g_{\kappa_{n}}^{\infty}\;\;f_{\kappa_{n}}^{\infty})^{\text{{T}}}$, (see, e.g., Ref. Yerokhin and Maiorova (2020)):
$G_{{\kappa_{n}}}(\omega,r_{1},r_{2})=\phi_{{\kappa_{n}}}^{\infty}(r_{1}){\phi_{{\kappa_{n}}}^{0}}^{\text{{T}}}(r_{2})\theta(r_{1}-r_{2})+\phi_{{\kappa_{n}}}^{0}(r_{1}){\phi_{{\kappa_{n}}}^{\infty}}^{\text{{T}}}(r_{2})\theta(r_{2}-r_{1}).$ (27)
Note that the solutions $\phi^{0}$ and $\phi^{\infty}$ are supposed here to be normalized so that their Wronskian equals one. Expressions similar to Eqs. (24)-(27) can also be obtained for the free-electron Green’s function, which we will denote as $G^{(0)}$.
With this in mind, we construct the many-potential part of the bound-electron Green’s function which involves two or more interactions with the binding potential and corresponds to the last term in the expression (11):
$G^{(2+)}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\int d^{3}\mathbf{x}\int d^{3}\mathbf{y}G^{(0)}(\omega,\mathbf{r}_{1},\mathbf{x})V(x)G(\omega,\mathbf{x},\mathbf{y})V(y)G^{(0)}(\omega,\mathbf{y},\mathbf{r}_{2}).$ (28)
Since the potential $V$ is assumed to be spherically symmetric, angular integrations in Eq. (28) can be easily carried out. As a result, one obtains the same PW expansion for $G^{(2+)}$ as that for $G$ in Eq. (24). The radial components of $G^{(2+)}$ read as
$G^{(2+)ij}_{\kappa_{n}}(\omega,r_{1},r_{2})=\sum_{k,m}\int dxx^{2}\int dyy^{2}G^{(0)ik}_{\kappa_{n}}(\omega,r_{1},x)V(x)G^{km}_{\kappa_{n}}(\omega,x,y)V(y)G^{(0)mj}_{\kappa_{n}}(\omega,y,r_{2}).$ (29)
In the GF case, because of the $\theta$-functions, radial integrations need to be handled carefully in Eq. (29), see, e.g., Ref. Artemyev et al. (2013). In the framework of the DKB approach, it is convenient to rewrite $G^{(2+)}$ in the form:
$G^{(2+)}(\omega)=\sum_{n}\frac{|\check{n}\rangle\langle\check{n}|}{\omega-\varepsilon_{n}^{-}},$ (30)
where the new states $|\check{n}\rangle$ are defined by:
$|\check{n}\rangle=\sum_{f}\frac{|f\rangle\langle f|V|n\rangle}{\omega-\varepsilon_{f}^{-}}\equiv{\begin{pmatrix}\check{g}_{i_{n}\kappa_{n}}(r)\chi_{\kappa_{n}m_{n}}(\mathbf{\hat{r}})\\ i\check{f}_{i_{n}\kappa_{n}}(r)\chi_{-\kappa_{n}m_{n}}(\mathbf{\hat{r}})\end{pmatrix}}$. (31)
Then, using Eq. (31), the many-potential contribution is expressed as:
$\Delta\varepsilon_{a}^{\text{Mp}}=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle a\check{n}|I(\omega)|\check{n}a\rangle}{\varepsilon_{a}-\omega-\varepsilon_{n}^{-}}.$ (32)
The expression (32) can be readily adjusted to be used within the GF approach. For this aim, one needs to reverse an expression similar to Eq. (26):
$\sum_{i_{n}}\frac{\check{g}_{{i_{n}}{\kappa_{n}}}(r_{1})\check{g}_{{i_{n}}{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}\rightarrow G^{(2+)11}_{\kappa_{n}}(\omega,r_{1},r_{2}).$ (33)
The matrix element of the operator $I(\omega)$ can be expressed as follows:
[…]
The sums over $J$ and $\kappa_{n}$ in Eq. (35) are not independent due to the triangular inequality. In our calculations, the sum over $\kappa_{n}$ is considered as the primary one. As a result, the many-potential contribution is represented by the sum of the often poorly converging PW-expansion series in $\kappa_{n}$. To calculate it, one has to resort to the procedure of extrapolation in $k=|\kappa_{n}|$. In what follows, we for brevity will use $\kappa$ instead of $\kappa_{n}$, if this does not lead to misunderstandings.
II.4 Partial-wave expansion acceleration schemes
The convergence of the PW-expansion series could be improved using different acceleration schemes. Their general idea is to separate out and subtract the term $\Delta\varepsilon_{a,x}^{\text{subtr}}$, which contains the slowly converging part of the PW expansion of the many-potential contribution. The index ‘‘$x$’’ in $\Delta\varepsilon_{a,x}^{\text{subtr}}$ indicates that the subtracted term should be calculated exactly in the same manner as $\Delta\varepsilon_{a}^{\text{Mp}}$, that is in coordinate space using the PW expansion. Naturally that the same term, evaluated separately, but not necessarily within the same approach, so we denote it as $\Delta\varepsilon_{a}^{\text{subtr}}$ without the additional index ‘‘$x$’’, should then be added back. The discussed idea can be illustrated by the formula:
$\Delta\varepsilon_{a}^{\text{Mp}}\rightarrow(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{\text{subtr}})+\Delta\varepsilon_{a}^{\text{subtr}}.$ (36)
It is implied that the PW-expansion series for the two contributions in the brackets in Eq. (36) are subtracted term by term and only then extrapolated in $k$. The resulting difference is expected to demonstrate a better convergence (at least, in terms of absolute values) than the initial series for the many-potential contribution. Therefore, if one can calculate $\Delta\varepsilon_{a}^{\text{subtr}}$ with high accuracy, it will determine the overall improvement in the calculations of the many-potential contribution to the SE correction.
Within the two-potential acceleration scheme (see the Introduction section), the subtracted term is the two-potential contribution, $\Delta\varepsilon_{a,x}^{\text{subtr}}=\Delta\varepsilon_{a,x}^{2\text{p}}$. When adding this term back, one must evaluate it with high precision in order to benefit from applying the scheme. A natural idea is to perform the corresponding calculations in momentum space without using the PW expansions. However, this problem is complicated and leads to complex multidimensional integrals, see, e.g., Ref. Yerokhin and Jentschura (2010b), where a similar approach was employed in the point-nucleus case for the SE correction to the hyperfine splitting. For this reason, an alternative was proposed in Refs. Artemyev et al. (2007, 2013), where the difference between the many- and two-potential terms was addressed using the FBS method, while the two-potential term itself was also considered in coordinate space but employing the GF method and truncating the corresponding PW summations at higher values of angular momenta. We note that in this case, it is not necessary to solve the system of differential equations numerically for an arbitrary potential, since the PW expansion of the free-electron propagator can be expressed in terms of the spherical Bessel functions Mohr et al. (1998) and the desired two-potential term can be obtained by combining three such propagators. In the present work, we study this two-potential scheme but for simplicity the difference between the many- and two-potential terms is also treated based on the GF approach. The PW expansion for this difference arises from the Green’s function
$G^{(3+)}=G^{(2+)}-\sum_{f_{1},f_{2},f_{3}}\frac{|f_{1}\rangle\langle f_{1}|V|f_{2}\rangle\langle f_{2}|V|f_{3}\rangle\langle f_{3}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{f_{2}}^{-})(\omega-\varepsilon_{f_{3}}^{-})}\,.$ (37)
We refer to the term corresponding to $G^{(3+)}$ as the three-plus-potential contribution, $\Delta\varepsilon_{a}^{\text{(3+)p}}\equiv(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{2\text{p}})$.
Figure 3: Diagrammatic representation of the quasi-two-potential term separation. The last term on the right side is calculated in momentum space without using any PW expansion.
The Sapirstein-Cheng (SC) scheme Sapirstein and Cheng (2023) is based on the idea that the dominant contribution from the free-electron propagators $G^{(0)}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ comes from the region where both their spatial arguments are close to each other Mohr (1974b), $\mathbf{r}_{1}\approx\mathbf{r}_{2}$. Therefore, one can obtain an approximation to the two-potential term by ‘‘transposing’’ the potentials from the inner electron line to the vertices, where the photon propagator is attached, see Fig. 3. We refer to this approximation as the quasi-two-potential term and denote it as $\Delta\varepsilon_{a,x}^{\text{subtr}}=\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$. The difference between the many- and quasi-two-potential terms, $\Delta\tilde{\varepsilon}_{a,x}^{(3+)\text{p}}\equiv\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$, referred to as the quasi-three-plus-potential contribution, can be studied in the coordinate space using the PW expansion for the Green’s function
$\widetilde{G}^{(3+)}=G^{(2+)}-\sum_{f}\frac{|Vf\rangle\langle fV|}{(\varepsilon_{a}-\omega-\varepsilon_{f}^{-})^{3}}$ (38)
Here, $|Vf\rangle=V|f\rangle$. A solid advantage of the SC scheme is the fact that it is fairly easy to calculate the quasi-two-potential contribution with high accuracy in momentum space. We denote the result as $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ and use it as $\Delta\varepsilon_{a}^{\text{subtr}}$. The potentials that have been transposed into vertices can be formally taken into account by multiplying the wave function of the external state $|a\rangle$ by them. An issue arises of calculating the Fourier transform of such products. Since $V$ is assumed to be spherically symmetric, this can be done by a straightforward generalization of Eqs. (13)-(15). We represent the result in the following form:
$\phi_{a}(\mathbf{p})=\int d^{3}{\mathbf{r}}e^{-i({\mathbf{p\cdot r}})}V(r)\psi_{a}({\mathbf{r}})=i^{-l_{a}}{\begin{pmatrix}\widetilde{t}_{a}(p)\chi_{\kappa_{a}m_{a}}(\mathbf{\hat{p}})\\ \widetilde{s}_{a}(p)\chi_{-\kappa_{a}m_{a}}(\mathbf{\hat{p}})\end{pmatrix}}$, (39)
$\widetilde{t}_{a}(p)=4\pi\int_{0}^{\infty}drr^{2}j_{l_{a}}(pr)V(r)g_{a}(r),$ (40)
$\widetilde{s}_{a}(p)=-4\pi\frac{\kappa_{a}}{|\kappa_{a}|}\int_{0}^{\infty}drr^{2}j_{2j_{a}-l_{a}}(pr)V(r)f_{a}(r).$ (41)
Using the identity
$\int d\mathbf{y}d\mathbf{z}G^{(0)}(\omega,\mathbf{x_{1}},\mathbf{y})G^{(0)}(\omega,\mathbf{y},\mathbf{z})G^{(0)}(\omega,\mathbf{z},\mathbf{x_{2}})=\frac{1}{2}\frac{\partial^{2}}{\partial\omega^{2}}G^{(0)}(\omega,\mathbf{x_{1}},\mathbf{x_{2}}),$ (42)
one obtains
$\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}=\frac{1}{2}\int\frac{d^{3}\mathbf{p}}{(2\pi)^{3}}\bar{\phi}_{a}(\mathbf{p})\frac{\partial^{2}\Sigma_{R}^{(0)}(\mathrm{p})}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\phi_{a}(\mathbf{p}),$ (43)
where $\mathrm{p}=(p_{0},\mathbf{p})$. The explicit form of the derivatives of the free-electron self-energy operator $\Sigma_{R}^{(0)}$ with respect to $p_{0}$ in both gauges is given in Appendix D. After performing the angular integration, which is carried out similarly to that in the zero-potential contribution, one can obtain:
$\begin{gathered}\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}=\frac{\alpha}{8\pi}\int\frac{dpp^{2}}{(2\pi)^{3}}\left[N_{1}\left(\widetilde{t}_{a}^{2}-\widetilde{s}_{a}^{2}\right)+N_{2}\left(\widetilde{t}_{a}^{2}+\widetilde{s}_{a}^{2}\right)+2pN_{3}\widetilde{t}_{a}\widetilde{s}_{a}\right]\end{gathered}$ (44)
where $N_{1}$, $N_{2}$, and $N_{3}$ are defined in Appendix D, $\widetilde{t}_{a}=\widetilde{t}_{a}(p)$ and $\widetilde{s}_{a}=\widetilde{s}_{a}(p)$.
III Computational details
The calculations are carried out by means of two approaches: the FBS method within the DKB basis and the GF method. The methods differ only in the treatment of the terms evaluated in coordinate space. The momentum-space contributions are the same in both approaches.
III.1 Momentum-space calculations
The contributions evaluated in momentum space include the zero- and one-potential terms resulting from the renormalization procedure, as well as the quasi-two-potential terms corresponding to the SC convergence-acceleration scheme. All the integrals in the zero-, one-, and quasi-two-potential terms are calculated using the Gauss-Legendre quadratures. Let us first briefly discuss an evaluation of the zero-potential terms. The domain of integration over $p$ is divided into two parts, $p\leq p_{0}$ and $p\geq p_{0}$, where $p_{0}=\alpha Z$. The integration over the second domain can be carried out using a change of variables of the form $t=\alpha Z/p$, as a result of which the ray $\alpha Z,\infty)$, is mapped onto the interval $[1,0]$. The first domain is finite, so its treatment causes no problems. Compared to the Feynman gauge, the zero-potential contribution in the Coulomb gauge contains an additional integration over the Feynman parameter $x$, see Eqs. (17), (60), and (61). The coefficient $F_{2}$ in Eq. (61) must be calculated as a principal-value integral. From a numerical point of view, it is helpful to isolate the singularity. We do that by adding and subtracting the same integrand but with the numerator evaluated at the singularity point $x=a$:
$\int dx\frac{F(x)}{x-a+i0}=\int dx\frac{F(x)-F(a)}{x-a+i0}+\int dx\frac{F(a)}{x-a+i0}.$ (45)
The first term in the right-hand side of Eq. (45) becomes regular. To avoid numerical problems at $x=a$, the exact expression can be replaced with its Taylor series in a small vicinity of this point. The second term in Eq. (45) is calculated analytically. Similar methods can be used to calculate $\Delta\tilde{\varepsilon}_{a}^{2\text{p}}$ in momentum space. To avoid possible numerical problems due to large cancellations, one should use the Taylor-series expansion for the corresponding integrands.
The evaluation of the one-potential contributions involves the integrations over the variables $p$, $p^{\prime}$, $z$, and a number of Feynman parameters. When integrating over the Feynman parameters in both gauges, to avoid possible numerical errors, the Taylor series for the integrands are employed. In the Coulomb-gauge case we use the results of Ref. Holmberg (2011), but expand all the coefficients $F_{1}$ - $F_{21}$ at ‘‘suspicious’’ points. The expression (20) has an integrable singularity for $q=0$ when $p=p^{\prime}$ and $z=1$. Therefore it is necessary to properly choose integration nodes for the variable $z$ to ensure the integral convergence. For this goal, we use the transformation $z=1-2t^{2}$ with $t\in[0,]$. To calculate the integrals over the variables $p$ and $p^{\prime}$, we change the variables according to $P=(p+p^{\prime})/\sqrt{2}$ and $P^{\prime}=(p-p^{\prime})/\sqrt{2}$. The integral over $P$ is calculated similar to the integral over $p$ for the zero-potential contribution, but the separation point $p_{0}=2\alpha Z$ is chosen instead. The integral over $P^{\prime}$ is taken within the finite interval [P,P][-P,P].
To estimate an uncertainty of the momentum-space calculations, the number of quadrature nodes and position of the separation points $p_{0}$ are varied.
III.2 Coordinate-space calculations
The calculation of all coordinate-space contributions, many-potential, two-potential, and (quasi-) three-plus-potential terms, has many common points. Let us illustrate them with an example of the many-potential contribution. Its evaluation essentially reduces to the calculation of three integrals: over the radial variables $r_{1}$ and $r_{2}$, see Eqs. (75) and (76), and over the energy parameter $\omega$. Taking into account the exponential decay of bound-electron wave functions at large distances, all radial integrations are performed over the finite interval $0,R_{\text{max}}]$. The value of the parameter $R_{\text{max}}$ is chosen so that the results are independent of it within the desired accuracy. All the radial integrations are carried out using the Gauss–Legendre quadratures. However, the specific details differ slightly in the DKB and GF approaches.
In the DKB approach, the domain $0,R_{\text{max}}]$ is divided into a finite number of intervals by the nodes used to construct the B splines. The integration method for the double integral over $r_{1}$ and $r_{2}$ depends on whether these variables belong to the same interval or not. If not, the integrations are carried out independently using the same quadrature for both variables. However, when $r_{1}$ and $r_{2}$ fall into the same interval, a more delicate integration is employed to take into account the fact that the integrand depends on $r_{<}=\min\{r_{1},r_{2}\}$ and $r_{>}=\max\{r_{1},r_{2}\}$. Namely, for some fixed value of $r_{1}$ the corresponding interval is divided into two subitervals lying to the left and to the right from $r_{1}$. The integration over $r_{2}$ is performed by using the Gauss-Legendre quadratures for both subintervals. Note that within the DKB approach, the integration over radial variables turns out to be ‘‘tied’’ to the size of the basis $N$. As the FBS size grows, the number of nodes also increases, that leads to the improvement of the integration accuracy. Thus, the single parameter $N$ affects both the accuracy of the Green’s function representation and the integration accuracy.
In the GF approach, the radial integrations become more complicated due to the aforementioned discontinuities of the electron Green’s functions. The integrations over $x$ and $y$ in Eq. (28) cannot be performed independently of the integrations over $r_{1}$ and $r_{2}$. To evaluate the radial integrals, we employ a variation of the method described, e.g., in Ref. Yerokhin and Maiorova (2020). Namely, a hierarchy of the radial integration grids is prepared, with each subsequent grid becoming finer and ‘‘lying’’ inside the previous one. Then, all the radial integrations are arranged in the following order: $r_{1}$, $r_{2}$, $x$, and $y$. Therefore, the finest grid corresponds to the variable $y$. The accuracy of the GF approach is determined mainly by the accuracy of the radial integrations.
$\text{Re}(\omega)$$\text{Im}(\omega)$$\varepsilon_{a}-\varepsilon_{1s}$
Figure 4: Deformed contour (blue solid line) in the complex plane $\omega$. Singularities and branching cuts of the electron Green’s function (black lines and black dots) and photon propagator (black dashed lines) are shown.
The photon and electron propagators have a complicated analytical structure as functions of the energy parameter $\omega$ in the complex plane. They are two-valued functions and are defined on the corresponding Riemann surfaces with two sheets (for each propagator). Fig. 4 shows the complex plane of the integration variable $\omega$, the branch points of both propagators with the cuts starting from them, and the poles of the electron propagator corresponding to the bound states. In the expression for the SE correction (3), the integration contour goes along the real axis, but it can be bended in the $\omega$ complex plane to ‘‘simplify’’ the calculations. Because of the zero photon mass, the contour is squeezed between two infinitely close branching points near zero. Introducing a finite mass would allow them to be pushed apart, which can be easily demonstrated in the Feynman gauge. Consequently, the contour must always pass through $\omega=0$. In principle, there is no other restrictions on the contour. Thus, it can be changed as needed, respecting the analytical structure of the integrand.
In this work, the approach proposed in Ref. Yerokhin and Shabaev (1999) is used. Namely, we employ the transformation that results in the contour shown in Fig. 4. The original contour along the real axis is rotated and deformed. The contributions from the circular arcs vanish as the radius of the circle tends to infinity, so we do not consider them. The rotation is done to avoid fast oscillations of the integrand encountered on the real axis: on the vertical parts of the deformed contour the integrand decays exponentially. The deformation of the contour near zero and its shift to the right from the imaginary axis is done to avoid going close to the poles of the electron Green’s function, which correspond to the bound states with binding energies no less than that of the state under consideration. For large enough values of $\kappa$, the contour simplifies and can be taken along the imaginary axis. The integration over $\omega$ is evaluated using the Simpson quadratures. The number of integration points is increased until the required accuracy is achieved.
III.3 Extrapolation of the PW-expansion sums
We obtain a coordinate-space contribution as a PW series in $\kappa$, which should inevitably be truncated at some $|\kappa_{\text{max}}|$. The remainder of the series can then be estimated using an extrapolation procedure. The extrapolation methods discussed below are applied to all treated coordinate-space contributions. Here and in what follows, when discussing partial sums, we will denote by the index $k>0$ the sum of all terms of the PW expansion with $|\kappa|\leq k$.
When using the DKB method, the individual contributions strongly depend on the FBS size $N$, therefore, an extrapolation over this parameter is also required. In this case, the most stable results are obtained if one extrapolates with respect to the basis size first. Let us denote as $S_{k}(N)$ the partial sum obtained for the given FBS size $N$. For each $k$, the extrapolation over $N$ is performed using a non-linear least-squares method using an ansatz of the form:
$S_{k}(N)=A_{k}\left(\frac{1}{N}\right)^{B_{k}}+S_{k},$ (46)
where $S_{k}$ are the extrapolated to the infinite basis set partial sums, and $A_{k}$ and $B_{k}$ are adjustable numerical coefficients. In the GF approach, the sums $S_{k}$ are obtained directly from the calculations, and this extrapolation is not required.
To extrapolate over $k$ in both DKB and GF approaches, we implement three following ansatzes for the partial sums $S_{k}$:
$\tilde{S}(k)=S_{\infty}+\frac{C_{2}}{k^{2}}+\frac{C_{3}}{k^{3}}+...+\frac{C_{m}}{k^{m}}\penalty 10000\ ,$ (47)
where the parameter $S_{\infty}$ and coefficients $C_{i}$ for $i=2,\ldots,m$ (with $m$ being $4$, 55 and $6$) are obtained by a kind of statistical method. Namely, for each $m$, $N_{\text{iter}}$ times a subset of $k$’s, $\{k_{i}\}_{i=1}^{m}$, is randomly selected. The parameter $N_{\text{iter}}$ is fixed by the condition that all possible subsets are considered for the largest $m$ ($m=6$). For each element of the above subsets we construct and solve the system of $m$ linear equations of the form
$\tilde{S}(k_{i})=S_{k_{i}},\quad i=1,\dots,m$ (48)
to determine $m$ variables $S_{\infty}$, $C_{2}$, …, $C_{m}$. Repeating this step $3N_{\text{iter}}$ times yields a set of approximations for $S_{\infty}$. Evaluating the mean value $\overline{S}_{\infty}$ and variance $\sigma_{k}$, one obtains the desired extrapolated sum and an estimation for its uncertainty. We note that we approximate the tails only (by excluding the first few $k$’s from consideration). Changing the tail size produces results that are consistent with each other.
Within the DKB approach, the uncertainties of extrapolation in $k$, $\sigma_{k}$, and extrapolation to an infinite basis set, $\sigma_{N}$, are summed quadratically
$\sigma_{\text{tot}}=\sqrt{\sigma_{\kappa}^{2}+\sigma_{N}^{2}}.$ (49)
Example of the use of this two-step extrapolation procedure can be found in the ‘‘results’’ section below.
IV Results
In the present work, we conduct two sets of SE-correction calculations. First, we consider the $1S_{1/2}$ state in three hydrogen-like ions: argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). We adopt the results from Ref. Hedendahl and Holmberg (2012) as a benchmark for comparison. For this reason, in the first part of calculations, the values of root-mean-square (rms) nuclear radii as well as the values of fundamental constants, e.g., the fine-structure constant $\alpha$ and the Hartree energy in eV, are taken to be the same as there. The obtained data are used to analyze the convergence of the many-potential contribution. Second, we study the SE corrections for the $1S_{1/2}$, $2S_{1/2}$, $2P_{1/2}$, and $2P_{3/2}$ states in a set of hydrogen-like ions with the nuclear charge $Z$ in the range from $Z=10$ to $Z=100$ and compare the obtained results with the ones presented in Ref. Yerokhin and Shabaev (2015a). In this case, the rms radii and models of the nuclear-charge distribution are chosen to coincide with those in the corresponding work, however, for $\alpha$ the numerical value from the CODATA (2018) Tiesinga et al. (2021) is used. In the second part, the SE correction is represented in terms of the dimensionless function $F(\alpha Z)$, defined as:
$\Delta\varepsilon_{a}=\frac{\alpha}{\pi}\frac{(\alpha Z)^{4}}{n^{3}}F(\alpha Z)mc^{2}.$ (50)
The data obtained in the second part of the work are used to analyze the acceleration-convergence schemes.
IV.1 Many-potential term: convergence analysis
Let us start with illustrating the two-step extrapolation procedure, described in the previous section, using the ground state of hydrogen-like xenon ($Z=54$) as an example, see Fig. 5. The many-potential term is considered, and the results for both gauges are presented. At the first stage, the extrapolation of the partial sums $S_{k}(N)$ with respect to the size of the basis $N$ is performed. For convenience, only the extrapolation curves for $k=10$, $15$, and $20$ are shown as functions of $1/N$. The corresponding extrapolated sums $S_{10}$, $S_{15}$, and $S_{20}$ are depicted as stars at $1/N\rightarrow 0$. The extrapolated partial sums $S_{k}$ obtained for other values of $k$ are depicted with dashes as well. At the second stage, the extrapolation over $k$ is performed. The final values for $\Delta\varepsilon_{a}^{\text{Mp}}=S_{\infty}$ are shown with black squares. Note that the GF method yields results that are indistinguishable at this scale from those obtained by the DKB method. Based on our experience, a small difference in the FBS size extrapolation in the two gauges appears in the small-$Z$ region, where the convergence behavior is slightly better in the Coulomb gauge. In the region of medium- and large-$Z$ there is no significant difference between the gauges. This can be seen, e.g., from Fig. 5, where the absolute values of the many-potential terms in two gauges differ by approximately one order of magnitude, while the rate of convergence in the FBS size is approximately the same.
Figure 5: The extrapolation of the many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like xenon ($Z=54$) obtained within the framework of the DKB approach in the Feynman (a) and Coulomb (b) gauges.
Figure 6: The individual many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}(k)$: (a) normalized to the total self-energy correction $\Delta\varepsilon_{a}$ and (b) additionally multiplied by $k^{3}$. Note the log-log scale.
Let us now examine in details the PW convergence of the many-potential contributions and perform a comparative analysis within two gauges. The zero-, one-, and many-potential contributions to the SE correction for the $1S_{1/2}$ state in hydrogen-like argon ($Z=18$) and uranium ($Z=92$) in Feynman and Coulomb gauges are presented in Table 1. For the many-potential contributions, the individual terms of the PW expansion, $\Delta\varepsilon_{a}^{\text{Mp}}(k)$, obtained by means of the GF approach are shown. These terms represent sums of contributions with $\kappa=k$ and $\kappa=-k$, i.e., the partial sums are given by $S_{k}=\sum_{i=1}^{k}\Delta\varepsilon_{a}^{\text{Mp}}(i)$. The row labeled $\sum_{k=1}^{24}$’’ gives the partial sums $S_{24}$, while the subsequent row $\sum_{k=25}^{\infty}$ [extr.]’’ provides our estimations for the remainders of PW series obtained by applying the statistical extrapolation procedure discussed above. The total SE corrections are presented in the last line.
For convenience, the absolute values of the individual terms $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ from Table 1, normalized to the total self-energy correction $\Delta\varepsilon_{a}$, are plotted in Fig. 6(a). From Table 1 and Fig. 6(a), it can be noticed that for both considered $Z$ the many-potential contributions as well as the first terms of the PW expansions are significantly smaller in the Coulomb gauge (in the case of $Z=18$, by two orders of magnitude) than in the Feynman gauge. This difference, however, becomes much less noticeable for large values of $k$, although the Feynman-gauge contributions still exceed their Coulomb-gauge counterparts.
Multiplying the individual terms $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ obtained for two gauges by the factor of $k^{3}$ provides a more informative comparison of their decay rate with the growth of $k$, that is demonstrated in Fig. 6(b). Let us discuss the leading asymptotic behavior of the PW-expansion terms for $k\rightarrow\infty$: $\Delta\varepsilon_{a}^{\text{Mp}}(k)\sim 1/{k^{p}}$. Our numerical results indicate that in both the Feynman and Coulomb gauges, the exponent $p$ exceeds 2 over a wide range of $Z$, with $p$ being close to $3$ in the large-$Z$ region. These conclusions do not contradict those from Ref. Yerokhin et al. (2005), where it was claimed that in the Feynman gauge $p=3$ regardless of the nuclear charge $Z$. The latter observation justifies the fact that the anzatz for the extrapolation over $k$ in (47) starts from ${1}/{k^{2}}$, since $\sum_{i=1}^{k}{1}/{i^{3}}\sim{1}/{k^{2}}$.
We also note that the numerical coefficients for the higher powers of $1/k$ in the expansion (47) in the Coulomb gauge are larger than their Feynman-gauge counterparts. That fact is also illustrated in the Fig. 6(b) – convergence to asymptotics in the Feynman gauge is significantly slower. As a result, within the $k$-range studied in the present work ($k\leq 20$ or $k\leq 24$), the leading asymptotic behavior is less pronounced in the Coulomb gauge. The latter fact apparently balances out the smallness of the absolute value of the Coulomb-gauge many-potential contribution. In practice, within the GF approach, given the same resource costs (number of partial waves, number of integration nodes for radial variables, etc.), the Coulomb and Feynman gauges are equally suitable for estimating the many-potential contribution to the SE correction within a wide range of nuclear charges $Z$. This conclusion slightly contradicts what was claimed in Ref. Yerokhin et al. (2025a). Nevertheless, we agree with the fact that the advantages of the Coulomb gauge are the improved convergence of the many-potential contributions over the basis-set size in the small-$Z$ region, relative smallness of the many-potential contribution Hedendahl and Holmberg (2012); Yerokhin et al. (2025a), and the absence of cancellation between different parts of the SE correction.
IV.2 Application of the convergence-acceleration schemes
Now let us turn to the comparative analysis of the SC and two-potential schemes used to accelerate the convergence of the PW expansions for the many-potential contributions in the Feynman and Coulomb gauges. In Fig. 7(a), ratios of the three-plus-potential $\Delta\varepsilon_{a,x}^{\text{(3+)p}}$ and quasi-three-plus-potential $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ contributions to the total many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}$ are plotted. In the low-$Z$ region, the two-potential scheme in Coulomb gauge yields the smallest remainder which is to be extrapolated. However, in this case the overall accuracy is limited by the two-potential contribution, which also has to be treated within the PW expansion.
Figure 7: Comparison of two convergence-acceleration schemes in two gauges, Feynman and Coulomb ones: (a) three-plus-potential $\Delta\varepsilon_{a,x}^{\text{(3+)p}}$ and quasi-three-plus-potential $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ contributions normalized to the total many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}$; (b) the same for the coefficients $C_{2}$ defined in Eq. (47).
Let us turn to the coefficients $C_{2}$ from Eq. (47) obtained for $\Delta{\varepsilon}_{a,x}^{\text{(3+)p}}$ and $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ (designated as $C_{2}^{\text{(3+)p}}$) by means of our statistical extrapolation approach. These coefficients normalized to the coefficients $C_{2}^{\text{Mp}}$ for the corresponding many-potential contributions are shown in Fig. 7(b). These data partially explain the convergence-acceleration mechanism. In the area of low $Z$, the coefficients $C_{2}$ for the (quasi-)three-plus-potential contributions decrease by one or two orders of magnitude compared to their counterparts for the many-potential contributions. In the region of high $Z$, they also decrease, but less effectively. It should be noted that for each gauge the reduction of the coefficient $C_{2}$ for both studied acceleration schemes is almost identical, which confirms that the SC trick indeed provide a good approximation to the two-potential contribution.
Figure 8: The individual contributions $\Delta\varepsilon_{a}^{\text{(3+)p}}(k)$ and $\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}(k)$ normalized to the total SE correction and multiplied by the factor of $k^{3}$: (a) for $Z=18$ argon; (b) for $Z=92$ uranium.
The subtraction of the two-potential or quasi-two-potential contributions from the many-potential contribution apparently does not change the leading asymptotic behavior of the PW-expansion terms for the remainders $\Delta\varepsilon^{\text{(3+)p}}_{a}$ and $\Delta\tilde{\varepsilon}^{\text{(3+)p}}_{a}$ and only decreases the absolute values of the corresponding coefficients. To illustrate this, in Fig. 8 we plot the individual terms of the PW expansions for these remainders normalized to the total SE correction and additionally multiplied by $k^{3}$, exactly as it was done in Fig. 6 for the many-potential contributions. One can see that the individual terms tend to constants in all the cases, i.e. $p=3$ for (quasi-)three-plus potential terms. That statement has also been verified empirically: the extrapolation procedure gives obviously incorrect results if one omits the term $C_{2}/k^{2}$ in the Eq. (47). We stress that our conclusion contradicts the one from work Sapirstein and Cheng (2023). Fig. 8 also demonstrates that for both acceleration schemes the approach to the asymptotics in the Coulomb gauge is much faster than in the Feynman gauge. The last statement makes it easier to extrapolate the corresponding contributions and certainly makes the Coulomb gauge more preferable for SE calculations within the convergence-acceleration schemes for a wide range of $Z$.
IV.3 Data: summary
In this section, we present and discuss the results of our calculations of SE corrections to the $n=1$ and $n=2$ energy levels in hydrogen-like ions.
First, in Table 2 we study the SE correction for the ground $1S_{1/2}$ state of hydrogen-like argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). As was noted above, Ref. Hedendahl and Holmberg (2012) is considered to be the benchmark in this case. Therefore, in this part of the calculations, we adopt the values of the fundamental constants and nuclear parameters to be the same as those used in that work. For each ion, we consider two gauges, the Coulomb gauge and the Feynman gauge, which are indicated in the second column as ‘‘C’’ and ‘‘F’’, respectively. The corresponding zero- and one-potential contributions, calculated in momentum space, are shown in the third and forth columns. For each gauge we compare different numerical approaches to the evaluation of the the many-potential contribution $\Delta\varepsilon_{a}^{\text{Mp}}$: (i) the standard method, labeled in the column ‘‘Scheme’’ as ‘‘DIR’’, in which the many-potential contribution is treated as a whole; (ii) the two-potential acceleration scheme, labeled as ‘‘TP’’, where this contribution is represented as the sum $\Delta\varepsilon_{a}^{\text{2p}}+\Delta\varepsilon_{a}^{\text{(3+)p}}$, and (iii) the Sapirstein-Cheng acceleration scheme, labeled as ‘‘SC’’, where it is given by the sum $\Delta\tilde{\varepsilon}_{a}^{\text{2p}}+\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}$. We also compare two methods for representing the electron propagators: the finite-basis-set method within the DKB basis and the Green’s function approach, which are indicated as ‘‘DKB’’ and ‘‘GF’’, respectively. All the coordinate-space contributions, $\Delta\varepsilon_{a}^{\text{Mp}}$, $\Delta\varepsilon_{a}^{\text{2p}}$, $\Delta\varepsilon_{a}^{\text{(3+)p}}$, and $\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}$, are obtained by applying the statistical extrapolation procedure described in Sec. III.3. Within the DKB approach, the PW-expansion series are truncated at $|\kappa_{\text{max}}|=20$. When using the GF approach, these series are truncated at $|\kappa_{\text{max}}|\sim 45$ in the case of the two-potential contribution $\Delta\varepsilon_{a}^{\text{2p}}$ and at $|\kappa_{\text{max}}|=24$ in the other cases. Finally, the total SE corrections obtained by means of different approaches are given in the last column. For each ion and gauge, the results from Ref. Hedendahl and Holmberg (2012) are shown for comparison as well.
As can be seen from Table 2, the main source of numerical errors, regardless of the gauge, is the many-potential contribution. Even when applying the acceleration schemes, this contributions still limits the accuracy of the total SE corrections. We note that the many-potential contribution is significantly smaller in the Coulomb gauge than in the Feynman gauge (see the penultimate column in Table 2), which was also noticed in Refs. Hedendahl and Holmberg (2012); Yerokhin et al. (2025a). However, for the reasons described above, this does not provide a significant increase in accuracy unless acceleration schemes are involved. Although obtaining the zero- and one-potential contributions with sufficient accuracy is not a technically difficult task, an attractive feature of the Coulomb gauge, as can be seen from Table 2, is the absence of large cancellations between these momentum-space contributions. Nevertheless, this observation is extremely useful for obtaining high-precision results for low-$Z$ systems.
Clearly, acceleration tricks reduce extrapolation uncertainties and allow for a more accurate evaluation of the many-potential contributions. Using either the two-potential scheme or the SC scheme reduces the numerical errors and allows one to obtain up to two additional significant digits. As a rule, the most accurate many-potential contribution can be obtained by using a combination of the Coulomb gauge and the SC acceleration scheme.
The GF approach provides a more accurate treatment of the many-potential contribution compared to the DKB approach. This results mainly not from the lower value of $|\kappa_{\text{max}}|$ employed in this case, but from the fact that the DKB method needs the approximation with respect to the size of the basis set. The low accuracy of the DKB approach does not detract from the already mentioned advantages, namely, the ability to exclude a specific state and construct Green’s functions for a wider class of systems that do not possess spherical symmetry. In some sense, the FBS method (not necessarily in the DKB realization) is more flexible and easier to implement. Therefore, if high accuracy is not required, it is an excellent method for calculating the SE correction.
Returning to the comparison with the benchmark calculations Hedendahl and Holmberg (2012), we note that our results obtained by means of different methods are in good agreement with each other and those from Ref. Hedendahl and Holmberg (2012). Small deviations have been revealed. However, the reason for them is unclear for us.
For a broader comparison, another set of calculations has been performed. Namely, we consider the SE correction for the ground $1S_{1/2}$ state in hydrogen-like ions with $Z\in\{10,18,26,36,54,82,92,100\}$. In this case, we compare our results with those from Ref. Yerokhin and Shabaev (2015a). Therefore, the nuclear parameters are taken from Ref. Yerokhin and Shabaev (2015a). Average values of rms nuclear radii from there were used. An estimation of the uncertainty of the SE correction with respect to the nuclear radius and nuclear model is beyond the scope of the current paper, the interested reader can find it in the original work Yerokhin and Shabaev (2015a). The values of fundamental constants, however, have been adopted from Ref. Tiesinga et al. (2021). The corresponding results are summarized in the Table 3, which is organized similarly to the Table 2. Overall, the results shown in Table 3 confirm the findings obtained earlier. The relative smallness of the Coulomb-gauge many-potential contribution is illustrated by Fig. 9(a), where different contributions to the SE correction for the $1S_{1/2}$ state in both gauges are plotted as functions of $Z$. This smallness as well as the absence of cancellation between the momentum-space contributions in the Coulomb gauge is demonstrated also in Fig. 9(b), where the sum of the zero- and one-potential contributions is shown. For neon ($Z=10$), a more ‘‘regular’’ behavior of the zero- and one-potential contributions in the Coulomb gauge is even more pronounced than for higher-$Z$ systems. From Table 3, it is seen that our results are in reasonable agreement with the results from Ref. Yerokhin and Shabaev (2015a). We conclude that the most accurate calculations of the many-potential contributions can indeed be performed using the combination of the Coulomb gauge and the SC convergence-acceleration scheme.
Using the most effective methods among those examined, namely, the GF approach and the SC convergence-acceleration scheme, we consider the SE corrections for the $n=2$ states: $2S_{1/2}$, $2P_{1/2}$, and $2P_{3/2}$. The SE corrections are calculated for hydrogen-like ions with $Z\in\{10,18,26,36,54,82,92,100\}$. The corresponding results for both Feynman and Coulomb gauges are given in Table LABEL:tab:we_yerokh_excited_states. Analyzing the table one can see that the accuracy of the total SE corrections for the excited states in hydrogen-like ions is still limited by the uncertainty of the many-potential contribution calculations. As in the case of the ground state, this contribution in the Coulomb gauge is significantly smaller than in the Feynman gauge. Moreover, the Coulomb gauge as before lacks strong cancellations between the momentum-space contributions. Finally, we note that our results are found to be in good agreement with the ones from Ref. Yerokhin and Shabaev (2015a).
V Conclusion
The self-energy correction for various states of hydrogen-like ions is considered. Two gauges, Feynman and Coulomb ones, are studied, the acceleration tricks are discussed, and a self-consistent set of useful for the corresponding calculations formulas are collected.
The acceleration schemes do improve the accuracy of calculations. The combination of the Coulomb gauge and the Sapirstein-Cheng partial-wave-expansion convergence-acceleration scheme Sapirstein and Cheng (2023) is most suitable (among those considered) for calculating the self-energy correction in hydrogen-like ions, especially when studying light nuclei.
We believe that the analysis performed in the present work will facilitate the calculation of a wide range of more complex radiative corrections with the self-energy loop.
VI Acknowledgments
This work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics ‘‘BASIS’’.
Appendix A Free-electron self-energy operator
We define the free-electron self-energy operator as:
[…]
$D_{\mu\nu}^{\text{F}}(\mathrm{k})=-\frac{g_{\mu\nu}}{\mathrm{k}^{2}}$ (52)
in the Feynman gauge and
$D_{\mu\nu}^{\text{C}}(\mathrm{k})=\frac{1}{\mathrm{k}^{2}}\left(-g_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{\mathbf{k}^{2}}+\frac{k^{0}(k_{\mu}\delta_{\nu 0}+k_{\nu}\delta_{\mu 0})}{\mathbf{k}^{2}}\right)$ (53)
in the Coulomb gauge. In Eq. (53), $\delta_{\nu\mu}$ is the four-dimensional Kronecker delta, $\delta_{\mu\nu}\equiv\delta^{\mu}_{\nu}$. The expression (51) is UV divergent in both gauges. To eliminate the UV divergences, we apply the renormalization procedure. To make sense of divergent expressions, we use the dimensional regularization $d=4-2\varepsilon$. In both gauges, we separate out the UV divergent part by writing the self-energy in the form:
[…]
where $\Delta_{\varepsilon}=1/\varepsilon-\gamma_{E}+\ln 4\pi+\ln m_{e}$, $\gamma_{E}$ is the Euler constant, and the mass counterterm is
$\delta m=\frac{3\alpha}{4\pi}m_{e}\left(\Delta_{\varepsilon}+\frac{4}{3}\right).$ (55)
Note that following Ref. Yerokhin and Shabaev (1999) we use this definition of $\Sigma^{(0)}_{R}$ instead of writing a more familiar expression in terms of the renormalization constant $Z_{2}$. In the Feynman gauge, the second approach leads to the IR singularities of individual contributions. In the Coulomb gauge, both approaches are equally convenient, but, for consistency, we write the expression for the free-electron self-energy operator in this form.
The renormalized part of the free-electron self-energy operator on the right-hand side of Eq. (54) can be written in the form:
[…]
The coefficients $a$, $b$ and $c$ are gauge-dependent. In the Feynman gauge, they can be found, e.g., in Ref. Yerokhin and Shabaev (1999):
$\begin{gathered}a(p_{0},p)=2m_{e}\left(1+\frac{2\rho}{1-\rho}\ln\rho\right),\\ b(p_{0},p)=-\frac{2-\rho}{1-\rho}\left(1+\frac{\rho}{1-\rho}\ln\rho\right),\\ c(p_{0},p)=0,\end{gathered}$ (57)
where $\rho=1-{\mathrm{p^{2}}}/{m_{e}^{2}}$. For the Coulomb gauge, they could be derived from the expression obtained in Ref. Adkins (1983):
[…]
where
$\begin{gathered}X=1+\frac{p^{2}}{m_{e}^{2}}(1-x),\\ Y=1-\frac{\mathrm{p}^{2}}{m_{e}^{2}}(1-x)-i0,\\ Z=1-\frac{p_{0}^{2}}{m_{e}^{2}}(1-u)+\frac{p^{2}}{m_{e}^{2}}(1-xu)-i0.\end{gathered}$ (59)
Some integrals in these expressions can be evaluated analytically, which was done, e.g., in Ref. Malenfant (1987). The latter results allows one to write for the coefficients $a$, $b$, and $c$:
$\begin{gathered}a(p_{0},p)=2m_{e}\left(1-F_{0}+\frac{\rho\log\rho}{1-\rho}\right),\\ b(p_{0},p)=\frac{(\rho-2)(1-\rho+\rho\ln\rho)}{(1-\rho)^{2}}-2F_{2}\rho+\frac{2m_{e}^{2}(F_{1}\rho\ln\rho-F_{0})}{p^{2}},\\ c(p_{0},p)=2\frac{p_{0}}{p^{2}}\big(F_{0}m_{e}^{2}-F_{1}m_{e}^{2}\rho\ln\rho+F_{2}\rho p^{2}\big),\end{gathered}$ (60)
where $F_{0}$, $F_{1}$, and $F_{2}$ are defined according to
$\begin{gathered}F_{0}=\frac{\sqrt{p^{2}+m_{e}^{2}}}{p}\ln\left|\frac{\sqrt{p^{2}+m_{e}^{2}}+p}{\sqrt{p^{2}+m_{e}^{2}}-p}\right|-2,\\ F_{1}=\frac{p_{0}}{p}\ln\left|\frac{p_{0}+p}{p_{0}-p}\right|-2,\\ F_{2}=\int_{0}^{1}dx\frac{\sqrt{x}\ln X}{X-\rho+i0}.\end{gathered}$ (61)
Appendix B Free-electron vertex operator
In order to calculate the one-potential term, one needs the free-electron vertex function given by:
[…]
This expression is UV divergent in both gauges. Using dimensional regularization, it can be written as:
$\Gamma^{\mu}(\mathrm{p}^{\prime},\mathrm{p})=\frac{\alpha}{4\pi}\Delta_{\varepsilon}\gamma^{\mu}+\Gamma_{R}^{\mu}(\mathrm{p}^{\prime},\mathrm{p}),$ (63)
where, as in the case of the free-electron self-energy operator, the UV finite part of $\Gamma_{R}$ also turns out to be IR finite. For brevity, we discuss below only the time component of the operator, $\Gamma^{0}_{R}$, for the case when $p_{0}=p^{\prime}_{0}=\varepsilon_{a}$:
[…]
In the Feynman gauge:
$\begin{gathered}A=C_{5}-2+\mathrm{p}^{\prime 2}C_{11}+\mathrm{p}^{2}C_{12}+4(\mathrm{p}^{\prime}\cdot\mathrm{p})(C_{00}+C_{11}+C_{12})+m_{e}^{2}(-2C_{00}+C_{11}+C_{12}),\\ B_{1}=-4(C_{11}+C_{23}),\;B_{2}=-4(C_{00}+C_{11}+C_{12}+C_{25}),\\ C_{1}=B_{2},\;C_{2}=-4(C_{12}+C_{24}),\\ D=2(C_{00}+C_{11}+C_{12}),\\ H_{1}=4m_{e}(C_{00}+2C_{11}),\;H_{2}=4m_{e}(C_{00}+2C_{12}),\;G_{1}=G_{2}=0,\end{gathered}$ (65)
where
$\begin{gathered}C_{ij}=\int_{0}^{1}\frac{dy}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}}S_{i}K_{j},\\ S_{0,1,2}=\{-\ln X^{\prime},1-Y^{\prime}\ln X^{\prime},-\frac{1}{2}+Y^{\prime}-{Y^{\prime}}^{2}\ln X^{\prime}\},\\ K_{0,1,2,3,4,5}=\{1,y,1-y,y^{2},(1-y)^{2},y(1-y)\},\\ C_{5}=-\int_{0}^{1}dy\ln(y^{2}\mathrm{k}^{2}/\penalty 50m_{e}^{2}-y\mathrm{k}^{2}/m_{e}^{2}+1),\\ X^{\prime}=1+\frac{1}{Y^{\prime}},\;Y^{\prime}=\frac{m_{e}^{2}-y\mathrm{p}^{\prime 2}-(1-y)\mathrm{p}^{2}}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}},\end{gathered}$ (66)
and $\mathrm{k}=\mathrm{p}-\mathrm{p}^{\prime}$. For the Coulomb gauge, these coefficients can be readily determined from Ref. Holmberg (2011):
[…]
and finally $A$:
$\begin{gathered}A=\varepsilon_{a}^{2}(2F_{1}-F_{2})-F_{22}+m_{e}^{2}(2F_{1}-3F_{2})+p^{\prime 2}(F_{11}-F_{8}+4F_{5}-5F_{3})-\\ -4(\mathbf{p^{\prime}}\cdot\mathbf{p})^{2}F_{13}+p^{2}(F_{10}-F_{11}-5F_{2}+5F_{3}+4F_{4}-4F_{5}-F_{7}+\\ +F_{8}+(\mathbf{p^{\prime}}\cdot\mathbf{p})(-2F_{12}+4F_{13}))+(-4F_{5}+4F_{6}+2F_{8}-2F_{9})k^{2}+\\ +(\mathbf{p^{\prime}}\cdot\mathbf{p})(4F_{1}-2F_{19}-2F_{2}-2F_{14}k^{2})-\varepsilon_{a}^{2}(B_{1}+C_{1}-D).\end{gathered}$ (68)
The coefficients $F_{1-22}$ introduced above can be determined as follows. Let us define $B_{i,j}^{k}$ as
$B_{i,j}^{k}=\begin{cases}1\,,&\text{if }i\leq k\leq j\,,\\ 0\,,&\text{otherwise}\,.\end{cases}$ (69)
then for $F_{1}-F_{6}$ one obtains
$\begin{gathered}F_{i}=\int_{0}^{1}\frac{du}{\mathrm{t}^{2}}\Big[C_{1}^{i}\delta_{1}+C_{2}^{i}-C_{3}^{i}\frac{A}{\mathrm{t}^{2}}\Big]C_{4}^{i},\\ \delta_{1}=\ln\left(\frac{\mathrm{t}^{2}+A}{A}\right),\,\,C_{1}^{i}=[1-B_{2,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)][1-B_{4,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)],\\ C_{2}^{i}=B_{2,3}^{i}+\frac{1}{2}B_{4,6}^{i},\,\,C_{3}^{i}=B_{4,6}^{i},\,\,C_{4}^{i}=\{1,1,u,1,u,u^{2}\}.\end{gathered}$ (70)
Here, $A=m_{e}^{2}-u\mathrm{p}^{2}-(1-u)\mathrm{p}^{2}$, $\mathrm{t}=u\mathrm{p}^{\prime}+(1-u)\mathrm{p}$. For $F_{7-11}$ (below the coefficient $F_{7}$ is obtained by substituting $i=1$, and so on, i.e., $F_{8}\leftrightarrow i=2$, …):
$\begin{gathered}F_{i}=\int_{0}^{1}du\Big[C_{5}^{i}\delta_{2}-C_{6}^{i}\frac{1}{t^{2}}\Big]C_{7}^{i},\\ \delta_{2}=\frac{2}{\sqrt{t^{2}C}}\tanh^{-1}\left[\left(\frac{t^{2}}{C}\right)^{1/2}\right],\,\,C_{5}^{i}=\frac{C}{t^{2}}B_{1,3}^{i}+B_{4,5}^{i},\\ C_{6}^{i}=2B_{1,3}^{i},\,\,C_{7}^{i}=\{1,u,u^{2},1,u\},\end{gathered}$ (71)
where $C=m_{e}^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+t_{0}^{2}$. For $F_{12}-F_{21}$ (below $F_{12}\leftrightarrow i=1$):
$\begin{gathered}F_{i}=\int_{0}^{1}dsdu\Big[C_{8}^{i}\delta_{3}+C_{9}^{i}\delta_{4}\Big]C_{10}^{i},\\ \delta_{3}=\frac{1}{s(t^{2})^{2}}\left\{\left(\frac{B}{st^{2}}\right)^{1/2}\tanh^{-1}\Big[\left(\frac{st^{2}}{B}\right)^{1/2}\Big]-1\right\},\,\,\delta_{4}=\frac{1}{t^{2}(B-st^{2})},\\ C_{8}^{i}=-3B_{1,7}^{i}+2t^{2}B_{8,10}^{i},\,\,C_{9}^{i}=B_{1,7}^{i},\\ C_{10}^{i}=\{1,u,u^{2},s,su,su^{2},su^{3},1,s,su\},\end{gathered}$ (72)
where $B=m_{e}^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+st_{0}^{2}$. And, finally, $F_{22}$ is given by
$F_{22}=\int_{0}^{1}du\ln(u^{2}\mathrm{k}^{2}/\penalty 50m_{e}^{2}-u\mathrm{k}^{2}/\penalty 50m_{e}^{2}+1).$ (73)
Appendix C Many-potential term
Reduced matrix elements of the operators $I_{\text{F}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ and $I_{\text{C}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ can be derived using the PW expansion of the photon propagator. Namely, for this aim one needs the standard expansion formulas Varshalovich et al. (1988) for the functions which depend on $r_{12}$, such as
$\frac{e^{i\omega r_{12}}}{r_{12}}=4\pi i\omega\sum_{L=0}^{\infty}\sum_{M=-L}^{L}j_{L}(\omega r_{<})h_{L}^{(1)}(\omega r_{>})Y_{LM}^{*}(\mathbf{\hat{r}_{1}})Y_{LM}(\mathbf{\hat{r}_{2}}).$ (74)
Here, $j_{L}$ and $h_{L}^{(1)}$ are the spherical Bessel function and spherical Hankel functions of the first kind, respectively and $r_{>}=\max(r_{1},r_{2})$ and $r_{<}=\min(r_{1},r_{2})$. Applying these formulas and performing the integrations over the angular variables, after some tedious manipulations one can obtain
$\begin{gathered}\langle ab||I_{\text{F}}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}r_{1}^{2}\int_{0}^{\infty}dr_{2}r_{2}^{2}\,\Big\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(\hat{\omega},r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}[J]g_{L}(\hat{\omega},r_{1},r_{2})D^{JL}_{ac}(r_{1})D^{JL}_{bd}(r_{2})\Big\}\end{gathered}$ (75)
for the Feynman gauge, and
$\begin{gathered}\langle ab||I_{\text{C}}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}r_{1}^{2}\int_{0}^{\infty}dr_{2}r_{2}^{2}\,\Big\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(0,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}a_{JL}g_{L}(\hat{\omega},r_{1},r_{2})D^{JL}_{ac}(r_{1})D^{JL}_{bd}(r_{2})\\ +(-1)^{J+1}b_{J}[g_{J}^{\text{ret}}(\hat{\omega},r_{1},r_{2})D^{JJ+1}_{ac}(r_{1})D^{JJ-1}_{bd}(r_{2})+g_{J}^{\text{ret}}(\hat{\omega},r_{2},r_{1})D^{JJ-1}_{ac}(r_{1})D^{JJ+1}_{bd}(r_{2})]\Big\}\end{gathered}$ (76)
for the Coulomb gauge. In the above expressions, $[J]=2J+1$ and the following functions are defined:
$g_{J}(\omega,r_{1},r_{2})$ $=i[J]\omega j_{J}({\omega}r_{<})h_{J}^{(1)}({\omega}r_{>}),$ (77)
$g_{J}(0,r_{1},r_{2})$ $=\frac{r_{<}^{J}}{r_{<}^{J+1}},$ (78)
[…]
The expressions depending on the radial wave functions read as
$A_{ab}(r)$ $=g_{a}(r)g_{b}(r)+f_{a}(r)f_{b}(r)\,$, (80)
$D^{JL}_{ab}(r)$ $=g_{a}(r)f_{b}(r)H_{L}^{J}(\kappa_{a},-\kappa_{b})-f_{a}(r)g_{b}(r)H_{L}^{J}(-\kappa_{a},\kappa_{b})\,.$ (81)
Finally, the angular coefficients are
[…]
$H_{L}^{J}(\kappa_{a},\kappa_{b})$ $=(-1)^{l_{a}}\sqrt{6[j_{a}][j_{b}][l_{a}][l_{b}]}{\begin{pmatrix}l_{a}&L&l_{b}\\ 0&0&0\end{pmatrix}}{\begin{Bmatrix}j_{a}&\frac{1}{2}&l_{a}\\ J&1&L\\ j_{b}&\frac{1}{2}&l_{b}\end{Bmatrix}},$ (83)
$a_{JL}$ $=\begin{cases}J+1,&\text{for}\ L=J-1,\\ 2J+1,&\text{for}\ L=J,\\ J,&\text{for}\ L=J+1,\end{cases}$ (84)
$b_{J}$ $=\sqrt{J(J+1)}\frac{\sqrt{[J+1][J-1]}}{[J]}.$ (85)
Note that (76) is consistent with Ref. Yerokhin et al. (2021), except for the typo in Eq. (B6) there.
Appendix D Derivative of the free-electron self-energy operator
Let us represent the second derivative of the function $\Sigma^{(0)}_{R}$ with respect to the time component of the four-vector $\mathrm{p}$ in the form:
$\frac{\partial^{2}\Sigma_{R}^{(0)}(\mathrm{p})}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}=\frac{\alpha}{4\pi}\left\{(N_{1}+N_{2}\gamma^{0}-N_{3}(\boldsymbol{\gamma}\cdot\mathbf{p})\right\},$ (86)
where the coefficients $N_{1}$, $N_{2}$, and $N_{3}$ can be obtained by differentiating the formula (56):
$\begin{gathered}N_{1}=\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=\varepsilon_{a}\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial b(p_{0},p)}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial^{2}c(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{3}=\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}.\end{gathered}$ (87)
In the Feynman gauge, the derivatives of the coefficients $a$ and $b$ are (see Sapirstein and Cheng (2023); Malyshev et al. (2024)):
$\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}=-\frac{8}{m_{e}(1-\rho)}\left\{1+\frac{1}{1-\rho}\left[\ln\rho-\frac{2p_{0}^{2}}{m_{e}^{2}}\left(\frac{1+\rho}{\rho}+\frac{2}{1-\rho}\ln\rho\right)\right]\right\},$ (88)
$\frac{\partial b(p_{0},p)}{\partial p_{0}}=\frac{2p_{0}}{m_{e}^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\ln\rho\right\},$ (89)
$\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}=\frac{2}{m_{e}^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\left[\ln\rho-\frac{p_{0}^{2}}{m_{e}^{2}}\left(\frac{2+5\rho-\rho^{2}}{\rho}+\frac{6}{1-\rho}\ln\rho\right)\right]\right\}.$ (90)
In the Coulomb gauge, it is more convenient to differentiate the free-electron self-energy operator in the form (58):
[…]
Then, for the coefficients $N_{1}$ - $N_{3}$ we have:
$\begin{gathered}N_{1}=-2m_{e}\int_{0}^{1}dx\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=2\int_{0}^{1}dx(1-x)\left(\varepsilon_{a}\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+2\frac{\partial\ln Y}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\\ N_{3}=2\int_{0}^{1}dx\left((1-x)\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}-\sqrt{x}\int_{0}^{1}du\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\end{gathered}$ (92)
where
$\begin{gathered}\frac{\partial\ln Y}{\partial p_{0}}=\frac{2p_{0}(x-1)}{m_{e}^{2}-(p_{0}^{2}-p^{2})(1-x)-i0},\\ \frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}=\frac{2(x-1)(m_{e}^{2}+(p_{0}^{2}+p^{2})(1-x))}{(m_{e}^{2}-(p_{0}^{2}-p^{2})(1-x)-i0)^{2}},\\ \frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}=\frac{2(u-1)(m_{e}^{2}+p_{0}^{2}(1-u)+p^{2}(1-xu))}{(m_{e}^{2}-p_{0}^{2}(1-u)+p^{2}(1-xu)-i0)^{2}}.\end{gathered}$ (93)
For the coefficient $Z$, the second derivative can be integrated analytically with respect to $u$:
$\begin{gathered}\int_{0}^{1}du\left(\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\right)=2\Big[\frac{p_{0}^{4}-3p_{0}^{2}(m_{e}^{2}+p^{2})-xp^{2}(m_{e}^{2}+p^{2}-3p_{0}^{2})}{m_{e}^{2}\rho(p_{0}^{2}-xp^{2})^{2}}\\ -\frac{(m_{e}^{2}+p^{2}(1-x))(3p_{0}^{2}+xp^{2})}{(p_{0}^{2}-xp^{2})^{3}}\ln\left(\frac{m_{e}^{2}+p^{2}-p_{0}^{2}}{m_{e}^{2}+p^{2}(1-x)}\right)\Big].\end{gathered}$ (94)
Some integrals over $x$ from the expressions (92) can be evaluated analytically using the following master integrals:
$\begin{gathered}\int_{0}^{1}dy\frac{y}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{p_{0}^{2}-p^{2}}\left(\frac{\ln\rho}{p_{0}^{2}-p^{2}}+\frac{1}{m_{e}^{2}\rho}\right),\\ \int_{0}^{1}dy\frac{y^{2}}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{m_{e}^{2}}{(p_{0}^{2}-p^{2})^{3}}\left(\frac{1}{\rho}+2\ln\rho-\rho\right),\\ \int_{0}^{1}dy\frac{y^{3}}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{(p_{0}^{2}-p^{2})^{4}}(p_{0}^{4}-2(p_{0}^{2}+2m_{e}^{2})p^{2}+\frac{2m_{e}^{4}}{\rho}+6m_{e}^{4}\ln\rho\\ +4m_{e}^{2}p_{0}^{2}+(p^{2})^{2}-2m_{e}^{4}),\\ \int_{0}^{1}dy\frac{y^{2}}{m_{e}^{2}-(p_{0}^{2}-p^{2})y}=-\frac{1}{2}\left(\frac{1}{p_{0}^{2}-p^{2}}+\frac{2m_{e}^{2}}{(p_{0}^{2}-p^{2})^{2}}+\frac{2m_{e}^{4}\ln\rho}{(p_{0}^{2}-p^{2})^{3}}\right).\end{gathered}$ (95)